### v2

# Qiang Tan, Hongyu Wang, Ken Wang

Abstract: In this paper, we define $\mathcal{D}^{+}_{J}$ operator that is a generalized $\partial\bar{\partial}$ operator on higher dimensional almost Kähler manifolds. In terms of $\mathcal{D}^{+}_{J}$ operator, we study $\bar{\partial}$-problem in almost Kähler geometry and the generalized Monge-Ampère equation on almost Kähler manifolds. Similarly to the Kähler case, we obtain $C^{\infty}$ a priori estimates for the solution of the generalized Monge-Ampère equation on the almost Kähler manifold $(M,\omega,J)$ depended only on $\omega$ and $J$. Then as done in Kähler geometry, we study Calabi conjecture for almost Kähler manifolds. Finally, we will pose some questions in almost Kähler geometry.
AMS Classification (2000): 53D35; 53C56; 53C65; 32Q60.
Keywords: almost Kähler form, $\mathcal{D}^{+}_{J}$ operator, $\bar{\partial}$-problem in almost Kähler geometry, the generalized Monge-Ampère equation.
1 Introduction
Yau’s Theorem for closed Kähler manifolds [72] states that one can prescribe the volume form of a Kähler metric within a given Kähler class. This result occupies a central place in the theory of Kähler manifolds (cf. Calabi [9, 10]), with wide-ranging applications in geometry and mathematical physics (cf. Yau [71] ).
More precisely, Yau’s Theorem is as follows. Let $(M^{2n},\omega,J,g_{J})$ be a closed Kähler manifold of real dimension $2n$, where $\omega$ denotes the Kähler form, $J$ an $\omega$-compatible complex structure and $g_{J}(\cdot,\cdot)=\omega(\cdot,J\cdot)$. Then given a smooth volume form $e^{F}\omega^{n}$ on $M^{2n}$ there exists a unique smooth function $\varphi$ satisfying
$(\omega+\sqrt{-1}\partial_{J}\bar{\partial}_{J}\varphi)^{n}$ $=$ $e^{F}\omega^{n},$
$\omega+\sqrt{-1}\partial_{J}\bar{\partial}_{J}\varphi$ $>$ $0\displaystyle 0,\,\,\,\sup_{M^{2n}}\varphi=$ 0 (1.1)
as long as $F$ satisfies the necessary normalization condition that
$\int_{M^{2n}}e^{F}\omega^{n}=\int_{M^{2n}}\omega^{n}.$ (1.2)
Equation (1) is known as the complex Monge-Ampère equation for Kähler manifolds. It is based on the method of continuity and in any given situation requires laborious a priori estimates for the derivatives of the solution $\varphi$ up to third order (cf. S.-T. Yau [72]). The uniqueness of the solution $\varphi$ has an elementary proof found by Calabi [10].
A different type of extension of Yau’s Theorem is to the case when $J$ is a non-integrable almost complex structure. Recall that an almost complex structure $J$ is said to be tamed by symplectic from $\omega$ when the bilinear form $\omega(\cdot,J\cdot)$ is positive definite. The almost complex structure $J$ is said to be compatible (or calibrate) with $\omega$ when the same bilinear form is also symmetric, that is, $\omega(\cdot,J\cdot)>0$ and $\omega(J\cdot,J\cdot)=\omega(\cdot,\cdot)$ (cf. McDuff-Salamon [45]). In 1990’s, Gromov posed the following problem to P. Delanoë [20]: Let $(M^{2n},\omega)$ be a closed symplectic manifold, $J$ an almost complex structure compatible with $\omega$ and $F$ a smooth function on $M^{2n}$ with
$\int_{M^{2n}}e^{F}\omega^{n}=\int_{M^{2n}}\omega^{n}.$
Can one find a smooth function $\varphi$ on $M$ such that $\omega+dJd\varphi$ is a symplectic form taming $J$ and satisfying
$(\omega+dJd\varphi)^{n}=e^{F}\omega^{n}?$ (1.3)
However, P. Delanoë [20] showed that when $n=2$ the answer to this question is negative, this was later extended to all dimensions by Wang-Zhu [65]. The key ingredient of their results in the construction of a smooth function such that $\omega+dJd\varphi_{0}$ is on the boundary of the set of taming symplectic forms (so its $(1,1)$ part is semipositive definite but not strictly positive), and yet $(\omega+dJd\varphi_{0})^{n}>0$. This is possible because in this case the $(2,0)+(0,2)$ part of $dJd\varphi_{0}$ contributes a strictly positive amount. This indicate that the problem with Gromov’s suggestion is that the 2-form $dJd\varphi$ is in general not of type $(1,1)$ with respect to $J$, due to the fact that $J$ may not be integrable. Its $(1,1)$ part will be denoted by
$\sqrt{-1}\partial_{J}\bar{\partial}_{J}\varphi=\frac{1}{2}(dJd\varphi)^{(1,1)}$
which agrees with the standard notation when $J$ is integrable.
Therefore, one can consider the Monge-Ampère equation for non-integrable almost complex structures. There are many important works when the base manifolds are almost complex manifolds. We refer to Harvey-Lawson [35], Tosatti-Wang-Weinkove-Yang [60], Chu-Tosatti-Weinkove [15], and the references therein.
In order to study Donaldson tameness question [26], Tan-Wang-Zhou-Zhu introduced operator $\mathcal{D}^{+}_{J}$ on tamed closed almost complex four-manifold $(M^{4},J)$ in [53]. Then by using $\mathcal{D}^{+}_{J}$, they solved this question with the condition of $h_{J}^{-}=b^{+}-1$. The operator $\mathcal{D}^{+}_{J}$ can be viewed as a generalization of $\partial_{J}\bar{\partial}_{J}$. Specifically, if $J$ is integrable, $\mathcal{D}^{+}_{J}=2\sqrt{-1}\partial_{J}\bar{\partial}_{J}$. Wang-Wang-Zhu [64], by using $\mathcal{D}^{+}_{J}$ operator, studied Nakai-Moishezon criterion for tamed almost Hermitian $4$-manifolds. In [66], Wang-Zhang-Zheng-Zhu studied the generalized Monge-Ampère equation
$(\omega+\mathcal{D}^{+}_{J}(f))^{2}=e^{F}\omega^{2}$
on almost Kähler surfaces and established a uniqueness and existence theorem for it.
It is well known that, for four dimension case, the operator $d^{-}_{J}d^{*}$ is a self-adjoint strongly elliptic linear operator and plays a key role in defining the operator $\mathcal{D}^{+}_{J}$. However, in high-dimensional situations, $d^{-}_{J}d^{*}$ is no longer elliptical. Fortunately, by direct calculation, we find that the principal symbol of $d^{-}_{J}d^{*}$ is the same as $\partial_{J}\partial_{J}^{*}+\bar{\partial}_{J}\bar{\partial}_{J}^{*}$ on higher dimensional almost Kähler manifolds. Thus, we get the operator $d^{-}_{J}d^{*}:{\rm D}(d^{-}_{J}d^{*})\longrightarrow{\rm R}(d^{-}_{J}d^{*})$ is an invertible formal self-adjoint and nonnegative operator. Hence, for any $f\in C^{\infty}(M^{2n})_{0}$, we have $d^{-}_{J}Jdf\,\bot\,{\rm coker}(d^{-}_{J}d^{*})$, and there exists a unique $\alpha_{f}\in{\rm D}(d^{-}_{J}d^{*})$ such that
$d^{-}_{J}d^{*}(\alpha_{f})=d^{-}_{J}Jdf.$
Then, using $d^{-}_{J}d^{*}$, we can also define $\mathcal{D}^{+}_{J}$ operator on higher dimensional almost Kähler manifolds by
$\mathcal{D}^{+}_{J}(f)$ $=$ $dJdf+dd^{*}\alpha_{f}$
$=$ $dd^{*}(f\omega)+dd^{*}\alpha_{f}$
satisfying $d^{-}_{J}d^{*}(f\omega)+d^{-}_{J}d^{*}\alpha_{f}=0$. Let $\mathcal{W}_{J}(f)=d^{*}(f\omega)+d^{*}\sigma(f)$, then
$\left\{\begin{array}[]{ll}d\mathcal{W}_{J}(f)=\mathcal{D}^{+}_{J}(f)&,\\ &\\ d^{*}\mathcal{W}_{J}(f)=0&,\end{array}\right.$ (1.4)
which is an elliptic system. Notice that $d_{J}^{-}\mathcal{W}_{J}(f)=0$, as $\bar{\partial}$-problem in classical complex analysis [36], $(\mathcal{W}_{J},d_{J}^{-})$-problem is whether $\mathcal{W}_{J}(f)=A$ has a solution for any $A$ satisfying $d_{J}^{-}A=0$. By using $L^{2}$-method [36], calculating the $L^{2}$-norm of $\mathcal{W}_{J}^{*}A$ and applying Riesz Representation Theorem [74], we solve $(\mathcal{W}_{J},d^{-}_{J})$-problem in Section 2 (see Theorem 2.5), which plays a very important role during the research process of the following generalized Monge-Ampère equation.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. We can define a generalized Monge-Ampère equation on $M^{2n}$,
$(\omega+\mathcal{D}^{+}_{J}(f))^{n}=e^{F}\omega^{n}$ (1.5)
for some real function $f\in C^{\infty}(M,J)_{0}$ with
$\omega+\mathcal{D}^{+}_{J}(f)>0,$
where $F\in C^{\infty}(M)$ satisfies
$\int_{M^{2n}}\omega^{n}=\int_{M^{2n}}e^{F}\omega^{n}.$ (1.6)
Then we establish a uniqueness and local existence theorem for it in Section 3 (see Theorem 3.6 and Theorem 3.5).
In Section 4, we devote to considering the global existence theorem for the generalized Monge-Ampère equation (1.5) on almost Kähler manifolds. Let $(M^{2n},\omega,J,g)$ be a closed almost Kähler manifold and $\omega_{1}=\omega+\mathcal{D}^{+}_{J}(f)$ another $J$-compatible almost Kähler form on $M^{2n}$ satisfying $\omega_{1}\in[\omega]$ and $\omega_{1}^{n}=e^{F}\omega^{n}$. Define a smooth function $f_{0}\in C^{\infty}(M,J)_{0}$ as follows:
$-\frac{1}{2n}\Delta_{g}f_{0}=\frac{\omega^{n-1}\wedge(\omega_{1}-\omega)}{\omega^{n}},$ (1.7)
where $\Delta_{g}$ is the Laplacian of the Levi-Civita connection with respect to the almost Kähler metric $g$. Using the result of Tosatti-Weinkove-Yau [61, Lemma 2.5], we can easily obtain
$-\omega^{n-1}\wedge dJdf_{0}=\frac{1}{2n}\Delta_{g}f_{0}\cdot\omega^{n}.$
Then, by Proposition A.1, we can find $a_{0}(f)\in\Omega^{1}(M)$ such that
$\mathcal{D}^{+}_{J}(f)=dJdf_{0}+da_{0}(f),$
and
$\left\{\begin{array}[]{ll}d^{*}a_{0}(f)=0&,\\ &\\ d^{-}_{J}Jdf_{0}+d^{-}_{J}a_{0}(f)=0&,\\ &\\ \omega^{n-1}\wedge da_{0}(f)=0&,\end{array}\right.$ (1.8)
where $d^{*}=-*_{g}d*_{g}$ and $*_{g}$ is the Hodge star operator with respect to the metric $g$. Hence, we have
$\omega_{1}-\omega$ $=$ $\mathcal{D}^{+}_{J}(f)$
$=$ $dJdf_{0}+da_{0}(f)$
$=$ $dd^{*}(f_{0}\omega)+da_{0}(f)$
$=$ $\mathcal{D}^{+}_{J,0}(f_{0}).$
Then we obtain our first main theorem ( Theorem 4.13 in Section 4),
Theorem 1.1.
Let $(M^{2n},\omega,J,g)$ be a closed almost Kähler manifold. Let $F$ be a real function on $M^{2n}$ with
$\int_{M^{2n}}\omega^{n}=\int_{M^{2n}}e^{F}\omega^{n}.$
Then if $\omega_{1}$ is an almost Kähler form with $[\omega_{1}]=[\omega]$ and solving the Calabi-Yau equation
$\omega_{1}^{n}=e^{F}\omega^{n},$
there are $C^{\infty}$ a priori bounds on $\omega_{1}$ depending only on $M^{2n},\omega,J$ and $F$.
Then, by classical continuity method, we have the following result (Theorem 4.15 in Section 4),
Theorem 1.2.
Let $(M^{2n},\omega,J,g)$ be a closed almost Kähler manifold of dimension $2n$. Then for $F\in C^{\infty}(M^{2n})$ satisfying
$\int_{M^{2n}}e^{F}\omega^{n}=\int_{M^{2n}}\omega^{n},$
there exists a unique almost Kähler form $\tilde{\omega}=\omega+\mathcal{D}^{+}_{J}(f)$ cohomologous to $\omega$ solving
$\tilde{\omega}^{n}=e^{F}\omega^{n},$
and there are $C^{\infty}$ a priori bounds on $\tilde{\omega}$ depending only on $\omega,J,F$.
In Section 5, we consider the generalized Calabi conjecture for almost Kähler setting. Let $M$ be a compact Kähler manifold with Kähler metric $\sum g_{i\bar{j}}dz^{i}\otimes d\bar{z}^{j}$. In [9], Calabi conjectured the existence of Kähler-Einstein metric on a Kähler manifold whose first Chern class is negative, zero or positive. Kähler Einstein metrics are the metrics whose Ricci form is proportional to the Kähler form. If the first Chern class is negative, then definition we can choose Kähler metric so that $\frac{\sqrt{-1}}{2\pi}\sum g_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ represents $-c_{1}(M)$. If $c_{1}(M)$ is zero, we choose $\frac{\sqrt{-1}}{2\pi}\sum g_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ to be arbitrary Kähler form. If $c_{1}(M)$ is positive, we choose $\frac{\sqrt{-1}}{2\pi}\sum g_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ to represent $c_{1}(M)$. In these case, the existence of Kähler-Einstein metric is equivalent to solve the following equation
$\det(g_{i\bar{j}}+\frac{\partial^{2}\varphi}{\partial z_{i}\partial\bar{z}_{j}})\det(g_{i\bar{j}})^{-1}=\exp(c\varphi+F),$ (1.9)
where $c=1$, 00 or $-1$ and $F$ is a smooth function defined on $M$. If $c_{1}(M)<0$, there exists a Kähler-Einstein metric (cf. Aubin [4], Yau [72]). If $c_{1}(M)=0$, the existence of Kähler-Einstein metric is equivalent the existence of the solution of the following complex Monge-Ampère equation
$\det(g_{i\bar{j}}+\frac{\partial^{2}\varphi}{\partial z_{i}\partial\bar{z}_{j}})\det(g_{i\bar{j}})^{-1}=\exp(F),$
where $F$ is a smooth function defined on $M$ satisfying Equation (5.2). For $c_{1}(M)>0$ case which is called Yau-Tian-Donaldson conjecture ([73, 56, 25]). In 2015, Chen-Donaldson-Sun [14] and Tian [57] solved Yau-Tian-Donaldson conjecture independently.
It is nature to consider Hermite-Einstein almost Kähler metrics. An almost Kähler metric $(\omega,J,g)$ is called Hermite-Einstein (HEAK for short) if the Hermite-Ricci form $\rho$ is a (constant) multiple of the symplectic form $\omega$, i.e.
$\rho=\frac{R}{2n}\omega,$
so the Hermitian scalar curvature $R$ is constant (cf. [41]). Similar to Kähler case, we have the following theorem for HEAK metrics (Theorem 5.4),
Theorem 1.3.
If $c_{1}(M^{2n},J)<0$, there exists a HEAK metric unique up to an homothety. If $c_{1}(M^{2n},J)=0$, there exists a HEAK metric unique up to an homothety in each positive $(1,1)$ cohomology chass.
If $c_{1}(M^{2n},J)=0$, it is just Theorem 1.2.
At last, as in Kähler case [27], on almost Kähler manifolds, we raise several existence question for four different kinds of special almost Kähler metrics, working within a fixed symplectic class on a compact manifold.
1) Extremal almost Kähler metrics
2) Constant Hermitian scalar curvature almost Kähler metrics
3) Hermite-Einstein almost Kähler metrics with $c_{1}>0$
4) Hermite-Ricci solitons
2 Preliminaries
Suppose that $M$ is an almost complex manifold with an almost complex structure $J$, then for any $x\in M$, $T_{x}M\otimes_{\mathbb{R}}\mathbb{C}$ which is the complexification of $T_{x}M$, has the following decomposition
$T_{x}M\otimes_{\mathbb{R}}\mathbb{C}=T^{1,0}_{x}+T^{0,1}_{x},$ (2.1)
where $T^{1,0}_{x}$ and $T^{0,1}_{x}$ are the eigenspaces of $J$ corresponding to the eigenvalues $\sqrt{-1}$ and $-\sqrt{-1}$, respectively. A complex tangent vector is of type $(1,0)$ (resp. $(0,1)$) if it belongs to $T^{1,0}_{x}$ (resp. $T^{0,1}_{x}$). Let $TM\otimes_{\mathbb{R}}\mathbb{C}$ be the complexification of the tangent bundle. Similarly, let $T^{*}M\otimes_{\mathbb{R}}\mathbb{C}$ denote the complexification of the cotangent bundle $T^{*}M$. $J$ can act on $T^{*}M\otimes_{\mathbb{R}}\mathbb{C}$ as follows:
$\forall\alpha\in T^{*}M\otimes_{\mathbb{R}}\mathbb{C},\,\,\,J\alpha(\cdot)=-\alpha(J\cdot).$
Hence $T^{*}M\otimes_{\mathbb{R}}\mathbb{C}$ has the following decomposition according to the eigenvalues $\mp\sqrt{-1}$:
$T^{*}M\otimes_{\mathbb{R}}\mathbb{C}=\Lambda^{1,0}_{J}\oplus\Lambda^{0,1}_{J}.$ (2.2)
We can form exterior bundle $\Lambda^{p,q}_{J}=\Lambda^{p}\Lambda^{1,0}_{J}\otimes\Lambda^{q}\Lambda^{0,1}_{J}$. Let $\Omega^{p,q}_{J}(M)$ denote the space of $C^{\infty}$ sections of the bundle $\Lambda^{p,q}_{J}$. The exterior differential operator acts on $\Omega^{p,q}_{J}$ as follows:
$d\Omega^{p,q}_{J}\subset\Omega^{p-1,q+2}_{J}+\Omega^{p+1,q}_{J}+\Omega^{p,q+1}_{J}+\Omega^{p+2,q-1}_{J}.$ (2.3)
Hence, $d$ has the following decomposition:
$d=A_{J}\oplus\partial_{J}\oplus\bar{\partial}_{J}\oplus\bar{A}_{J},$ (2.4)
where each component is a derivation with bi-degrees given by
$|A_{J}|=(-1,2),\,\,|\partial_{J}|=(1,0),\,\,|\bar{\partial}_{J}|=(0,1),\,\,|\bar{A}_{J}|=(2,-1).$
By observation, the operators $\partial_{J},\bar{\partial}_{J}$ are of the first order and the operators $A_{J},\bar{A}_{J}$ are of the zero order (see [46]).
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. Let $\Omega^{2}_{\mathbb{R}}(M)$ denote the space of real smooth $2$-forms on $M^{2n}$, that is, the real $C^{\infty}$ sections of the bundle $\Lambda^{2}_{\mathbb{R}}(M)$. The almost complex structure $J$ acts on the space $\Omega_{\mathbb{R}}^{2}$ of smooth $2$-forms on $M^{2n}$ as an involution by
$\alpha\longmapsto\alpha(J\cdot,J\cdot),\quad\alpha\in\Omega_{\mathbb{R}}^{2}(M).$ (2.5)
This gives the $J$-invariant, $J$-anti-invariant decomposition of $2$-forms (see [26]):
$\Omega_{\mathbb{R}}^{2}=\Omega^{+}_{J}\oplus\Omega^{-}_{J},\quad\alpha=\alpha_{J}^{+}+\alpha_{J}^{-}$
as well as the splitting of corresponding vector bundles
${\Lambda}_{\mathbb{R}}^{2}={\Lambda}_{J}^{+}\oplus{\Lambda}_{J}^{-}.$ (2.6)
We define the following operators:
$d^{+}_{J}$ $=$ $P^{+}_{J}d:\,\,\,\Omega_{\mathbb{R}}^{1}\longrightarrow\Omega^{+}_{J},$
$d^{-}_{J}$ $=$ $P^{-}_{J}d:\,\,\,\Omega_{\mathbb{R}}^{1}\longrightarrow\Omega^{-}_{J},$ (2.7)
where $P^{\pm}_{J}:\Omega_{\mathbb{R}}^{2}\longrightarrow\Omega^{\pm}_{J}$. A differential $k$-form $B_{k}$ with $k\leq n$ is called primitive if $L^{n-k+1}_{\omega}B_{k}=0$, or equivalently, $\Lambda_{\omega}B_{k}=0$ (see [62, 70]). Here $L_{\omega}$ is the Lefschetz operator (see [8, 62, 70]) which is defined acting on a $k$-form $A_{k}\in\Omega_{\mathbb{R}}^{k}(M)$ by
$L_{\omega}(A_{k})=\omega\wedge A_{k}.$
The dual Lefschetz operator is defined by $\Lambda_{\omega}:\Omega_{\mathbb{R}}^{k}(M)\rightarrow\Omega_{\mathbb{R}}^{k-2}(M)$. Indeed, $\Lambda_{\omega}$ is a contraction map associated to the symplectic form $\omega$. Define the space of primitive $k$-forms by $\Omega^{k}_{0}(M)$. Specifically,
$\Omega^{2}_{0}(M)=\{\alpha\in\Omega_{\mathbb{R}}^{2}(M)\,|\,\omega^{n-1}\wedge\alpha=0\}.$
Then
$\Omega^{2}_{0}(M)=\Omega^{-}_{J}(M)\oplus\Omega^{+}_{J,0}(M)$
and
$\Omega_{\mathbb{R}}^{2}(M)$ $=$ $\Omega^{2}_{1}(M)\oplus\Omega^{2}_{0}(M)$ (2.8)
$=$ $\Omega^{2}_{1}(M)\oplus\Omega^{-}_{J}(M)\oplus\Omega^{+}_{J,0}(M),$
where $\Omega^{+}_{J,0}(M)$ is the space of the primitive $J$-invariant $2$-forms and
$\Omega^{2}_{1}(M):=\{f\omega\,|\,f\in C^{\infty}(M,\mathbb{R})\}.$
Let
$d^{-}_{J}d^{*}:\Omega^{-}_{J}(M)\longrightarrow\Omega^{-}_{J}(M),$
where $d^{*}=-*_{g}d*_{g}$ and $*_{g}$ is the Hodge star operator with respect to the metric $g$. For any $\alpha,\beta\in\Omega^{-}_{J}(M)$, it is easy to see that
$<d^{-}_{J}d^{*}\alpha,\beta>_{g}=<dd^{*}\alpha,\beta>_{g}=<\alpha,dd^{*}\beta>_{g}=<\alpha,d^{-}_{J}d^{*}\beta>_{g}.$
Hence, $d^{-}_{J}d^{*}$ is a self-adjoint operator. If $\alpha\in\ker(d^{-}_{J}d^{*})\subset\Omega^{-}_{J}(M)$,
$0=<d^{-}_{J}d^{*}\alpha,\alpha>_{g}=<d^{*}\alpha,d^{*}\alpha>_{g}.$
So we obtain that,
$\ker(d^{-}_{J}d^{*})={\rm coker}(d^{-}_{J}d^{*})=\{\alpha\in\Omega^{-}_{J}(M)\,\,|\,\,d^{*}\alpha=0\}.$
By Weil’s identity [62, 68],
$*_{g}\alpha=\frac{1}{(n-2)!}\omega^{n-2}\wedge\alpha.$
So
$d*_{g}\alpha=\frac{1}{(n-2)!}\omega^{n-2}\wedge d\alpha=0.$
Hence,
$\ker(d^{-}_{J}d^{*})={\rm coker}(d^{-}_{J}d^{*})$ $=$ $\{\alpha\in\Omega^{-}_{J}(M)\,\,|\,\,d^{*}\alpha=0\}$ (2.9)
$=$ $\{\alpha\in\Omega^{-}_{J}(M)\,\,|\,\,\omega^{n-2}\wedge d\alpha=0\}.$
If $n=2$, $d^{*}\alpha=0$ implies $d\alpha=0$, that is, $\alpha$ is $J$-anti-invariant harmonic $2$-form in $\Omega^{-}_{J}(M^{4})$ (cf. [39]). Hence, $\ker(d^{-}_{J}d^{*})=\mathcal{H}^{-}_{J}=\mathcal{H}^{2}\cap\Omega^{-}_{J}$. If $n\geq 3$, it is easy to see that $\mathcal{H}^{-}_{J}\subsetneq\ker(d^{-}_{J}d^{*})$.
Notice that $C^{\infty}(M)$ is dense in $L^{2}_{2}(M)$, we can extend $d^{-}_{J}d^{*}$ to become a closed, densely defined operator (see [36]),
$d^{-}_{J}d^{*}:\Lambda^{-}_{J}\otimes L^{2}_{2}(M)\longrightarrow\Lambda^{-}_{J}\otimes L^{2}(M).$
It is easy to see that, in the sense of distributions,
$\ker(d^{-}_{J}d^{*})=\{\alpha\in\Lambda^{-}_{J}\otimes L^{2}_{2}(M^{2n})\,\,|\,\,d^{-}_{J}d^{*}\alpha=0\}$
is closed. Actually, let $\{\alpha_{i}\}$ be a sequence in $\ker(d^{-}_{J}d^{*})$ such that $\alpha_{i}$ is converging in $L^{2}_{2}$ to some $\alpha\in\Lambda^{-}_{J}\otimes L^{2}_{2}(M)$. Then $\{d^{-}_{J}d^{*}\alpha_{i}\}=\{0\}$ is a constant sequence which naturally converges to 00. Hence, $d^{-}_{J}d^{*}\alpha=0$ and $\alpha\in\ker(d^{-}_{J}d^{*})$ since $d^{-}_{J}d^{*}$ is a closed operator. Let
${\rm D}(d^{-}_{J}d^{*})=\Lambda^{-}_{J}\otimes L^{2}_{2}(M)\setminus\ker(d^{-}_{J}d^{*})$
and
${\rm R}(d^{-}_{J}d^{*})=\Lambda^{-}_{J}\otimes L^{2}(M)\setminus{\rm coker}(d^{-}_{J}d^{*}).$
Then
$d^{-}_{J}d^{*}:{\rm D}(d^{-}_{J}d^{*})\longrightarrow{\rm R}(d^{-}_{J}d^{*})$
is invertible. For any
[…]
by direct calculation and Proposition 1.13.1 in [33], we will get
$Jdf=d^{*}(f\omega),\,\,d^{-}_{J}Jdf=d^{-}_{J}d^{*}(f\omega),\,\,d^{+}_{J}Jdf=2\sqrt{-1}\partial_{J}\bar{\partial}_{J}f.$
Notice that
$d^{-}_{J}Jdf=d^{-}_{J}d^{*}(f\omega)\,\bot\,{\rm coker}(d^{-}_{J}d^{*}),$
then there exists a unique $\alpha_{f}\in{\rm D}(d^{-}_{J}d^{*})$ such that
$d^{-}_{J}d^{*}(f\omega+\alpha_{f})=0.$
We have the following theorem which has been proven by Lejmi in a $4$-dimensional case [39].
Theorem 2.1.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. The operator $d^{-}_{J}d^{*}:{\rm D}(d^{-}_{J}d^{*})\longrightarrow{\rm R}(d^{-}_{J}d^{*})$ is an invertible formal self-adjoint and nonnegative operator. Hence, for any $f\in C^{\infty}(M^{2n})_{0}$, we have $d^{-}_{J}Jdf\,\bot\,{\rm coker}(d^{-}_{J}d^{*})$, and there exists a unique $\alpha_{f}\in{\rm D}(d^{-}_{J}d^{*})$ such that
$d^{-}_{J}d^{*}(\alpha_{f})=d^{-}_{J}Jdf.$
Remark 2.2.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. We have the following second order linear differential operator
$P_{J}:\Omega^{2}_{0}(M)$$\rightarrow$ $\Omega^{2}_{0}(M)$
$\psi$ $\mapsto$ $\Delta_{g}\psi-\frac{1}{n}g(\Delta_{g}\psi,\omega)\omega,$
where $\Delta_{g}$ is the Riemannian Laplacian with respect to the almost Kähler metric $g$ (here we use the convention $<\omega,\omega>_{g}=n$). $P_{J}$ is a self-adjoint strongly elliptic linear operator with kernel the primitive $g$-harmonic $2$-forms (Lejmi [39], Tan-Wang-Zhou [51]). In case $n=2$, Lejmi proved that $P_{J}$ preserves the decomposition
$\Omega^{2}_{0}=\Omega^{+}_{J,0}\oplus\Omega^{-}_{J}.$
Furthermore, $P_{J}|_{\Omega^{+}_{J,0}}(\psi)=\Delta_{g}\psi$ and $P_{J}|_{\Omega^{-}_{J}}(\psi)=2d^{-}_{J}d^{*}\psi$. He also pointed out that $P_{J}|_{\Omega^{-}_{J}}(\psi)=2d^{-}_{J}d^{*}\psi$ is a self-adjoint strongly elliptic linear operator from $\Omega^{-}_{J}$ to $\Omega^{-}_{J}$ on a closed almost Kähler $4$-manifold. In general case ($n>2$), we have
$P_{J}|_{\Omega^{-}_{J}}:\Omega^{-}_{J}\longrightarrow\Omega^{2}_{0}$
is an elliptic operator since its symbol is injective, although not invertible [61].
As done in Tan-Wang-Zhou-Zhu [53], Wang-Wang-Zhu [64], by using Theorem 2.1, we define $\mathcal{D}^{+}_{J}$ operator on higher dimensional almost Kähler manifolds (see Wang-Zhang-Zheng-Zhu [66, Remark 1.4]):
Definition 2.3.
Define $\mathcal{D}^{+}_{J}$ as follows
$\mathcal{D}^{+}_{J}:L^{2}_{2}(M)_{0}\longrightarrow\Lambda^{+}_{J}\otimes L^{2}(M),$
$\mathcal{D}^{+}_{J}(f)=dJdf+dd^{*}\sigma(f)=dd^{*}(f\omega)+dd^{*}\sigma(f),\quad\sigma(f)\in\Omega_{J}^{-}(M^{2n}),$
satisfying $d^{-}_{J}d^{*}(f\omega)+d^{-}_{J}d^{*}\sigma(f)=0$. Let
$\mathcal{W}_{J}:L^{2}_{2}(M)_{0}\longrightarrow\Lambda_{\mathbb{R}}^{1}\otimes L^{2}_{1}(M),$
$\mathcal{W}_{J}(f)=d^{*}(f\omega+\sigma(f)).$
Then $d\mathcal{W}_{J}(f)=\mathcal{D}^{+}_{J}(f)$ and $d^{-}_{J}\mathcal{W}_{J}(f)=0$. The function $f$ is called almost Kähler potential with respect to the almost Kähler metric $g$.
Remark 2.4.
If $J$ is integrable, that is $N_{J}=0$, then $\mathcal{D}^{+}_{J}(f)=2\sqrt{-1}\partial_{J}\bar{\partial}_{J}f$. So we can regard $\mathcal{D}^{+}_{J}$ as a generalized $\partial\bar{\partial}$ operator.
Denote the space of harmonic $2$-forms by $\mathcal{H}^{2}_{dR}$ (cf. [13]). Let
$\mathcal{H}^{-}_{J}=\mathcal{H}^{2}_{dR}\cap\Omega^{-}_{J},\,\,\,h^{-}_{J}=\dim\mathcal{H}^{-}_{J};$
$\mathcal{H}^{+}_{J,0}=\mathcal{H}^{2}_{dR}\cap\Omega^{+}_{J,0},\,\,\,h^{+}_{J,0}=\dim\mathcal{H}^{+}_{J,0}.$
Here $\mathcal{H}^{-}_{J}$ and $\mathcal{H}^{+}_{J,0}$ are harmonic representation groups of their respective real de Rham cohomology groups on $M^{2n}$ (see Draghici-Li-Zhang [29] for $n=2$). Similar to Theorem 4.3 for complex de Rham cohomology groups in [16] ,
$h^{-}_{J}+h^{+}_{J,0}\leq b^{2}-1.$
If $h^{-}_{J}+h^{+}_{J,0}=b^{2}-1$, then
$\mathcal{H}^{2}_{dR}=Span\{\omega\}\oplus\mathcal{H}^{+}_{J,0}\oplus\mathcal{H}^{-}_{J}.$
Notice that if $n=2$, $\mathcal{H}^{2,0}_{dR}=\{0\}=\mathcal{H}^{0,2}_{dR}$ (cf. [16, Lemma 5.6]). As done in almost Kähler $4$-manifolds (see Tan-Wang-Zhang-Zhu [50] and Tan-Wang-Zhou [52]), one can define
$\mathcal{H}^{\perp}_{J,0}=\mathbb{R}\cdot\omega\oplus\{\alpha_{f}=f\omega+d^{-}_{J}(v_{f}+\bar{v}_{f})\in\mathcal{H}^{2}_{dR}\,\,|\,\,f\in C^{\infty}(M)_{0},\,v_{f}\in\Omega^{0,1}_{J}(M)\}.$
Then $h^{\perp}_{J,0}=\dim\mathcal{H}^{\perp}_{J,0}\geq 1$ and
$\mathcal{H}^{\perp}_{J,0}\oplus\mathcal{H}^{+}_{J,0}\oplus\mathcal{H}^{-}_{J}\subseteq\mathcal{H}^{2}_{dR}.$
It is easy to see that
$\ker\mathcal{W}_{J}=Span_{\mathbb{R}}\{f_{2},\cdot\cdot\cdot,f_{h^{\perp}_{J,0}-1}\,\,|\,\,\alpha_{f_{i}}\in\mathcal{H}^{\perp}_{J,0}\}.$
If $n=2$, $\dim\mathcal{H}^{\perp}_{J,0}=b^{+}-\dim\mathcal{H}^{-}_{J}$, $10\leq\dim\mathcal{H}^{-}_{J}\leq b^{+}-$ 1 and
$\mathcal{H}^{\perp}_{J,0}\oplus\mathcal{H}^{+}_{J,0}\oplus\mathcal{H}^{-}_{J}=\mathcal{H}^{2}_{dR},$
where $b^{+}$ is the self-dual second Betti number of $M^{4}$ (cf. Tan-Wang-Zhou [52]).
If $n\geq 3$, for $\alpha_{f}\in\mathcal{H}^{\perp}_{J,0}$, $d^{*}\alpha_{f}=0$ and $d\alpha_{f}=\omega\wedge df+dd^{-}_{J}(v_{f}+\bar{v}_{f})=0$. Hence,
00 0 $=$ $d*_{g}\alpha_{f}$
$=$ $d*_{g}[f\omega+d^{-}_{J}(v_{f}+\bar{v}_{f})]$
$=$ $d[\frac{1}{(n-1)!}f\omega^{n-1}+\frac{1}{(n-2)!}\omega^{n-2}\wedge d^{-}_{J}(v_{f}+\bar{v}_{f})]$
$=$ $\frac{1}{(n-1)!}\omega^{n-1}\wedge df+\frac{1}{(n-2)!}\omega^{n-2}\wedge dd^{-}_{J}(v_{f}+\bar{v}_{f})$
$=$ $\frac{1}{(n-1)!}\omega^{n-1}\wedge df-\frac{1}{(n-2)!}\omega^{n-2}\wedge\omega\wedge df$
$=$ $(\frac{1}{(n-1)!}-\frac{1}{(n-2)!})\omega^{n-1}\wedge df.$
By Corollary $2.7$ in Yan [70], $L_{\omega}^{n-1}:\Omega_{\mathbb{R}}^{1}\rightarrow\Omega_{\mathbb{R}}^{2n-1}$ is an isomorphism. So we obtain $df=0$ and $f=0$. Therefore, $d^{-}_{J}(v_{f}+\bar{v}_{f})\in\mathcal{H}^{2}_{dR}$. Moreover, $d^{-}_{J}(v_{f}+\bar{v}_{f})\in\mathcal{H}^{-}_{J}$. Hence, $d^{-}_{J}(v_{f}+\bar{v}_{f})=0$ and $\mathcal{H}^{\perp}_{J,0}=\mathbb{R}\cdot\omega$. Thus, $\ker\mathcal{W}_{J}=\{0\}$.
Notice that $d_{J}^{-}\mathcal{W}_{J}(f)=0$, as $\bar{\partial}$-problem in classical complex analysis [36], $(\mathcal{W}_{J},d_{J}^{-})$-problem is whether $\mathcal{W}_{J}(f)=A$ has a solution for any $A$ satisfying $d_{J}^{-}A=0$. If we can use the theory of Hilbert space, considering
$L_{2}^{2}(M)_{0}\stackrel{{\scriptstyle\mathcal{W}_{J}}}{{\rightarrow}}\wedge_{\mathbb{R}}^{1}\otimes L_{1}^{2}(M)\stackrel{{\scriptstyle d_{J}^{-}}}{{\rightarrow}}\wedge_{J}^{-}\otimes L^{2}(M),$ (2.10)
then the above problem is equivalent to whether the kernel of $d_{J}^{-}$ is equal to the image of $\mathcal{W}_{J}$. Recall the definition of $\mathcal{W}_{J}$: $f\in L^{2}_{2}(M)_{0}$, $\sigma(f)\in{\rm D}(d^{-}_{J}d^{*})$ such that
$\mathcal{W}_{J}(f)=d^{*}(f\omega+\sigma(f))$
satisfying $d^{-}_{J}\mathcal{W}_{J}(f)=0$, $d^{*}\mathcal{W}_{J}(f)=0$ and
$d\mathcal{W}_{J}(f)=d^{+}_{J}\mathcal{W}_{J}(f)\in\Lambda^{+}_{J}\otimes L^{2}(M).$
Without loss of generality, we may assume that if $A\in\Omega^{1}_{\mathbb{R}}(M)$, $d^{*}A=0$ and $d^{-}_{J}A=0$, then
$<\mathcal{W}_{J}(f),A>_{g}$ $=$ $\int_{M}A\wedge*_{g}d^{*}(f\omega+\sigma(f))$
$=$ $-\int_{M}A\wedge d(f\frac{\omega^{n-1}}{(n-1)!}+\sigma(f)\wedge\frac{\omega^{n-2}}{(n-2)!})$
[…]
$=$ $-\int_{M}d^{+}_{J}A\wedge f\frac{\omega^{n-1}}{(n-1)!}$
$=$ $<f,\mathcal{W}_{J}^{*}A>_{g}.$
Thus, the formal $L^{2}$-adjoint operator of $\mathcal{W}_{J}$ is
$\mathcal{W}_{J}^{*}A=\frac{-n\omega^{n-1}\wedge d^{+}_{J}A}{\omega^{n}}=-(\Lambda_{\omega}d^{+}_{J}A).$ (2.11)
As done on closed almost Kähler $4$-manifolds, by using $L^{2}$-method [36], calculating the $L^{2}$-norm of $\mathcal{W}_{J}^{*}A$ and applying Riesz Representation Theorem [74], it is easy to get the following theorem (cf. [53, Appendix A.3] or [64, Section 6]):
Theorem 2.5.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. One can solve $(\mathcal{W}_{J},d^{-}_{J})$-problem.
Suppose $\psi\in\Lambda^{1,1}_{\mathbb{R}}\otimes L^{2}(M)$ is $d$-exact, hence
$\psi=d(u_{\psi}+\bar{u}_{\psi})=d^{+}_{J}(u_{\psi}+\bar{u}_{\psi}),\,\,\,i.e.,\,\,\,d^{-}_{J}(u_{\psi}+\bar{u}_{\psi})=0,$ (2.12)
for some $u_{\psi}\in\Lambda^{0,1}_{J}\otimes L^{2}_{1}(M)$. By Theorem 2.5, there exists $f\in L_{2}^{2}(M)_{0}$ such that $\mathcal{W}_{J}(f)=u_{\psi}+\bar{u}_{\psi}$. Hence,
$\psi=d(u_{\psi}+\bar{u}_{\psi})=d\mathcal{W}_{J}(f)=\mathcal{D}^{+}_{J}(f).$
We summarize the above discussion to obtain the following corollary:
Corollary 2.6.
Suppose that $\psi\in\Lambda^{+}_{J}\otimes L^{2}(M)$ is $d$-exact, that is, there is $u_{\psi}\in\Lambda^{0,1}_{J}\otimes L^{2}_{1}(M)$ such that $\psi=d(u_{\psi}+\bar{u}_{\psi})$. Then $\psi$ is $\mathcal{D}^{+}_{J}$-exact, that is, there exists $f_{\psi}\in L^{2}_{2}(M)_{0}$ such that $\psi=\mathcal{D}^{+}_{J}(f_{\psi})$.
With the above lemma, we can easily get the following proposition (cf. Tan-Wang-Zhou-Zhu [53]):
Proposition 2.7.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. $\mathcal{D}^{+}_{J}:L^{2}_{2}(M)_{0}\longrightarrow\Lambda^{+}_{J}\otimes L^{2}(M)$ has closed range.
Proof.
Let $\{f_{i}\}$ be a sequence of real functions on $M^{2n}$ in $L^{2}_{2}(M)_{0}$. By Definition 2.3, $\{\mathcal{W}_{J}(f_{i})\}$ is a sequence of real $1$-forms on $M^{2n}$ with coefficients in $L^{2}_{1}$ such that
$\psi_{i}=d\mathcal{W}_{J}(f_{i})=\mathcal{D}^{+}_{J}(f_{i})\in\Lambda^{+}_{J}\otimes L^{2}(M)$
is converging in $L^{2}$ to some $\psi\in\Lambda^{+}_{J}\otimes L^{2}(M)$. It is clear that $d^{*}\mathcal{W}_{J}(f_{i})=0$ and $\mathcal{W}_{J}(f_{i})$ is perpendicular to the harmonic $1$-forms. It is well known that $d+d^{*}$ is an elliptic system (cf. Donaldson-Kronheimer [28]). Hence there exists a constant $C$ such that
$\|\mathcal{W}_{J}(f_{i})\|^{2}_{L^{2}_{1}(M)}$$\leq$ $C(\|d\mathcal{W}_{J}(f_{i})\|^{2}_{L^{2}(M)}+\|d^{*}\mathcal{W}_{J}(f_{i})\|^{2}_{L^{2}(M)})$
$=$ $C\|d\mathcal{W}_{J}(f_{i})\|^{2}_{L^{2}(M)}<Const.$
Hence, $\{\mathcal{W}_{J}(f_{i})\}$ is bounded in $L^{2}_{1}$, so a subsequence of $\{\mathcal{W}_{J}(f_{i})\}$ converges weakly in $L^{2}_{1}$ to some $\mathcal{W}_{J}\in\Lambda^{1}_{\mathbb{R}}\otimes L^{2}_{1}(M)$. Since $d\mathcal{W}_{J}(f_{i})\in\Lambda^{+}_{J}\otimes L^{2}(M)$, it follows that
$d\mathcal{W}_{J}=\psi\in\Lambda^{+}_{J}\otimes L^{2}(M).$
By Lemma 2.6, there exists $f\in L^{2}_{2}(M)_{0}$ such that $\mathcal{D}^{+}_{J}(f)=d\mathcal{W}_{J}(f)=\psi$. This completes the proof of Proposition 2.7. ∎
3 Generalized Monge-Ampère Equation on Almost Kähler Manifolds
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. Denote by
[…]
Define
$\mathcal{D}^{+}_{J}:C^{\infty}(M,J)_{0}\longrightarrow\Omega^{+}_{J}(M),$
by
$\mathcal{D}^{+}_{J}(f)=d\mathcal{W}_{J}(f)=dJdf+dd^{*}\sigma(f)$
satisfying $d^{-}_{J}\mathcal{W}_{J}(f)=0$, where $\mathcal{W}_{J}(f)=Jdf+d^{*}\sigma(f)$ and $\sigma(f)\in\Omega^{-}_{J}(M)$. It is easy to see that $\mathcal{D}^{+}_{J}$ is a linear operator and $d^{*}\mathcal{W}_{J}(f)=0$ with the fact that $Jdf=d^{*}(f\omega)$.
It is well known that Monge-Ampère equations are a class of important fully nonlinear equations profoundly related to many fields of analysis and geometry (see [30, 71]). The solution of the Monge-Ampère equation has been studied extensively. There are many existence, uniqueness and regularity results of the Monge-Ampère equation under different conditions and we refer the reader to Tosatti-Weinkove [59], Chu-Tosatti-Weinkove [15], Demailly-Pali [19], Pliś [47], Zhang-Zhang [75] and others. As done in almost Kähler surface in Wang-Zhang-Zheng-Zhu [66], we can define a generalized Monge-Ampère equation on $(M^{2n},\omega,J,g)$,
$(\omega+\mathcal{D}^{+}_{J}(f))^{n}=e^{F}\omega^{n}$ (3.1)
for some real function $f\in C^{\infty}(M,J)_{0}$ with
$\omega+\mathcal{D}^{+}_{J}(f)>0,$
where $F\in C^{\infty}(M)$ satisfies
$\int_{M^{2n}}\omega^{n}=\int_{M^{2n}}e^{F}\omega^{n}.$ (3.2)
If $J$ is integrable, $\mathcal{D}^{+}_{J}(f)=2\sqrt{-1}\partial_{J}\bar{\partial}_{J}f$. Then Equation (3.1) evolves into the famous Monge-Ampère equation (cf. [9, 72]), which is called the generalized Monge-Ampère equation on almost Kähler $2n$-manifolds. If $n=2$, see Wang-Zhang-Zheng-Zhu [66].
By Definition 2.3, the generalized Monge-Ampère equation (3.1) is equivalent to the following Calabi-Yau equation for $1$-forms:
$(\omega+da)^{n}=e^{F}\omega^{n},$ (3.3)
where $a\in\Omega_{\mathbb{R}}^{1}(M)$, $d^{*}a=0$ and $d^{-}_{J}a=0$. In fact, we can choose $a=\mathcal{W}_{J}(f)$, $f\in C^{\infty}(M,J)_{0}$. For form-type Calabi-Yau equations, see Weinkove [69], Tosatti-Weinkove-Yau [61], Fu-Wang-Wu [30], Delanoë [20], Wang-Zhu [65] and others.
We want to consider the local theory of the Calabi-Yau equation on almost Kähler $2n$-manifolds (cf. [20, 65]).
Definition 3.1.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. The sets $A$, $B$, $A_{+}$ and $B_{+}$ are defined as follows:
$A:=\{u\in\Omega^{1}_{\mathbb{R}}(M)\,\,|\,\,d^{-}_{J}u=0,\,d^{*}u=0\};$
$A_{+}:=\{u\in A\,\,|\,\,\omega(u)=\omega+du>0\};$
$B:=\{f\in C^{\infty}(M)\,\,|\,\,\int_{M^{2n}}f\omega^{n}=\int_{M^{2n}}\omega^{n}\};$
$B_{+}:=\{f\in B\,\,|\,\,f>0\,\,on\,\,M^{2n}\}.$
We define an operator $\mathcal{F}$ from $A$ to $B$ as follows:
$u\mapsto\mathcal{F}(u),$
where
$\mathcal{F}(u)\omega^{n}=(\omega(u))^{n},\,\,\,\omega(u)=\omega+du.$ (3.4)
Restricting the operator $\mathcal{F}$ to $A_{+}$, we get $\mathcal{F}(A_{+})\subset B_{+}$. Thus, the existence of a solution to Equation (3.1) is equivalent to that the restricted operator
$\mathcal{F}|_{A_{+}}:A_{+}\rightarrow B_{+}$ (3.5)
is surjective. Moerover, we have the following result (cf. [20, Proposition 5] or [65, Proposition 2.4]):
Proposition 3.2.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$.
(1) If $u\in A_{+}$, then $\mathcal{F}(u)>0$;
(2) define
$B_{\varepsilon}(0):=\{u\in A\,|\,\|u\|_{C^{1}}\leq\varepsilon\},$
where $C^{1}$ is $C^{1}$-norm introduced by the metric $g$; if $\varepsilon<<1$, then $B_{\varepsilon}(0)\subset A_{+}$.
Suppose that $u_{0}\in A_{+}$. By Proposition 3.2, there exists a small neighborhood $U(u_{0})$ in $A$ such that $U(u_{0})\subset A_{+}$, and if $u_{1}\in A$, $\varepsilon<<1$, we can obtain that $\mathcal{F}(u_{0}+\varepsilon u_{1})>0$.
By solvability of $(\mathcal{W}_{J},d^{-}_{J})$-problem (see Theorem 2.5), for any $\phi\in A$, there exists $f_{\phi}\in C^{\infty}(M,J)_{0}$ such that $\mathcal{W}_{J}(f_{\phi})=\phi$. For any $u\in A_{+}$, since $A_{+}$ is an open subset of $A$, it is easy to see that the tangent space at $u$, $T_{u}A_{+}$, is $A$. For $\phi\in T_{u}A_{+}$, define $L(u)(\phi)$ by
$L(u)(\phi)=\frac{d}{dt}\mathcal{F}(u+t\phi)|_{t=0}.$
By simple calculation, we can obtain
$L(u)(\phi)$ $=$ $\frac{n\omega(u)^{n-1}\wedge\mathcal{D}^{+}_{J}(f_{\phi})}{\omega^{n}}.$
It is easy to see that $L(u)$ is a linear elliptic system on $A$ (cf. [20, 65]). Moreover, we have $\ker L(u)=\{0\}$. Indeed, for $u\in A_{+}$,
$\omega(u)=\omega+du=\omega+\mathcal{D}^{+}_{J}(f_{u})>0,$
that is, a symplectic form compatible with $J$ on $M^{2n}$, where $\mathcal{W}_{J}(f_{u})=u$. Let $g_{u}(\cdot,\cdot)=\omega(u)(\cdot,J\cdot)$. Then $(\omega(u),J,g_{u})$ is an almost Kähler structure on $M^{2n}$. By the primitive decomposition,
$\Omega^{+}_{J}(M)=\Omega^{2}_{1}(u)(M)\oplus\Omega^{+}_{J,0}(u)(M),$
where
$\Omega^{2}_{1}(u):=\{f\omega(u)\,|\,f\in C^{\infty}(M)\},$
and
$\Omega^{+}_{J,0}(u)(M)=\{\alpha\in\Omega^{+}_{J}(M)\,|\,\omega(u)^{n-1}\wedge\alpha=0\}.$
If $L(u)(\phi)=0$, then $\mathcal{D}^{+}_{J}(f_{\phi})$ us a primitive 2-form on $M$. By Weil’s identity [62],
$*_{g_{u}}\mathcal{D}^{+}_{J}(f_{\phi})=-\mathcal{D}^{+}_{J}(f_{\phi})\wedge\frac{\omega(u)^{n-2}}{(n-2)!}.$
Thus,
$<\mathcal{D}^{+}_{J}(f_{\phi}),\mathcal{D}^{+}_{J}(f_{\phi})>_{g_{u}}$ $=$ $\int_{M^{2n}}\mathcal{D}^{+}_{J}(f_{\phi})\wedge*_{g_{u}}\mathcal{D}^{+}_{J}(f_{\phi})$
$=$ $-\int_{M^{2n}}\mathcal{D}^{+}_{J}(f_{\phi})\wedge\mathcal{D}^{+}_{J}(f_{\phi})\wedge\frac{\omega(u)^{n-2}}{(n-2)!}$
$=$ $0$
Therefore, $\mathcal{D}^{+}_{J}(f_{\phi})=d\mathcal{W}_{J}(f_{\phi})=0$. On the other hand, we have $d^{*}\mathcal{W}_{J}(f_{\phi})=0$. So $\mathcal{W}_{J}(f_{\phi})$ is a harmonic $1$-form [13]. By the definition of $\mathcal{W}_{J}(f_{\phi})$, we know that $\mathcal{W}_{J}(f_{\phi})$ is a $d^{*}$-exact form. Therefore, by Hodge decomposition, $\mathcal{W}_{J}(f_{\phi})=0$, that is, $\phi=0$. Hence, we have the following lemma (cf. [20, Proposition 1] or [65, Lemma 2.5]):
Lemma 3.3.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. Then the restricted operator
$\mathcal{F}:A_{+}\longrightarrow B_{+}$
is elliptic type on $A_{+}$. Moreover, the tangent map, $d\mathcal{F}(u)=L(u)$, of $\mathcal{F}$ at $u\in A_{+}$ is a linear elliptic system on $A$ and $\ker L(u)=\{0\}$.
Obviously, $A_{+}\subset A$ is a convex open set. Suppose that
$\mathcal{F}(u_{0})=\mathcal{F}(u_{1})$
for $u_{0},u_{1}\in A_{+}$. Let $u_{t}=tu_{1}+(1-t)u_{0}$, $t\in[0,1]$. So
[…]
Then
[…]
Therefore $\mathcal{F}:A_{+}\rightarrow B_{+}$ is an injectivity map, and by nonlinear analysis (cf. Aubin [5]), we obtain the following result (cf. [20, Theorem 2] or [65, Proposition 2.6]):
Theorem 3.4.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. Then the restricted operator
$\mathcal{F}:A_{+}\rightarrow\mathcal{F}(A_{+})\subset B_{+}$
is a diffeomorphic map.
Let $F\in C^{\infty}(M^{2n},\mathbb{R})$ satisfying $e^{F}\in\mathcal{F}(A_{+})\subset B_{+}$, and
$\int_{M}\omega^{n}=\int_{M}e^{F}\omega^{n}.$
By the above theorem, there exists $u\in A_{+}$ such that $\mathcal{F}(u)=e^{F}$ and $\mathcal{F}(u)\omega^{n}=\omega(u)^{n}$. Then
$e^{F}\omega^{n}=\omega(u)^{n}=(\omega+du)^{n}=(\omega+d\mathcal{W}_{J}(f_{u}))^{n}=(\omega+\mathcal{D}^{+}_{J}(f_{u}))^{n}.$
Hence, with Theorem 3.4, we have the following local existence theorem for solution of generalized Monge-Ampère equation on the closed almost Kähler manifold $(M^{2n},\omega,J,g)$.
Theorem 3.5.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold. Let $F\in C^{\infty}(M^{2n},\mathbb{R})$ satisfying
$\int_{M^{2n}}e^{F}\omega^{n}=\int_{M^{2n}}\omega^{n},$
and
$e^{F}\in\mathcal{F}(A_{+})\subset B_{+}.$
Then, there exists a smooth function $f\in C^{\infty}(M,J)_{0}$ satisfying
$(\omega+\mathcal{D}^{+}_{J}(f))^{n}=e^{F}\omega^{n}.$
The remainder of this section is devoted to studying a uniqueness theorem for the solution of generalized Monge-Ampère equation on closed almost Kähler manifold $(M^{2n},\omega,J,g)$. If there are two solutions $f_{0},f_{1}$ for equation (3.1), then
$(\omega+\mathcal{D}^{+}_{J}(f_{1}))^{n}=(\omega+\mathcal{D}^{+}_{J}(f_{2}))^{n}=e^{F}\omega^{n}.$
Let $f_{t}=tf_{1}+(1-t)f_{0}$, $t\in[0,1]$.
$\mathcal{D}^{+}_{J}(f_{t})$ $=$ $d\mathcal{W}_{J}(f_{t})$
$=$ $td\mathcal{W}_{J}(f_{1})+(1-t)d\mathcal{W}_{J}(f_{0})$
$=$ $td\phi_{1}+(1-t)d\phi_{0}$
$=$ $d\phi_{t},$
where $\phi_{0}=\mathcal{W}_{J}(f_{0})$, $\phi_{1}=\mathcal{W}_{J}(f_{1})$ and $\phi_{t}=t\phi_{1}+(1-t)\phi_{0}$ are all in $A_{+}$. So by the definition of operator $L$, we have
[…]
$=$ $L(\phi_{0})(\phi_{1}-\phi_{0})\omega^{n}.$
By Lemma 3.3, we know that $\ker L(u)=\{0\}$ for $u\in A_{+}$. Hence, $\phi_{1}=\phi_{0}$ and $\mathcal{W}_{J}(f_{1})=\mathcal{W}_{J}(f_{0})$. So we obtain a uniqueness theorem for the generalized Monge-Ampère equation up to $\ker\mathcal{W}_{J}$ (cf. Calabi [9], Wang-Zhang-Zheng-Zhu [66] and Weinkove [69]).
Theorem 3.6.
The generalized Monge-Ampère equation (3.1) on almost Kähler $2n$-manifold has at most one solution up to $\ker\mathcal{W}_{J}$.
In the next section, we will consider the existence theorem for the generalized Monge-Ampère equation.
4 Existence Theorem for the Generalized Monge-Ampère Equation
This section is devoted to considering existence theorem for the generalized Monge-Ampère equation on almost Kähler manifolds (cf. [15, 61, 66]).
Let $(M^{2n},\omega,J,g)$ be a closed almost Kähler manifold and $\omega_{1}$ another $J$-compatible almost Kähler form on $M^{2n}$ satisfying $\omega_{1}=\omega+\mathcal{D}^{+}_{J}(f)\in[\omega]$ and $\omega_{1}^{n}=e^{F}\omega^{n}$. By the definition of $\mathcal{D}_{J}^{+}$, for any $f\in C^{\infty}(M)_{0}$,
$\mathcal{D}_{J}^{+}(f)=dJdf+dd^{*}\sigma(f),$
where $d^{*}=-*_{g}d*_{g}$, $\sigma(f)\in\Omega^{-}_{J}(M)$ and
$d_{J}^{-}dJdf+d_{J}^{-}d^{*}\sigma(f)=0.$
Define a smooth function $f_{0}\in C^{\infty}(M,J)_{0}$ as follows:
$-\frac{1}{2n}\Delta_{g}f_{0}=\frac{\omega^{n-1}\wedge(\omega_{1}-\omega)}{\omega^{n}},$ (4.1)
where $\Delta_{g}$ is the Laplacian of the Levi-Civita connection with respect to the almost Kähler metric $g$. In general, $f_{0}\neq f$. Using the result of Tosatti-Weinkove-Yau [61, Lemma 2.5], we can easily obtain
$-\omega^{n-1}\wedge dJdf_{0}=\frac{1}{2n}\Delta_{g}f_{0}\cdot\omega^{n}.$
Then, by Proposition A.1, we can find $a_{0}(f)\in\Omega^{1}(M)$ such that
$\mathcal{D}^{+}_{J}(f)=dJdf_{0}+da_{0}(f),$
and
$\left\{\begin{array}[]{ll}d^{*}a_{0}(f)=0&,\\ &\\ d^{-}_{J}Jdf_{0}+d^{-}_{J}a_{0}(f)=0&,\\ &\\ \omega^{n-1}\wedge da_{0}(f)=0&,\end{array}\right.$ (4.2)
where $d^{*}=-*_{g}d*_{g}$ and $*_{g}$ is the Hodge star operator with respect to the metric $g$. In general, $a_{0}(f)\neq d^{*}\sigma(f)$. Hence, we have
$\omega_{1}-\omega$ $=$ $\mathcal{D}^{+}_{J}(f)$
$=$ $dJdf+dd^{*}\sigma(f)$
$=$ $dJdf_{0}+da_{0}(f)$
$=$ $dd^{*}(f_{0}\omega)+da_{0}(f)$
$=$ $\mathcal{D}^{+}_{J,0}(f_{0}).$
Therefore, $\mathcal{W}_{J}(f)=d^{*}(f\omega)+d^{*}\sigma(f)$ can be rewritten as
$\mathcal{W}_{J}(f)=d^{*}(f_{0}\omega)+a_{0}(f).$
Then
$\left\{\begin{array}[]{ll}d\mathcal{W}_{J}(f)=\mathcal{D}^{+}_{J}(f)=\omega_{1}-\omega&,\\ &\\ d^{*}\mathcal{W}_{J}(f)=0&.\end{array}\right.$ (4.3)
Let $g_{1}=\omega_{1}(\cdot,J\cdot)$ be an almost Kähler metric on $M^{2n}$ with respect to the almost Kähler form $\omega_{1}$. Similarly, define a real function $f_{1}\in C^{\infty}(M,J)_{0}$ as follows:
$-\frac{1}{2n}\Delta_{g_{1}}f_{1}=\frac{\omega_{1}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{1}^{n}},$ (4.4)
where $\Delta_{g_{1}}$ is the Laplacian of the Levi-Civita connection with respect to the almost Kähler metric $g_{1}$ ([61]). Similarly, there exists $a_{1}(f)\in\Omega^{1}(M)$ such that
$\mathcal{D}^{+}_{J}(f)$ $=$ $dJdf_{1}+da_{1}(f)$
$=$ $\mathcal{D}^{+}_{J,1}(f_{1}),$
and
$\left\{\begin{array}[]{ll}d^{*_{1}}a_{1}(f)=0&,\\ &\\ d^{-}_{J}Jdf_{1}+d^{-}_{J}a_{1}(f)=0&,\\ &\\ \omega_{1}^{n-1}\wedge da_{1}(f)=0&,\end{array}\right.$ (4.5)
where $d^{*_{1}}=-*_{g_{1}}d*_{g_{1}}$ and $*_{g_{1}}$ is the Hodge star operator with respect to the metric $g_{1}$.
By elementary linear algebra, using simultaneous diagonalization (see McDuff-Salamon [45]), for any $p\in M^{2n}$, it is possible to find complex coordinates $z_{1},\cdot\cdot\cdot,z_{n}$ on $M^{2n}$ near $p$ such that:
Lemma 4.1.
$g(p)=2(|dz_{1}|^{2}+\cdot\cdot\cdot+|dz_{n}|^{2})$;
$g_{1}(p)=2(a_{1}|dz_{1}|^{2}+\cdot\cdot\cdot+a_{n}|dz_{n}|^{2})$;
$\omega(p)=\sqrt{-1}(dz_{1}\wedge d\bar{z}_{1}+\cdot\cdot\cdot+dz_{n}\wedge d\bar{z}_{n})$;
$\omega_{1}(p)=\sqrt{-1}(a_{1}dz_{1}\wedge d\bar{z}_{1}+\cdot\cdot\cdot+a_{n}dz_{n}\wedge d\bar{z}_{n})$, where $0<a_{1}\leq\cdot\cdot\cdot\leq a_{n}$.
By using orthonormal coordinate system [13] and the result of Tosatti-Weinkove-Yau [61, Lemma 2.5], we have
$\Delta^{1}_{g}f(p)=\sqrt{-1}\sum^{n}_{j=1}(dJdf)^{(1,1)}(p)(v_{j},\bar{v}_{j}),$ (4.6)
where $v_{j}=\partial/\partial z_{j}$ at $p$ and $\Delta^{1}_{g}$ is the complex Laplacian of the second canonical connection with respect to the almost Kähler metric $g$ at $p$. Also by Tosatti-Weinkove-Yau [61, Lemma 2.6], since $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$, then
$\Delta_{g}f=\Delta^{1}_{g}f,$ (4.7)
where $\Delta_{g}$ is the Laplacian of the Levi-Civita connection with respect to the almost Kähler metric $g$. Thus, we can relate $a_{1},\cdot\cdot\cdot,a_{n}$ to $e^{F}$ and $\Delta_{g}f=\Delta^{1}_{g}f$. If $\omega_{1}^{n}=e^{F}\omega^{n}$ (that is Equation (3.1)), we have the following lemma:
Lemma 4.2.
For any $p\in M^{2n}$, find complex coordinates $z_{1},\cdot\cdot\cdot,z_{n}$ near $p$ such that:
$\Pi^{n}_{j=1}a_{j}=e^{F(p)},\,\,\,\frac{\partial^{2}f_{1}}{\partial z_{j}\partial\bar{z}_{j}}(p)=a_{j}-1,\,\,\,\frac{\partial^{2}f_{1}}{\partial z_{i}\partial\bar{z}_{j}}(p)=0\,\,\,for\,\,\,i\neq j$
and
$\Delta_{g_{1}}f_{1}(p)=n-\sum^{n}_{j=1}a_{j}.$
Remark 4.3.
The function $f_{1}$ is similar to the real function $\varphi$ defined in Tosatti-Weinkove-Yau [61] which is called almost Kähler potential with respect to the almost Kähler form $\omega$.
By the previous lemmas, it is easy to obtain the following lemma:
Lemma 4.4.
At $p\in M^{2n}$, we have
$|d\mathcal{W}_{J}(f)(p)|^{2}_{g}=|\mathcal{D}^{+}_{J}(f)(p)|^{2}_{g}=2\sum^{n}_{j=1}(a_{j}-1)^{2},\,\,\,|g_{1}(p)|^{2}_{g}=2\sum^{n}_{j=1}a_{j}^{2}$
and
$|g_{1}^{-1}(p)|^{2}_{g}=2\sum^{n}_{j=1}a_{j}^{-2}.$
By the first and last equations in Lemma 4.2 and the fact that the geometric mean $(a_{1}\cdot\cdot\cdot a_{n})^{\frac{1}{n}}$ is less than or equal to the arithmetic mean $\frac{1}{n}(a_{1}+\cdot\cdot\cdot+a_{n})$, it is easy to obtain the following inequality:
$\Delta_{g_{1}}f_{1}(p)\leq n-ne^{\frac{F(p)}{n}}<n.$ (4.8)
As $\Delta_{g_{1}}f_{1}(p)=n-\sum^{n}_{j=1}a_{j}$ by Lemma 4.2 and $a_{j}$ are positive, from Lemma 4.4 one can show that at $p$,
$|d\mathcal{W}_{J}(f)|^{2}_{g}=|\partial_{J}\bar{\partial}_{J}f_{1}|^{2}_{g}\leq 2n+2(n-\Delta_{g_{1}}f_{1})^{2}$ (4.9)
and
$2ne^{\frac{2F}{n}}\leq|g_{1}|^{2}_{g}\leq 2(n-\Delta_{g_{1}}f_{1})^{2}.$ (4.10)
Since $\Pi^{n}_{j=1}a_{j}=e^{F(p)}$ by Lemma 4.2, we see that
$a^{-1}_{j}=e^{-F(p)}\Pi^{n}_{1\leq k\leq n,j\neq k}a_{k}.$
Notice that the inequality
$a_{k}\leq n-(\Delta_{g_{1}}f_{1})(p)$
derived from the third equation in Lemma 4.2, we find that
$a^{-2}_{j}\leq e^{-2F(p)}(n-(\Delta_{g_{1}}f_{1})(p))^{2n-2}.$ (4.11)
Hence, we follow the argument given in Chu-Tosatti-Weinkove [15, Proposition 3.1] to give a zero order estimate of the almost Kähler potential $f_{1}$.
Proposition 4.5.
Let $f_{1}$ solve the equivalent equation of the generalized Monge-Ampère equation (3.1), that is
$(\omega+dJdf_{1}+da_{1}(f))^{n}=e^{F}\omega^{n},$
and satisfies (4.4). Then there exists a constant $C$ depending only on $M^{2n},\omega,J$ and bounds of $F$, such that
$|f_{1}|\leq C,$
that is,
$\|f_{1}\|_{C^{0}(g)}\leq C.$
Proof.
Let $c=\sup_{M}f_{1}$ and define $f_{1}^{\prime}=f_{1}-c$. Hence, $\sup_{M}f_{1}^{\prime}=0$, and $f_{1}^{\prime}$ solves the equivalent equation of the generalized Monge-Ampère equation (3.1). Then, we want to prove $|f_{1}^{\prime}|\leq C$ which is equivalent to $|f_{1}|\leq C$. The following proof is completely in accordance with the proof of Proposition $3.1$ in [15]. Readers familiar with this proof can ignore this paragraph.
Since $\sup_{M}f_{1}^{\prime}=0$, it suffices to derive a uniform lower bound for $I:=\inf_{M}f_{1}^{\prime}$. Let $p\in M^{2n}$ be a point where this infimum is achieved, and choose local coordinates $\{x_{1},\cdot\cdot\cdot,x_{2n}\}$ centered at $p$ defined on an open set containing the unit ball $B\subset\mathbb{R}^{2n}$ in its interior. On $B$ let
$\varphi(x)=f_{1}^{\prime}(x)+\varepsilon\sum^{2n}_{i=1}(x^{i})^{2},$ (4.12)
where $\varepsilon>0$ will be chosen later. Then
$\varphi(0)=I=\inf_{B}\varphi,\,\,\,\inf_{\partial B}\varphi\geq\varphi(0)+\varepsilon.$ (4.13)
We define
$P=\{x\in B\,\,|\,\,|D\varphi(x)|<\varepsilon/2,\,\,{\rm and}\,\,\varphi(y)\geq\varphi(x)+D\varphi(x)\cdot(y-x),\,\,\forall y\in B\}.$
Note that $0\in P$, and that $D^{2}\varphi(x)\geq 0$ as well as $|Df_{1}^{\prime}(x)|\leq 5\varepsilon/2$ for all $x\in P$, where $D^{2}\varphi(x)$ denote the second derivative matrix at $x\in M^{2n}$. Now at any $x\in B$ the symmetric bilinear form
$H(\varphi)(x)(X,Y)$ ::=:= $(dJd\varphi)^{1,1}(\varphi)(x)(X,JY)$
$=$ $(2\sqrt{-1}\partial_{J}\bar{\partial}_{J}\varphi)(x)(X,JY)$
is equal to
$H(\varphi)=(D^{2}\varphi)^{J}+E(\varphi),$ (4.14)
where $(D^{2}\varphi)^{J}$ is the $J$-invariant part of $D^{2}\varphi$,
$(D^{2}\varphi)^{J}(x):=\frac{1}{2}(D^{2}\varphi+J^{T}\cdot D^{2}\varphi\cdot J)(x),$
and $E(\varphi)$ is an error matrix which depends linearly on $D\varphi$ (more details, see Harvey-Lawson [35] ). Therefore, using the fact that $\det(A+B)\geq\det A+\det B$ for symmetric non-negative definite matrices $A,B$, we have for all $x\in P$,
$\det(D^{2}\varphi)(x)\leq 2^{2n-1}\det((D^{2}\varphi)^{J})(x)=2^{2n-1}\det(H(\varphi)-E(\varphi))(x).$ (4.15)
Moreover, by (4.12), we have $D^{2}f_{1}^{\prime}(x)=D^{2}\varphi(x)-2\varepsilon{\rm Id}$. Therefore,
$(D^{2}f_{1}^{\prime})^{J}(x)$ $=$ $\frac{1}{2}(D^{2}f_{1}^{\prime}+J^{T}\cdot D^{2}f_{1}^{\prime}\cdot J)(x)$
$=$ $\frac{1}{2}(D^{2}\varphi-2\varepsilon{\rm Id}+J^{T}\cdot D^{2}\varphi\cdot J-2\varepsilon{\rm Id})(x)$
$=$ $D^{2}\varphi^{J}(x)-2\varepsilon{\rm Id}.$
Note that $D^{2}\varphi^{J}(x)\geq 0$. Then
$(D^{2}f_{1}^{\prime})^{J}(x)=(D^{2}\varphi)^{J}(x)-2\varepsilon{\rm Id}\geq-2\varepsilon{\rm Id},$
and using $|Df_{1}^{\prime}(x)|\leq 5\varepsilon/2$ together with (4.14) we obtain
$H(f_{1}^{\prime})(x)\geq-C_{1}\varepsilon{\rm Id}$
for a uniform constant $C_{1}$. Therefore, by (4.9), at any $x\in P$ we have
$\omega+\mathcal{D}^{+}_{J}(\varphi)\geq\frac{1}{2}\omega$ (4.16)
provided $\varepsilon$ is sufficiently small (this fixes the value of $\varepsilon$). Here, by the definition of $\mathcal{D}^{+}_{J}$ and Proposition A.1, we have
$\mathcal{D}^{+}_{J}(\varphi)$ $=$ $dJd\varphi+dd^{*}\sigma(\varphi)$
$=$ $dJd\varphi_{1}^{\prime}+da_{1}(\varphi)$
and
$\left\{\begin{array}[]{ll}d^{*}a_{1}(\varphi)=0&,\\ &\\ d^{-}_{J}Jd\varphi_{1}^{\prime}+d^{-}_{J}a_{1}(\varphi)=0&,\\ &\\ \omega^{n-1}\wedge da_{1}(\varphi)=0&.\end{array}\right.$ (4.17)
Let $a_{1}\leq\cdot\cdot\cdot\leq a_{n}$ be the eigenvalues of $\omega+\mathcal{D}^{+}_{J}(\varphi)$ with respect to $\omega$ at point $x$. Then, by (4.16), we can easily get $a_{n}\geq\cdot\cdot\cdot\geq a_{1}\geq\frac{1}{2}$ and $a_{1}\cdot\cdot\cdot a_{n}\geq\frac{1}{2^{n-1}}a_{n}$. On the other hand, by the generalized Monge-Ampère equation (3.1), we have
$\frac{1}{2^{n-1}}a_{n}\leq a_{1}\cdot\cdot\cdot a_{n}=e^{F(x)}.$
Thus, $a_{1}\leq\cdot\cdot\cdot\leq a_{n}\leq 2^{n-1}e^{F(x)}=C_{2}$. Then, for any $x\in P$,
$\omega+\mathcal{D}^{+}_{J}(\varphi)\leq C_{2}\omega.$
From this, using again (4.12) and $|D\varphi(x)|\leq\varepsilon/2$, we conclude that for all $x\in P$ we have
$0\leq(H(\varphi)-E(\varphi)(x)\leq C_{3}{\rm Id},$
and (4.15) gives
$\det(D^{2}\varphi)(x)\leq C_{4}.$
Applying the modified Aleksandrov-Bakelman-Pucci maximum principle in [49, Proposition 10] we obtain
$\varepsilon^{2n}\leq C_{2n}\int_{P}\det(D^{2}\varphi)$
for a constant $C_{2n}$ which depends only on the dimension of $M^{2n}$, and so
$\varepsilon^{2n}\leq C_{5}|P|,$
where $|P|$ denotes the Lebesgue measure of $P$. But for any $x\in P$ by (4.13) we also have
$\varphi(x)\leq\varphi(0)+\varepsilon/2=I+\varepsilon/2,$
and we may assume that $I+\varepsilon/2<0$, so
$|P|\leq\frac{\int_{P}(-\varphi)}{|I+\varepsilon/2|}\leq\frac{C_{6}}{|I+\varepsilon/2|}$
in view of Proposition $2.3$ in [15], where $C_{6}$ is a constant depending only on $(M^{2n},\omega,J)$. At last, we obtain
$|I+\varepsilon/2|\leq\frac{C}{\varepsilon^{2n}},$
where $C=C_{5}\cdot C_{6}$ is a constant depending only on $M^{2n},\omega,J$ and $F$. This gives us the desired uniform lower bound for $I$. This completes the proof of Proposition 4.5. ∎
Remark 4.6.
Let $\omega_{\frac{1}{2}}=\frac{1}{2}(\omega+\omega_{1})$, $g_{\frac{1}{2}}(\cdot,\cdot)=\omega_{\frac{1}{2}}(\cdot,J\cdot)$. Define $f_{\frac{1}{2}}$ as follows:
$-\frac{1}{2n}\Delta_{g_{\frac{1}{2}}}f_{\frac{1}{2}}=\frac{\omega_{\frac{1}{2}}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{\frac{1}{2}}^{n}},$
it implies
$-4n<\Delta_{g_{\frac{1}{2}}}f_{\frac{1}{2}}<4n,$
where $\Delta_{g_{\frac{1}{2}}}$ is the Laplacian of the Levi-Civita connection with respect to the almost Kähler metric $g_{\frac{1}{2}}$.
As done in Tosatti-Weinkove-Yau [61, Theorem 3.1] or Weinkove [69, Theorem 3.1], we can give a estimate for ${\rm tr}_{g}g_{1}$ (this is viewed as the generalized second order estimate of the almost Kähler potential) as follows. Our proof is completely in accordance with the proof of Theorem $3.1$ in [61]. Readers familiar with this proof can ignore the following sketch and go directly to see Proposition 4.10.
We introduce some notation. Let $\theta^{i}$ and $\tilde{\theta}^{i}$ be local unitary coframes for $g$ and $g_{1}$, respectively. Denote by $\nabla^{1}_{g}$ and $\nabla^{1}_{g_{1}}$ the associated second canonical connections. We will use $\Theta$, $\Psi$ ($\tilde{\Theta}$, $\tilde{\Psi}$ ) to denote the torsion, curvature and so on with respect to $\nabla^{1}_{g}$ ($\nabla^{1}_{g_{1}}$). Define local matrices $(a^{i}_{j})$ and $(b^{i}_{j})$ by
$\tilde{\theta}^{i}=a^{i}_{j}\theta^{j},$ (4.18)
$\theta^{j}=b^{j}_{i}\tilde{\theta}^{i},$ (4.19)
such that $a^{i}_{j}b^{k}_{i}=\delta^{k}_{j}$. Define a function $u$ by
$u=a^{i}_{j}\bar{a}^{i}_{j}=\frac{1}{2}{\rm tr}_{g}g_{1}.$
We will prove Proposition 4.10 using the maximum principle. The key result which we need for this is the following lemma (cf. [61, Lemma 3.2]).
Lemma 4.7.
Suppose that $g_{1}$ is an almost Kähler metric and solves the generalized Monge-Ampère equation (3.1). Then
$\Delta_{g_{1}}\log u\geq\frac{1}{u}(\Delta_{g}F-2R+8N^{l}_{\bar{p}\bar{i}}\overline{N^{p}_{\bar{l}\bar{i}}}+2\bar{a}^{p}_{i}a^{p}_{j}b^{k}_{q}\bar{b}^{l}_{q}\mathcal{R}_{i\bar{j}k\bar{l}}),$
where $\mathcal{R}_{i\bar{j}k\bar{l}}=R^{j}_{ik\bar{l}}+4N^{r}_{\bar{l}\bar{j}}\overline{N^{i}_{\bar{r}\bar{k}}}$, $N^{l}_{\bar{p}\bar{i}}$ is the Nijenhuis tensor of $J$, $R^{j}_{ik\bar{l}}$ is the curvature tensor of the second canonical connection $\nabla^{1}_{g}$, and $R$ is the Hermitian scalar curvature of the second canonical connection $\nabla^{1}_{g}$. For more details we refer to Section $2$ in [61].
To prove this lemma, we use some general identities (Lemmas 4.8 and 4.9 below) which are independent of the generalized Monge-Ampère equation.
First, differentiating (4.18) and using the first structure equations we obtain
$-\tilde{\theta}^{i}_{k}\wedge\tilde{\theta}^{k}+\tilde{\Theta}^{i}=da^{i}_{j}\wedge\theta^{j}-a^{i}_{j}\theta^{j}_{k}\wedge\theta^{k}+a^{i}_{j}\Theta^{j}.$
Using (4.19) and rearranging, we have
$(b^{j}_{k}da^{i}_{j}-a^{i}_{j}b^{l}_{k}\theta^{j}_{l}+\tilde{\theta}^{i}_{k})\wedge\tilde{\theta}^{k}=\tilde{\Theta}^{i}-a^{i}_{j}\Theta^{j}.$ (4.20)
By the definition of the canonical connection, the right-hand side of (4.20) has no $(1,1)$ part. Hence there exist functions $a^{i}_{kl}$ such that
$b^{j}_{k}da^{i}_{j}-a^{i}_{j}b^{l}_{k}\theta^{j}_{l}+\tilde{\theta}^{i}_{k}=a^{i}_{kl}\tilde{\theta}^{l},$ (4.21)
which can be rewritten as
$da^{i}_{m}-a^{i}_{j}\theta^{j}_{m}+a^{k}_{m}\tilde{\theta}^{i}_{k}=a^{i}_{kl}a^{k}_{m}\tilde{\theta}^{l}.$ (4.22)
Also, here $g$ and $g_{1}$ are almost Kähler, from (4.20) we see that we have $a^{i}_{kl}=a^{i}_{lk}$. We will now calculate a formula for $\Delta_{g_{1}}u$.
Lemma 4.8.
(cf. [61, Lemma 3.3]) For $g_{1}$ and $g$ almost Kähler metrics, we have
$\frac{1}{2}\Delta_{g_{1}}u=a^{i}_{kl}\bar{a}^{i}_{pl}a^{k}_{j}\bar{a}^{p}_{j}-\bar{a}^{i}_{j}a^{k}_{j}\tilde{R}^{i}_{kl\bar{l}}+\bar{a}^{i}_{j}a^{i}_{r}b^{q}_{l}\bar{b}^{s}_{l}R^{r}_{jq\bar{s}}.$
Now let $\nu=\det(a^{i}_{j})$ and set $v=|\nu|^{2}=\nu\bar{\nu}$, which is the ratio of the volume forms of $g_{1}$ and $g$. We have the following lemma.
Lemma 4.9.
(cf. [61, Lemma 3.4]) For $g_{1}$ and $g$ almost Kähler metrics, the following identities hold:
1) $(d\partial_{J}\log v)^{(1,1)}=-R_{k\bar{l}}\theta^{k}\wedge\bar{\theta}^{l}+\tilde{R}_{k\bar{l}}a^{k}_{i}\bar{a}^{l}_{j}\theta^{i}\wedge\bar{\theta}^{j}$;
2) $\Delta_{g}\log v=2R-2\tilde{R}_{k\bar{l}}a^{k}_{i}\bar{a}^{l}_{i}$,
where $R$ is the Hermitian scalar curvature of $g$ and $R_{k\bar{l}}$ ($\tilde{R}_{k\bar{l}}$) is the Hermite-Ricci curvature tensor of the second canonical connection $\nabla^{1}_{g}$ ($\nabla^{1}_{g_{1}}$).
Using Lemmas 4.8 and 4.9 we can now prove Lemma 4.7. Notice that the generalized Monge-Ampère equation (3.1) : $(\omega+\mathcal{D}^{+}_{J}(f))^{n}=\omega_{1}^{n}=e^{F}\omega^{n}$ can be written locally as
$\det g_{1}=e^{2F}\det g,$ (4.23)
which is the same as the equation $(3.17)$ in the proof of Lemma $3.2$ in [61]. Then the proof of Lemma 4.7 is the same as the proof of Lemma $3.2$ in [61]. By Lemma 4.7 , we can give the following Proposition.
Proposition 4.10.
(cf. Weinkove [69, Theorem 3.1], Tosatti-Weinkove-Yau [61, Theorem 3.1]) Let $g_{1}(\cdot,\cdot)=\omega_{1}(\cdot,J\cdot)$ be an almost Kähler metric solving the generalized Monge-Ampère equation (3.1), where $g(\cdot,\cdot)=\omega(\cdot,J\cdot)$ is an almost Kähler metric and
$\omega_{1}$ $=$ $\omega+\mathcal{D}^{+}_{J}(f)$
$=$ $\omega+dJdf_{1}+da_{1}(f).$
Then there exist constants $C$ and $A$ depending only $\omega,J,\sup|F|$ and the lower bound of $\Delta_{g}F$ such that
${\rm tr}_{g}g_{1}\leq Ce^{A(f_{1}-\inf_{M}f_{1})}=C(M^{2n},\omega,J,F,\Delta_{g}F).$
Proof.
(sketch) It is easy to see that $u=\frac{1}{2}{\rm tr}_{g}g_{1}$ is bounded below away from zero by a positive constant depending only on $\inf F$. Hence there exists a constant $C^{\prime}$ depending only on $J$, $\inf F$, $\Delta_{g}F$ and $R$ such that
$|\frac{1}{u}(\Delta_{g}F-2R+8N^{l}_{\bar{p}\bar{i}}\overline{N^{p}_{\bar{l}\bar{i}}})|\leq C^{\prime}.$ (4.24)
Choose $A^{\prime}$ sufficiently large such that
$\mathcal{R}_{i\bar{j}k\bar{l}}+A^{\prime}\delta_{ij}\delta_{kl}\geq 0.$
Then
$\frac{1}{u}2\bar{a}^{p}_{i}a^{p}_{j}b^{k}_{q}\bar{b}^{l}_{q}\mathcal{R}_{i\bar{j}k\bar{l}}\geq-\frac{1}{u}2A^{\prime}\bar{a}^{p}_{i}a^{p}_{i}b^{k}_{q}\bar{b}^{k}_{q}=-A^{\prime}{\rm tr}_{g_{1}}g.$ (4.25)
Combining (4.24) and (4.25) with Lemma 4.7 we obtain
$\Delta_{g_{1}}\log u\geq-C^{\prime}-A^{\prime}{\rm tr}_{g_{1}}g.$
We apply the maximum principle to $(\log u-2A^{\prime}f_{1})$, where
$-\frac{1}{2n}\Delta_{g_{1}}f_{1}=\frac{\omega_{1}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{1}^{n}}$
which follows that
$\Delta_{g_{1}}f_{1}=2n-{\rm tr}_{g_{1}}g.$ (4.26)
Suppose that the maximum of this quantity is achieved at a point $x_{0}$. Then at this point, using (4.26)
$0\geq\Delta_{g_{1}}(\log u-2A^{\prime}f_{1})\geq-C^{\prime}+A^{\prime}{\rm tr}_{g_{1}}g-4A^{\prime}n.$
Hence,
${\rm tr}_{g_{1}}g(x_{0})\leq\frac{4A^{\prime}n+C^{\prime}}{A^{\prime}}.$
Suppose that $0<a_{1}\leq a_{2}\leq\cdot\cdot\cdot\leq a_{n}$. If $n\geq 3$,
$\sum^{n}_{i=1}\frac{1}{a_{1}\cdot\cdot\cdot\hat{a}_{i}\cdot\cdot\cdot a_{n}}\leq\frac{1}{(n-1)!}(\sum^{n}_{i=1}\frac{1}{a_{i}})^{n-1}.$
Using the inequality above, we see that
$\frac{{\rm tr}_{g}g_{1}}{2}\sqrt{\frac{\det g}{\det g_{1}}}\leq\frac{1}{(n-1)!}(\frac{{\rm tr}_{g_{1}}g}{2})^{n-1}.$
Hence, using Equation (3.1) and $\omega_{1}=\omega+dJdf_{1}+da_{1}(f)$ again, $u$ can be bounded from above in terms of ${\rm tr}_{g_{1}}g$ and $\sup F$, and so we obtain an estimate
$u(x_{0})\leq C^{\prime\prime}.$
It follows that for any $x\in M^{2n}$,
$\log u(x)-2A^{\prime}f_{1}(x)\leq\log C^{\prime\prime}-2A^{\prime}\inf_{M^{2n}}f_{1}$
and, after exponentiating, this proves the Proposition.
Then by (4.9), (4.10), (4.11), Proposition 4.5 and Proposition 4.10, we have the following proposition:
Proposition 4.11.
$\|n(e^{\frac{F}{n}}-1)\|_{C^{0}(g)}\leq\|dJdf_{1}+da_{1}(f)\|_{C^{0}(g)}\leq c_{1},\,\,\,\|ne^{\frac{F}{n}}\|_{C^{0}(g)}\leq\|g_{1}\|_{C^{0}(g)}\leq c_{2}$
and
$\|ne^{\frac{-F}{n}}\|_{C^{0}(g)}\leq\|g^{-1}_{1}\|_{C^{0}(g)}\leq c_{3},$
where $c_{1},c_{2}$ and $c_{3}$ are constants depending only on $M^{2n},\omega,J$ and $F$. Hence, the almost Kähler metric $c^{-1}g\leq g_{1}\leq cg$, where $c$ is a positive constant depending only on $M^{2n},\omega,J$ and $F$. If $F=0$, then $a_{1}=\cdot\cdot\cdot=a_{n}=0$ and $g_{1}=g$.
[…]
$C_{1}\leq|\Delta_{g}f_{1}|\leq C_{2},$ (4.27)
where $C_{1}$ and $C_{2}$ are constants depending only on $M^{2n},\omega,J$ and $F$. Then, with Proposition 4.5, we will obtain a $C^{2}$-estimate of $f_{1}$ depending only on $M^{2n},\omega,J$ and $F$.
Now we give an estimate on the derivative of an almost Kähler metric $g_{1}$ solving the generalized Monge-Ampère equation (3.1). By Proposition 4.10, we have obtained
${\rm tr}_{g}g_{1}\leq C,$
where $C$ is constant depending only $\omega,J$ and $F$. We consider the first order estimate for $g_{1}$, this is a generalization of the third-order estimate for almost Kähler potential [72] (see also [60]).
Proposition 4.12.
(cf. Theorem 4.1 in [61]) Let $g_{1}(\cdot,\cdot)=\omega_{1}(\cdot,J\cdot)$ be a solution of the generalized Monge-Ampère equation (3.1). Then there exists a constant $C_{0}$ depending only on $M^{2n},\omega,J,F$ such that
$|\nabla^{1}_{g}g_{1}|^{2}_{g}\leq C_{0},$
where $\nabla^{1}_{g}$ is the second canonical connection associated to $g,J$.
By Proposition 4.10 and Proposition 4.12, we obtain the following theorem.
Theorem 4.13.
Let $(M^{2n},J)$ be a closed almost complex manifold with a $J$-compatible symplectic form $\omega$. Let $F$ be a real function on $M^{2n}$ with
$\int_{M^{2n}}\omega^{n}=\int_{M^{2n}}e^{F}\omega^{n}.$
Then if $\omega_{1}$ is an almost Kähler form with $[\omega_{1}]=[\omega]$ and solving the Calabi-Yau equation
$\omega_{1}^{n}=e^{F}\omega^{n},$
there are $C^{\infty}$ a priori bounds on $\omega_{1}$ depending only on $M^{2n},\omega,J$ and $F$.
Proof.
(sketch) From Proposition 4.10, Proposition 4.11 and Proposition 4.12 we have the estimate
$\|g_{1}\|_{C^{1}(g)}\leq C,$
where $C$ depends on $M^{2n},\omega,J$ and $F$. It remains to prove the higher order estimates. Our approach is along the lines used by Tosatti, Weinkove and Yau to prove Theorem 1.3 in [61]. Note that
$\omega_{1}-\omega$ $=$ $\mathcal{D}^{+}_{J}(f)$
$=$ $dJdf_{1}+da_{1}(f),$
and $a_{1}(f)$ is defined only up to the addition of a harmonic $1$-form. From the definition of $f_{1}$ it follows that
$\left\{\begin{array}[]{ll}\omega_{1}^{n-1}\wedge da_{1}(f)=0,&\\ &\\ d^{-}_{J}Jdf_{1}+d^{-}_{J}a_{1}(f)=0,&\\ &\\ d^{*_{1}}a_{1}(f)=0,&\end{array}\right.$ (4.28)
which is an elliptic system. Note that the kernel of (4.28) consists of the harmonic $1$-forms. Then if $a_{1}(f)$ is in the kernel of (4.28), we have $\|da_{1}(f)\|^{2}_{L^{2}(g_{1})}\leq C$ after integrating by parts. Since $d^{*_{1}}a_{1}(f)=0$, we see that $a_{1}(f)$ is harmonic with respect to $g_{1}$ (cf. Remark A.2).
Fix any $10<\beta<$ 1. Since $g_{1}$ is uniformly bounded in $C^{\beta}$, by Proposition 4.5, we can apply the elliptic Schauder estimates to (4.26) to get a bound $\|f_{1}\|_{C^{2+\beta}(g_{1})}\leq C$. Then by Proposition 4.11, we also have $\|f_{1}\|_{C^{2+\beta}(g)}\leq C$. Hence the right-hand side of (4.28) is bounded in $C^{1+\beta}(g_{1})$ (also in $C^{1+\beta}(g)$), and the coefficients of the system have a $C^{\beta}(g_{1})$ bound; so assuming that $a_{1}(f)$ is orthogonal to the harmonic $1$-forms, the elliptic estimates applied to (4.28) give $C^{2+\beta}(g_{1})$ (also $C^{2+\beta}(g)$) bounds on $a_{1}(f)$. By differentiating the Calabi-Yau equation in a direction $\partial/\partial x^{i}$ we obtain
$\Delta_{g_{1}}(\partial_{i}f_{1})+\{{\rm lower-order\,\,terms}\}=2\partial_{i}F+g^{pq}\partial_{i}g_{pq},$ (4.29)
where the lower-order terms may contain up to two derivatives of $f_{1}$ or $a_{1}(f)$, and so are bounded in $C^{\beta}$. Applying the Schauder estimates again we get $\|f_{1}\|_{C^{3+\beta}(g_{1})}\leq C$, and using (4.26) again we get $\|a_{1}(f)\|_{C^{3+\beta}(g_{1})}\leq C$ since
$\left\{\begin{array}[]{ll}d^{-}_{J}a_{1}(f)=-d^{-}_{J}Jdf_{1}&\\ &\\ d^{*_{1}}a_{1}(f)=0&\end{array}\right.$
is an elliptic system. Hence, $\mathcal{D}^{+}_{J}(f)=dJdf_{1}+da_{1}(f)$ has $C^{3+\beta}(g_{1})$ bound. Now a bootstrapping argument using (4.29) and (4.28) gives the required higher-order estimates. This completes the proof of Theorem 4.13. ∎
Remark 4.14.
Our theorem above is similar to the existence theorem for Monge-Ampère equation on Kähler manifolds (cf. Yau [72]). By Proposition 4.5 (cf. [61, Lemma 5.1]), the $C^{\infty}$ a priori estimates are independent of $\alpha$-integral [54],
[…]
where $\alpha$ is a positive constant, $\varphi$ is a $C^{\infty}$ function on $M^{2n}$ satisfying
$\omega+\mathcal{D}^{+}_{J}(\varphi)>0.$
We now consider the following equation
$\omega_{t}^{n}=e^{tF+c_{t}}\omega^{n},$ (4.30)
where $\omega_{t}$ is required to be cohomologous to $\omega$ and compatible with $J$, and $c_{t}$ is the constant given by
$c_{t}=\log(\frac{\int_{M^{2n}}\omega^{n}}{\int_{M^{2n}}e^{tF}\omega^{n}}).$
Let
[…]
As done in Yau’s argument [72], by Theorem 3.5 and Theorem 4.13 we can show that $T$ is open in $[0,1]$ and $T$ is also closed in $[0,1]$. Thus, we have the following theorem:
Theorem 4.15.
Let $(M^{2n},\omega,J,g)$ be a closed almost Kähler manifold of dimension $2n$. Then for $F\in C^{\infty}(M^{2n})$ satisfying
$\int_{M^{2n}}e^{F}\omega^{n}=\int_{M^{2n}}\omega^{n},$
there exists a unique almost Kähler form $\tilde{\omega}=\omega+\mathcal{D}^{+}_{J}(f)$ cohomologous to $\omega$ solving
$\tilde{\omega}^{n}=e^{F}\omega^{n},$
and there are $C^{\infty}$ a priori bounds on $\tilde{\omega}$ depending only on $\omega,J,F$.
5 Calabi Conjecture for Almost Kähler Manifolds
In this section, we devote to considering the generalized Calabi conjecture for almost Kähler setting.
Let $M$ be a compact Kähler manifold with Kähler metric $\sum g_{i\bar{j}}dz^{i}\otimes d\bar{z}^{j}$. Let $\sum R_{i\bar{j}}dz^{i}\otimes d\bar{z}^{j}$ be the Ricci tensor of $M$. Then, the $(1,1)$ form $\frac{\sqrt{-1}}{2\pi}\sum R_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ is closed and represents the first Chern class of $M$. In [9, 10], Calabi asked, given a closed $(1,1)$ form $\frac{\sqrt{-1}}{2\pi}\sum\widetilde{R}_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ which represents the first Chern class of $M$, can one find $\sum\widetilde{g}_{i\bar{j}}dz^{i}\otimes d\bar{z}^{j}$ a Kähler metric such that whose Kähler form determines the same cohomology class as $\frac{\sqrt{-1}}{2\pi}\sum g_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$.
For almost Kähler manifolds, we have the similar conjecture. As done in Kähler case, the conjecture is therefore equivalent to solving the generalized Monge-Ampère equation.
In [9], Calabi also conjectured the existence of Kähler-Einstein metric on a Kähler manifold whose first Chern class is negative, zero or positive. Kähler-Einstein metrics are the metrics whose Ricci form is proportional to the Kähler form. If the first Chern class is negative, then definition we can choose Kähler metric so that $\frac{\sqrt{-1}}{2\pi}\sum g_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ represents $-c_{1}(M)$. If $c_{1}(M)$ is zero, we choose $\frac{\sqrt{-1}}{2\pi}\sum g_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ to be arbitrary Kähler form. If $c_{1}(M)$ is positive, we choose $\frac{\sqrt{-1}}{2\pi}\sum g_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ to represent $c_{1}(M)$. In these case, the existence of Kähler-Einstein metric is equivalent to solve the following equation
$\det(g_{i\bar{j}}+\frac{\partial^{2}\varphi}{\partial z_{i}\partial\bar{z}_{j}})\det(g_{i\bar{j}})^{-1}=\exp(c\varphi+F),$ (5.1)
where $c=1$, 00 or $-1$ and $F$ is a smooth function defined on $M$ that satisfies the following equality when $c=0$,
$\int_{M}\exp Fdvol_{g}=vol_{g}(M).$ (5.2)
If $c_{1}(M)<0$, there exists a Kähler-Einstein metric (cf. Aubin [4], Yau [72]). If $c_{1}(M)=0$, the existence of Kähler-Einstein metric is equivalent the existence of the solution of the following complex Monge-Ampère equation
$\det(g_{i\bar{j}}+\frac{\partial^{2}\varphi}{\partial z_{i}\partial\bar{z}_{j}})\det(g_{i\bar{j}})^{-1}=\exp(F),$
where $F$ is a smooth function defined on $M$ satisfying Equation (5.2).
For the $c_{1}(M)>0$ case, which is called Yau-Tian-Donaldson conjecture ([73, 56, 25]). In [58], Tian and Yau proved the for any $n$ between $3$ and $8$, there is a compact complex surface $M$ diffeomorphic to $\mathbb{CP}^{2}\#n\overline{\mathbb{CP}^{2}}$ such that $c_{1}(M)>0$ admits a Kähler-Einstein metric. In [55], Tian obtained the following result: Any compact complex surface with $c_{1}(M)>0$ admits a Kähler-Einstein metric of $Lie(Aut(M))$ is reductive. In [56], Tian obtained a new kind of obstruction involving Geometric Invariant theory (see also Donaldson [25]). In 2015, Chen-Donaldson-Sun [14] and Tian [57] solved Yau-Tian-Donaldson conjecture independently.
It is nature to consider Hermite-Einstein almost Kähler metrics (HEAK for short) and extremal almost Kähler metrics. Let $(M^{2n},\omega)$ be a closed symplectic manifold of dimension $2n$. An almost complex structure $J$ is compatible with $\omega$ if the tensor field $g(\cdot,\cdot)=\omega(\cdot,J\cdot)$ defines a Riemannian metric on $M^{2n}$, then, $(\omega,J,g)$ is called an almost Kähler structure on $M^{2n}$. The canonical Hermitian connection $\nabla^{1}_{g}$ (which is also called the second canonical connection) corresponding to $(\omega,J,g)$ on the complex tangent bundle $TM^{2n}\otimes_{\mathbb{R}}\mathbb{C}$ is defined by
${\nabla^{1}_{g}}_{X}Y=D^{g}_{X}Y-\frac{1}{2}J(D^{g}_{X}J)Y,$
where $D^{g}$ is the Levi-Civita connection with respect to $g$ and $X,Y$ are vector fields on $M^{2n}$. Denote by $\Psi$ the curvature of $\nabla^{1}_{g}$ (as the definition in Section 4 ). Recall that
$(\Psi^{j}_{i})^{1,1}=R^{j}_{ik\bar{l}}\theta^{k}\wedge\bar{\theta}^{l};$
$(\Psi^{j}_{i})^{2,0}=K^{i}_{jkl}\theta^{k}\wedge\theta^{l};$
$(\Psi^{j}_{i})^{0,2}=K^{i}_{j\bar{k}\bar{l}}\bar{\theta}^{k}\wedge\bar{\theta}^{l}.$
For details, see [53, C.4] or [61]. Then, the Hermite-Ricci form $\rho$ is the trace of $\Psi_{X,Y}$ viewed as an anti-Hermitian linear operator
$\rho(X,Y)=-{\rm tr}(J\circ\Psi_{X,Y}).$ (5.3)
Hence, the $2$-form $\rho$ is a closed (real) $2$-form and it is a deRham representative of $2\pi c_{1}(M^{2n})$. We define the Hermitian scalar curvature $R$ of an almost Kähler metric $(\omega,J,g)$ as the trace of $\rho$ with respect to $\omega$,
$R\omega^{n}=2n(\rho\wedge\omega^{n-1}),$ (5.4)
or, in an equivalent way,
$R=2\Lambda_{\omega}\rho.$
Denote by $AK_{\omega}$ the Fréchet space of $\omega$-compatible almost complex structures. The space $AK_{\omega}$ comes naturally equipped with a formal Kähler structure ([44, 31, 23]). Let $Ham(M^{2n},\omega)$ be the group of Hamiltonian symplectomorphisms of $(M^{2n},\omega)$. The Lie algebra of $Ham(M^{2n},\omega)$ is identified with the space of smooth functions on $M^{2n}$ with zero mean value. A key observation, made by Fujiki [31] in the integrable case and by Donaldson [23] in the general almost Kähler case, asserts that the natural action of $Ham(M^{2n},\omega)$ on $AK_{\omega}$ is Hamiltonian with momentum given by the Hermitian scalar curvature. More precisely, the moment map
$\mu:AK_{\omega}\rightarrow(Lie(Ham(M^{2n},\omega)))^{*}$
is
$\mu_{J}(f)=\int_{M^{2n}}R(J)f\frac{\omega^{n}}{n!},$
where $R(J)$ is the Hermitian scalar curvature of $(\omega,J,g)$ and $f$ is a smooth function with zero mean value viewed as an element of $Lie(Ham(M^{2n},\omega))$ (cf. Lejmi [40]). The square-norm of the Hermitian scalar curvature defines the generalized Calabi functional on $AK_{\omega}$ [33]
$J\mapsto\int_{M^{2n}}R(J)^{2}\frac{\omega^{n}}{n!}.$ (5.5)
Definition 5.1.
The critical points of the functional (5.5) are called extremal almost Kähler metrics.
This observation was used [3] to deduce that the critical points of (5.5) are precisely the almost Kähler metrics $(\omega,J,g)$ for which the symplectic gradient of their Hermitian scalar curvature ${\rm grad}_{\omega}R(J)$, defined as the symplectic dual of $ds$ and thus ${\rm grad}_{\omega}R(J)=J{\rm grad}_{g}R(J)$.
Proposition 5.2.
A metric $(\omega,J,g)$ is a critical point of (5.5) if and only if ${\rm grad}_{\omega}R(J)$ is a Killing vector field with respect to $g$.
A proof of Proposition 5.2 is given in [3, 33, 40].
Remark 5.3.
(cf. Lejmi [40]) It is clear that the symplectic gradient of $R(J)$ of such metric is a $J$-holomorphic vector field in the sense that its flow preserves the corresponding almost complex structure. In particular, extremal Kähler metrics in the sense of Calabi [11] and almost Kähler metrics with constant Hermitian scalar curvature (cHsc for short) are extremal.
An almost Kähler metric $(\omega,J,g)$ is called Hermite-Einstein (HEAK for short) if the Hermite-Ricci form $\rho$ is a (constant) multiple of the symplectic form $\omega$, i.e.
$\rho=\frac{R}{2n}\omega,$
so the Hermitian scalar curvature $R$ is constant (cf. [41]). For an almost Kähler structure $(\omega,J,g)$ on $M^{2n}$, $\rho$ can be written as
$\rho=\rho^{+}+\rho^{-},$
where $\rho^{+}$ is $J$-invariant part of $\rho$, $\rho^{-}$ is $J$-anti-invariant part of $\rho$. In general, $\rho^{-}\neq 0$ if $J$ is not integrable. The form $\rho$ is a de Rham representative of $2\pi c_{1}(M^{2n},J)$ in $H^{2}(M^{2n},\mathbb{R})$ where $c_{1}(M^{2n},J)$ is the first (real) Chern class. If we suppose that $\omega$ and $\tilde{\omega}$ are symplectic forms compatible with the same almost-complex structure $J$ and satisfy $\tilde{\omega}^{n}=e^{f}\omega^{n}$ for some real-valued function $f$ then
$\tilde{\rho}=\rho-\frac{1}{2}dJdf,$ (5.6)
where $\tilde{\rho}$ (resp. $\rho$) is the Hermite-Ricci form of $(\tilde{\omega},J,\tilde{g})$ (resp. $(\omega,J,g)$). Let $v=\frac{\tilde{\omega}^{n}}{\omega^{n}}$, then $f=\log v$. If we choose $\tilde{\omega}=\omega+\mathcal{D}^{+}_{J}(\varphi)$, $\varphi\in C^{\infty}(M,J)_{0}$, then the generalized Monge-Ampère operator is
$v$ $=$ $M(\varphi)$ (5.7)
$=$ $\begin{vmatrix}1+\mathcal{D}^{+}_{J}(\varphi)_{1\bar{1}}&\mathcal{D}^{+}_{J}(\varphi)_{1\bar{2}}&\cdots&\mathcal{D}^{+}_{J}(\varphi)_{1\bar{n}}\\ \mathcal{D}^{+}_{J}(\varphi)_{2\bar{1}}&1+\mathcal{D}^{+}_{J}(\varphi)_{2\bar{2}}&\cdots&\mathcal{D}^{+}_{J}(\varphi)_{2\bar{n}}\\ \vdots&\vdots&\cdots&\vdots\\ \mathcal{D}^{+}_{J}(\varphi)_{n\bar{1}}&\mathcal{D}^{+}_{J}(\varphi)_{n\bar{2}}&\cdots&1+\mathcal{D}^{+}_{J}(\varphi)_{n\bar{n}}\\ \end{vmatrix}.$
Here, by Proposition A.1, we have
$\mathcal{D}^{+}_{J}(\varphi)$ $=$ $dJd\varphi+dd^{*}\sigma(\varphi)$
$=$ $dJd\tilde{\varphi}+d\tilde{a}(\varphi),$
where
$-\frac{1}{2n}\Delta_{\tilde{g}}\tilde{\varphi}=\frac{\tilde{\omega}^{n-1}\wedge(\tilde{\omega}-\omega)}{\tilde{\omega}^{n}},$
and
$\left\{\begin{array}[]{ll}d^{\tilde{*}}\tilde{a}(\varphi)=0&,\\ &\\ d^{-}_{J}Jd\tilde{\varphi}+d^{-}_{J}\tilde{a}(\varphi)=0&,\\ &\\ \tilde{\omega}^{n-1}\wedge d\tilde{a}(\varphi)=0&.\end{array}\right.$ (5.8)
Further more, if $\tilde{g}$ is a HEAK metric, that is, $\tilde{\rho}=\lambda\tilde{\omega}$, then
$\frac{\lambda}{2}dJd\tilde{\varphi}$ $=$ $\tilde{\rho}-\rho-\frac{1}{2}dJdf-\frac{\lambda}{2}d\tilde{a}(\varphi)$ (5.9)
$=$ $-\frac{1}{2}dJd\log M(\varphi)-\frac{1}{2}dJdf-\frac{\lambda}{2}d\tilde{a}(\varphi).$
Then (5.9) implies that $-\lambda\tilde{\varphi}=\log M(\varphi)+f$ since
$\frac{\lambda}{2}dJd\tilde{\varphi}\wedge\tilde{\omega}^{n-1}=(-\frac{1}{2}dJd\log M(\varphi)-\frac{1}{2}dJdf)\wedge\tilde{\omega}^{n-1}$
and
$-\lambda\Delta_{\tilde{g}}\tilde{\varphi}=\Delta_{\tilde{g}}\log M(\varphi)+\Delta_{\tilde{g}}f.$
Then the problem of the existence of HEAK metrics on almost Kähler manifold $(M^{2n},\omega,J,g)$ is equivalent to solve the equations
$\left\{\begin{array}[]{ll}\log M(\varphi)=\tilde{\varphi}-f,&~{}if~{}~{}c_{1}(M^{2n},J)<0;\\ &\\ \log M(\varphi)=-f+c,&~{}if~{}~{}c_{1}(M^{2n},J)=0;\\ &\\ \log M(\varphi)=-\tilde{\varphi}-f,&~{}if~{}~{}c_{1}(M^{2n},J)>0.\end{array}\right.$ (5.10)
Similar to Kähler case, we have the following theorem for HEAK metrics.
Theorem 5.4.
If $c_{1}(M^{2n},J)<0$, there exists a HEAK metric unique up to an homothety. If $c_{1}(M^{2n},J)=0$, there exists a HEAK metric unique up to an homothety in each positive $(1,1)$ cohomology chass.
Remark 5.5.
If $c_{1}(M^{2n},J)>0$, it is called symplectic Fano manifold. For Kähler case, Bando and Mabuchi [6] proved the uniqueness of Kähler Einstein metric on Fano manifolds. It is natural to ask the uniqueness of HEAK metrics for symplectic Fano manifolds.
Definition 5.6.
$\varphi$ admissible means that $\tilde{\omega}=\omega+\mathcal{D}^{+}_{J}(\varphi)$ is positive definite. $A$ will be the set of the $C^{2}$ admissible functions.
Proof of Theorem 5.4. If $c_{1}(M^{2n},J)=0$, it is just Theorem 4.15.
For the case of $c_{1}(M^{2n},J)<0$ in Theorem 5.4, it is equivalent to solve the generalized Monge-Ampère equation
$M(\varphi)=e^{\tilde{\varphi}-f}.$
Hence, we have
$\Delta_{\tilde{g}}\log M(\varphi)=\Delta_{\tilde{g}}\tilde{\varphi}-\Delta_{\tilde{g}}f.$ (5.11)
At $q$ a point of $M^{2n}$, where $\varphi$, a $C^{2}$ solution of (5.11), has a minimum, $\mathcal{D}^{+}_{J}(\varphi)_{\lambda\bar{\lambda}}\geq 0$ (that is $\Delta_{\tilde{g}}\tilde{\varphi}\geq 0$) for all directions $\lambda$. So at $q$, $\tilde{g}$ is positive definite, where $\tilde{g}=\tilde{\omega}(\cdot,J\cdot)$, $\tilde{\omega}=\omega+\mathcal{D}^{+}_{J}(\varphi)$ By continuity, no eigenvalue of $\tilde{g}$ can be zero since $M(\varphi)>0$. Hence $\varphi$ is admissible.
Consider the following mapping of the $C^{2}$ functions to $C^{0}$:
$\Gamma:\varphi\longrightarrow\tilde{\varphi}-f-\log M(\varphi).$ (5.12)
$\Gamma$ is continuously differentiable. Let $d\Gamma_{\varphi}$ denote its differential at $\varphi$:
$d\Gamma_{\varphi}(\psi)=\tilde{\psi}+\Delta_{\tilde{g}}\tilde{\psi}.$ (5.13)
Here $\Delta_{\tilde{g}}\tilde{\psi}=-\tilde{g}^{i\bar{j}}\mathcal{D}^{+}_{J}(\psi)_{i\bar{j}}$, and $\tilde{g}^{i\bar{j}}$ denotes the components of the inverse matrix of ${\tilde{g}}_{i\bar{j}}$.
Since $\varphi$ is admissible, Equation (5.11) is elliptic at $\varphi$ Hence, by ([5, Theorem 3.56]), $\varphi\in C^{2}$ $(\tilde{\varphi}\in C^{2})$ implies $\varphi\in C^{\infty}$ $(\tilde{\varphi}\in C^{\infty})$. Since
$\left\{\begin{array}[]{ll}d^{-}_{J}\tilde{a}(\varphi)=-d^{-}_{J}Jd\tilde{\varphi}&\\ &\\ d^{\tilde{*}}\tilde{a}(\varphi)=0&\end{array}\right.$
is an elliptic system, then $\tilde{a}(\varphi)\in C^{\infty}$.
Proposition 5.7.
Equation (5.11) has at most one solution, possibly up to $\ker\mathcal{W}_{J}$.
Proof.
This follows from the maximum principle ([5, Theorem 3.74]). Let $\varphi_{1}$ and $\varphi_{2}$ be two solutions of (5.11). According to the mean value Theorem [5, Theorem 3.6], there exists a function $\theta$ ($10<\theta<$ 1) such that $\psi=\varphi_{2}-\varphi_{1}$ satisfies
$\Delta_{g_{\gamma}}\tilde{\psi}+\tilde{\psi}=0\,\,\,{\rm with}\,\,\,\gamma=\varphi_{1}+\theta(\varphi_{2}-\varphi_{1}),$ (5.14)
where $g_{\gamma}(\cdot,\cdot)=\omega_{\gamma}(\cdot,J\cdot)$, $\omega_{\gamma}=\omega+\mathcal{D}^{+}_{J}(\gamma)$ Equation (5.14) has no solution except zero. ∎
Now we return to the proof of Theorem 5.4. We shall use the continuity method. For $t>0$ a parameter, let us consider the equation:
$\log M(\varphi)=\tilde{\varphi}-tf,$ (5.15)
with $f\in C^{3+\alpha}$. If for some t, Equation (5.15) has a $C^{2}$ solution $\varphi_{t}$, then $\varphi_{t}$ is unique, admissible, and belongs to $C^{5+\alpha}$, by the regularity and Proposition 5.7.
(1) The set of functions $f$, for which Equation (5.11) has a $C^{5+\alpha}$ solution is open in $C^{3+\alpha}$. To prove this, let us consider $\Gamma$, the mapping of the set $\Theta$ of the $C^{5+\alpha}$ admissible functions in $C^{3+\alpha}$ defined by:
$C^{5+\alpha}\supset\Theta\ni\varphi\stackrel{{\scriptstyle\Gamma}}{{\rightarrow}}\tilde{\varphi}-\log M(\varphi)\in C^{3+\alpha}.$
$\Gamma$ is continuously differentiable. Its differential at $\varphi$ is
$d\Gamma_{\varphi}(\psi)=\tilde{\psi}+\Delta_{g_{\varphi}}\tilde{\psi}.$
By ([5, Theorem 4.18 ]) the operator $d\Gamma_{\varphi}$ is invertible since the coefficient of $\psi$ is $1$. Since $C^{5+\alpha}$ and $C^{3+\alpha}$ are Banach spaces, we can use the inverse function theorem. Thus if $\varphi\in C^{5+\alpha}$ satisfies
$\log M(\varphi)=\tilde{\varphi}-f,$
there exists $V$, a $C^{3+\alpha}$ neighborhood of $f$, such that Equation (5.11) has a $C^{5+\alpha}$ solution when $f\in V$.
Return to Equation (5.15), where f is given. Because $\varphi_{0}=0$ is the solution of (5.15) for $t=0$, (5.15) has a solution for some interval $t\in[0,\tau]$, where $\tau>0$. Let $\tau$ be the largest real number such that Equation (5.15) has a solution for all $t\in[0,\tau]$. If $\tau\geq 1$, then $\varphi_{1}$ is the solution of (5.11) and it is proved. So suppose $\tau<1$, and come to a contradiction.
(2) We claim that the set $B$ of functions $\varphi_{t}$, $t\in[0,\tau]$, is bounded in $C^{2+\alpha}$, $10<\alpha<$ 1. If $\varphi_{t}$ has a maximum at $p$, then $M(\varphi_{t})\leq 1$. Thus,
$\tilde{\varphi}_{t}(q)-tf(q)\leq 0.$
Similarly, we have that if $\varphi_{t}$ has a minimum at $q$, then $M(\varphi_{t})\geq 1$ and
$\tilde{\varphi}_{t}(q)-tf(q)\geq 0.$
Hence $\sup|\tilde{\varphi}_{t}|\leq\tau\sup|f|$. Since
$\left\{\begin{array}[]{ll}d^{-}_{J}\tilde{a}(\varphi_{t})=-d^{-}_{J}Jd\tilde{\varphi}_{t}&\\ &\\ d^{\tilde{*}}\tilde{a}(\varphi_{t})=0&\end{array}\right.$
is an elliptic system, then $\tilde{a}(\varphi_{t})\in C^{1+\alpha}$, $\tilde{\varphi}_{t}\in C^{2+\alpha}$. The set $B$ is bounded in $C^{0}$, and $\mathcal{D}^{+}_{J}(\varphi)=dJd\tilde{\varphi}+d\tilde{a}(\varphi)$ is bounded in $C^{\alpha}$ for $\varphi\in B$.
(3) We now show that (5.15) has a solution for $t=\tau$ and hence for some $t>\tau$ by (1). This will give the desired contradiction. According to Ascoli’s theorem, the imbedding $C^{2+\alpha}\subset C^{2}$ is compact. Thus there exists $\varphi_{t}\in C^{2}$ and $t_{i}\rightarrow\tau$ an increasing sequence such that $\varphi_{t_{i}}$ converges to $\varphi_{\tau}$ in $C^{2}$.
Letting $i\rightarrow\infty$ in $\log M(\varphi_{t_{i}})=\tilde{\varphi}_{t_{i}}-t_{i}f$ we prove that $\tilde{\varphi}_{\tau}$, is the solution of (5.15) for $t=\tau$. According to the regularity theorem $\tilde{\varphi}_{\tau}\in C^{5+\alpha}$, $\tilde{a}_{\tau}(\varphi_{\tau})\in C^{5+\alpha}$, and the contradiction follows from (1), since (5.15) has a solution for $t$ in a neighborhood of $\tau$.
According to (5.10), finding an HEAK metric when $c_{1}(M^{2n},J)<0$ is equivalent to solving Equation (5.11). Thus, we complete the proof of Theorem 5.4.
6 Further Remarks and Questions
If $\omega$ is a Kähler form, let $K_{\omega}$ be the space of $\omega$-compatible complex structures, $AK_{\omega}$ the space of $\omega$-compatible almost complex structures. It is easy to see that $K_{\omega}\subsetneq AK_{\omega}$, where $AK_{\omega}$ is viewed as a contractible Fréchet manifold equipped with a formal Kähler structure. Suppose that $(g,\omega,J)$ is an almost Kähler metric on closed manifold $M^{2n}$. Let $\mathcal{H}(\omega,J)$ be the almost Kähler potential space, that is,
$\mathcal{H}(\omega,J):=\{f\in C^{\infty}(M,J)\,\,|\,\,\omega+\mathcal{D}^{+}_{J}(f)>0\}.$ (6.1)
[…]
$(\omega+\mathcal{D}^{+}_{\tilde{J}}(u))^{n+1}=0,$ (6.2)
where $\tilde{J}=(J_{st},J)$ is an almost complex structure on ×$\mathbb{C}\times M^{2n}$, $\tilde{J}$ is the standard complex structure. In order to study almost Kähler geometry, we need extend Futaki invariant [32], Mabuchi functional [43, 44], Tian $\alpha$-integral [54] and Ding functional [22].
As in Kähler case [27], on almost Kähler manifolds, we can consider existence question for four different kinds of special almost Kähler metrics, working within a fixed symplectic class on a compact manifold.
1) Extremal almost Kähler metrics
As in Kähler setting, extremal almost Kähler metrics are critical points of Calabi functional
$(\omega,J)\mapsto\int_{M^{2n}}R(J)^{2}\frac{\omega^{n}}{n!},$ (6.3)
where $R(J)$ is the Hermitian scalar curvature with respect the almost Kähler metric. It is interesting to consider the uniqueness of extremal almost Kähler metrics. For Kähler case, see [7]. Also, one can consider the relation between extremal Kähler metrics and extremal almost Kähler metrics. In the toric case, the existence of an extremal Kähler metric is conjecturallye equivalent to the existence of nonintegrable extremal almost Kähler metric [25] (see also [2]).
2) Constant Hermitian scalar curvature almost Kähler metrics (CHSC for short)
These metrics are just those whose Hermitian scalar curvature $R$ is constant. The constant Hermitian scalar curvature almost Kähler metrics are centainly extremal. Recently, Keller and Lejmi study $L^{2}$-norm of the Hermitian scalar curvature [38].
3) Hermite-Einstein almost Kähler metrics with $c_{1}>0$
This is an important question which is regared as the generalized Yau-Tian-Donaldson conjecture for symplectic Fano manifolds. For Kähler Fano manifolds, see Chen-Donaldson-Sun[14], Tian [57] and others.
4) The generalzied almost Kähler-Ricci solitons (GeAKRS for short)
For Kähler-Ricci solitons, there are a lot of paper for this topic, see [76, 67] and others. Inoue [37] fixed a compact subgroup $G$ of $Ham(M^{2n},\omega)$ and consider the group $Ham^{G}(M^{2n},\omega)$ of Hamiltonian symplectomorphisms commuting with $G$. Let $AK^{G}_{\omega}$ be the space of $G$-invariant almost complex structures compatible with $\omega$. He proved that the action of $Ham^{G}(M^{2n},\omega)$ on $AK^{G}_{\omega}$ with moment map given by
$R^{G}_{\xi}(J)=R^{G}(J)-n-2\Delta_{g}f_{\xi}+2f_{\xi}-2|f_{\xi}|^{2}_{g},$ (6.4)
where $f_{\xi}$ is a Hamiltonian potential of $\xi$ which is a fixed element in the center of the Lie algebra of $G$ on a compact symplectic Fano manifold. Inoue proved that the zeros of this moment map correspond to Kähler-Ricci solitons (cf. [37, Proposition 3.2]). If an almost Kähler metric $(g,J,\omega)$ satisfies the condition $R^{G}_{\xi}(J)=0$, we call it generalized almost-Kähler-Ricci soliton (GeAKRS for short). For the study of GeAKRS, see [1].
Appendix A The property of $\mathcal{D}_{J}^{+}$ operator
Suppose that $(M,\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. By the definition of $\mathcal{D}_{J}^{+}$, for any $f\in C^{\infty}(M)_{0}$,
$\mathcal{D}_{J}^{+}(f)=dJdf+dd^{*}\sigma(f),$
where $d^{*}=-*_{g}d*_{g}$, $\sigma(f)\in\Omega^{-}_{J}(M)$ and
$d_{J}^{-}dJdf+d_{J}^{-}d^{*}\sigma(f)=0.$
If $\omega_{1}=\omega+\mathcal{D}_{J}^{+}(f)>0$, then $\omega_{1}$ is a $J$-compatible symplectic form. Let
[…]
Then $\omega_{t}$ is a family of $J$-compatible symplectic forms in the same symplectic class $[\omega]$ and $g_{t}(\cdot,\cdot)=\omega_{t}(\cdot,J\cdot)$ is a family of almost Kähler metrics. It is easy to see that $(\omega,J,g)=(\omega_{0},J,g_{0})$. By direct calculation and Proposition 1.13.1 in [33], we have $d^{*_{t}}(f\omega_{t})=Jdf$, where $d^{*_{t}}=-*_{g_{t}}d*_{g_{t}}$. Then
$\mathcal{D}_{J}^{+}(f)$ $=$ $dJdf+dd^{*}\sigma(f)$ (A.1)
$=$ $dd^{*_{t}}(f\omega_{t})+da^{\prime}_{t}(f),$
where $a^{\prime}_{t}(f)=d^{*}\sigma(f)+d\varphi_{t}$ satisfying
$d^{*_{t}}a^{\prime}_{t}(f)=0,\,\,\,da^{\prime}_{t}(f)=dd^{*}\sigma(f),$
and $\varphi_{t}\in C^{\infty}(M)$. Hence, $Jdf+a^{\prime}_{t}(f)$ is $d^{*_{t}}$ closed.
Define
$\Omega_{1}^{2,t}(M)=\{f\omega_{t}\mid f\in C^{\infty}(M)\}.$
Recall the decomposition, we have
$\Omega^{2}(M)=\Omega_{1}^{2,t}(M)\oplus\Omega_{J}^{-}(M)\oplus\Omega_{J,0}^{+,t}(M)$
where
$\Omega_{J,0}^{+,t}(M)=\{\beta\in\Omega_{J}^{+}(M)\mid\omega_{t}^{n-1}\wedge\beta=0\}.$
A differential $k$-form $B_{k}$ with $k\leq n$ on $M^{2n}$ is called $\omega_{t}$-primitive if it satisfies $\omega_{t}^{n-k+1}\wedge B_{k}=0$ [68, 62]. Let $B_{k}$ with $k<n$, then $dB_{k}=B^{0}_{k+1}+\omega_{t}\wedge B_{k-1}^{1}$ [62, Lemma 2.4]. For $\omega_{t}$-primitive k-forms,
$*_{g_{t}}\frac{1}{r!}\omega_{t}^{r}\wedge B_{k}=(-1)^{\frac{k(k+1)}{2}}\frac{1}{(n-k-r)!}\omega_{t}^{n-k-r}\mathcal{J}(B_{k}),$ (A.2)
where
$\mathcal{J}=\sum_{p,q}(\sqrt{-1})^{p-q}\prod^{p,q}$
projects a k-form on its (p,q) parts times the multiplicative factor $(\sqrt{-1})^{p-q}$ [68, 62]. By (A.2), it follows that for any $\alpha\in\Omega_{J}^{-}(M^{2n})$ and $\beta\in\Omega_{J,0}^{+,t}(M^{2n})$, we have
$*_{g_{t}}\alpha=\frac{\omega_{t}^{n-2}}{(n-2)!}\wedge\alpha,\quad*_{g_{t}}\beta=-\frac{\omega_{t}^{n-2}}{(n-2)!}\wedge\beta.$
Define a smooth function $f_{t}\in C^{\infty}(M)$ as follows:
$-\frac{1}{2n}\Delta_{g_{t}}f_{t}=\frac{\omega_{t}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{t}^{n}},$ (A.3)
where $\Delta_{g_{t}}$ is the Laplacian of the Levi-Civita connection with respect to the almost Kähler metric $g_{t}$. In general, $f\neq f_{0}$. Using the result of Tosatti-Weinkove-Yau [61, Lemma 2.5], we can easily obtain
$-\omega_{t}^{n-1}\wedge dJdf_{t}=\frac{1}{2n}\Delta_{g_{t}}f_{t}\cdot\omega_{t}^{n}.$
On the other hand, by a pair $(r,k)$ corresponding to the space
$\mathcal{L}^{r,k}(t)=\left\{A\in\Omega^{2r+k}(M)\mid A=\frac{1}{r!}L_{\omega_{t}}^{r}B_{k}\text{ with }\Lambda_{\omega_{t}}B_{k}=0\right\}.$ (A.4)
We have the result [63] that $d$ acting on $\mathcal{L}^{r,k}$ leads to most two terms
$d:\mathcal{L}^{r,k}(t)\longrightarrow\mathcal{L}^{r,k+1}(t)\oplus\mathcal{L}^{r+1,k-1}(t).$ (A.5)
Indeed we can define the decomposition of $d$ into linear differential operators $(\partial_{+,t},\partial_{-,t})$ with respect to $\omega_{t}$ by writing
$d=\partial_{+,t}+L_{\omega_{t}}\partial_{-,t}.$
By Lemma 2.5 in [63], we find that on a symplectic manifold $(M^{2n},\omega_{t})$, the symplectic differential operator $(\partial_{+,t},\partial_{-,t})$ satisfies the following:
  1. (i)
    $(\partial_{+,t})^{2}=(\partial_{-,t})^{2}=0$;
  2. (ii)
    $L_{\omega_{t}}(\partial_{+,t}\partial_{-,t})=-L_{\omega_{t}}(\partial_{-,t}\partial_{+,t})$;
  3. (iii)
    $[\partial_{+,t},L_{\omega_{t}}]=[L_{\omega_{t}}\partial_{-,t},L_{\omega_{t}}]=0$.
For any $1$-form $b$, we have
$db=\partial_{+,t}b+\omega_{t}\wedge\partial_{-,t}b,$ (A.6)
where $\partial_{+,t}b\in\Omega_{J}^{-}(M)\oplus\Omega_{J,0}^{+,t}(M)$ and $\partial_{-,t}b\in C^{\infty}(M,J)_{0}$. By (A.3), we have
$\frac{\omega_{t}^{n-1}\wedge d(Jdf+a_{t}^{\prime}(f))}{\omega_{t}^{n}}=\frac{\omega_{t}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{t}^{n}}=-\frac{1}{2n}\Delta_{g_{t}}f_{t}.$
By (A.6), we have
$\omega_{1}-\omega=d(Jdf+a_{t}^{\prime}(f))=\partial_{+,t}(Jdf+a_{t}^{\prime}(f))+\omega_{t}\wedge\partial_{-,t}(Jdf+a_{t}^{\prime}(f)).$
Then, we will get
$\partial_{-,t}(Jdf+a_{t}^{\prime}(f))=-\frac{1}{2n}\Delta_{g_{t}}f_{t}.$ (A.7)
Similarly,
$dJdf_{t}=\partial_{+,t}Jdf_{t}+\omega_{t}\wedge\partial_{-,t}Jdf_{t}.$
Moreover,
$P^{+}_{J}(\partial_{+,t}Jdf_{t})\in\Omega_{J,0}^{+,t}(M),\,\,\,\partial_{-,t}Jdf_{t}=-\frac{1}{2n}\Delta_{g_{t}}f_{t}.$ (A.8)
By (A.7) and (A.8), we have
$\omega_{1}-\omega-dJdf_{t}$ $=$ $d(Jdf+a_{t}^{\prime}(f))-dJdf_{t}$
$=$ $\partial_{+,t}(Jdf+a_{t}^{\prime}(f))-\partial_{+,t}Jdf_{t},$
and
$P^{-}_{J}(\omega_{1}-\omega-dJdf_{t})=-d^{-}_{J}Jdf_{t}.$
Let $a_{t}(f)=Jdf+a_{t}^{\prime}(f)-Jdf_{t}$. Then
[…]
where
$\left\{\begin{array}[]{ll}d^{*_{t}}a_{t}(f)=0,&\\ &\\ d^{-}_{J}a_{t}(f)=-d^{-}_{J}Jdf_{t},&\\ &\\ \omega^{n-1}_{t}\wedge da_{t}(f)=0.\end{array}\right.$
Hence, one can find a linear map
$L(\omega,t):C^{\infty}(M,J)_{0}$$\rightarrow$ $C^{\infty}(M,J)_{0}$
[…]
which is injective.
Proposition A.1.
For any $f\in C^{\infty}(M)$, if $\omega_{1}=\omega+\mathcal{D}_{J}^{+}(f)>0$, one can define a family of $J$-compatible symplectic foems $\omega_{t}=t\omega_{1}+(1-t)\omega$, $0\leq t\leq 1$ and smooth functions $f_{t}\in C^{\infty}(M,J)_{0}$ by the following equations
$-\frac{1}{2n}\Delta_{g_{t}}f_{t}=\frac{\omega_{t}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{t}^{n}}.$
Then
$\omega_{1}-\omega=dJdf_{t}+da_{t}(f)=\mathcal{D}_{J,t}^{+}(f_{t}),$ (A.9)
where
$\left\{\begin{array}[]{ll}d^{*_{t}}a_{t}(f)=0,&\\ &\\ d^{-}_{J}a_{t}(f)=-d^{-}_{J}Jdf_{t},&\\ &\\ \omega^{n-1}_{t}\wedge da_{t}(f)=0.\end{array}\right.$ (A.10)
Thus, $L(\omega,t):C^{\infty}(M,J)_{0}\rightarrow L(\omega,t)(C^{\infty}(M,J)_{0})$ is an isomorphism.
Remark A.2.
1) By Theorem 2.1, we can find $\sigma(f_{t})\in\Omega^{-}_{J}(M)$ satisfying
$d^{-}_{J}Jdf_{t}+d^{-}_{J}d^{*_{t}}\sigma(f_{t})=0.$
But in general, $a_{t}(f)\neq d^{*_{t}}\sigma(f_{t})$. In fact, we have
$d^{-}_{J}(a_{t}(f)-d^{*_{t}}\sigma(f_{t}))=0.$
By Theorem 2.5, one can find $f^{\prime}_{t}\in C^{\infty}(M)$ such that
$a_{t}(f)-d^{*_{t}}\sigma(f_{t})=\mathcal{W}_{J}(f^{\prime}_{t})=Jdf^{\prime}_{t}+d^{*_{t}}\sigma(f^{\prime}_{t}).$
Thus, $a_{t}(f)=d^{*_{t}}\sigma(f_{t})+Jdf^{\prime}_{t}+d^{*_{t}}\sigma(f^{\prime}_{t})$.
2) Formula (A.10) is similar to (5.4) in [61]. The kernel of (A.10) consists of the harmonic $1$-forms. If $a_{t}(f)$ is in the kernel of (A.10), we have
$\|da_{t}(f)\|^{2}_{L^{2}(g_{t})}$ $=$ $\int_{M^{2n}}da_{t}(f)\wedge*_{t}da_{t}(f)$
[…]
$=$ $0$
Since $d^{*_{t}}a_{t}(f)=0$, we see that $a_{t}(f)$ is harmonic with respect to $g_{t}$.
3) For $n=2$, one can use self-dual equation to prove Proposition A.1 (see [69, 66]).
Acknowledgements.  The third author is very grateful to his advisor Z. Lü for his support; the authors thank Haisheng Liu and Zuyi Zhang for some helpful discussions.