### v4

# Qiang Tan, Hongyu Wang, Ken Wang, Zuyi Zhang

Abstract: In this paper, we introduce $\mathcal{D}^{+}_{J}$, a generalization of $\partial\bar{\partial}$ operator on higher dimensional almost Kähler manifolds. Using the $\mathcal{D}^{+}_{J}$ operator, we investigate the $\bar{\partial}$-problem in almost Kähler geometry and explore the generalized Monge-Ampère equation on almost Kähler manifolds. We establish a uniqueness up to the addition of a constant and local existence theorem for this equation. At last, we find an elliptical system for $\mathcal{D}^{+}_{J}$ operator. As an application, we reorganize the result of Tosatti-Weinkove-Yau in [46].
AMS Classification (2000): 53D35; 53C56; 53C65; 32Q60.
Keywords: almost Kähler form, $\mathcal{D}^{+}_{J}$ operator, $\bar{\partial}$-problem in almost Kähler geometry, the generalized Monge-Ampère equation.
1 Introduction
Yau’s Theorem for closed Kähler manifolds [59] asserts that one can prescribe the volume form of a Kähler metric within a given Kähler class. This result is fundamental in the theory of Kähler manifolds (cf. Calabi [7, 8]) and has broad applications in both geometry and mathematical physics (cf. Yau [58] ).
More precisely, Yau’s Theorem is stated as follows: Let $(M^{2n},\omega,J,g)$ be a closed Kähler manifold of real dimension $2n$, where $\omega$ denotes the Kähler form, $J$ is an $\omega$-compatible complex structure, and $g(\cdot,\cdot)=\omega(\cdot,J\cdot)$. Given a smooth volume form $e^{F}\omega^{n}$ on $M^{2n}$, there exists a unique smooth function $\varphi$ that satisfies
$(\omega+\sqrt{-1}\partial_{J}\bar{\partial}_{J}\varphi)^{n}$ $=$ $e^{F}\omega^{n},$
$\omega+\sqrt{-1}\partial_{J}\bar{\partial}_{J}\varphi$ $>$ $0,\,\,\,\sup_{M^{2n}}\varphi=0$ (1.1)
as long as $F$ satisfies the necessary normalization condition that
$\int_{M^{2n}}e^{F}\omega^{n}=\int_{M^{2n}}\omega^{n}.$ (1.2)
Equation (1) is known as the complex Monge-Ampère equation for Kähler manifolds. The solution method involves the technique of continuity and requires detailed a priori estimates for the derivatives of $\varphi$ up to third order (cf. Yau [59]). Finally, the uniqueness of the solution $\varphi$ was established with an elementary proof by Calabi [8].
An alternative extension of Yau’s Theorem arises in the case where $J$ is a non-integrable almost complex structure. Recall that an almost complex structure $J$ is said to be tamed by a symplectic form $\omega$ if the bilinear form $\omega(\cdot,J\cdot)$ is positive definite. The structure $J$ is termed compatible (or calibrated) with $\omega$ when the same bilinear form is symmetric, that is, when $\omega(\cdot,J\cdot)>0$ and $\omega(J\cdot,J\cdot)=\omega(\cdot,\cdot)$ (cf. McDuff-Salamon [36]). In the 1990s, Gromov posed the following problem to P. Delanoë [15]: Let $(M^{2n},\omega)$ be a closed symplectic manifold, $J$ an almost complex structure compatible with $\omega$, and $F$ a smooth function on $M^{2n}$ such that
$\int_{M^{2n}}e^{F}\omega^{n}=\int_{M^{2n}}\omega^{n}.$
Is it possible to find a smooth function $\varphi$ on $M^{2n}$ such that $\omega+dJd\varphi$ is a symplectic form taming $J$ and satisfies
$(\omega+dJd\varphi)^{n}=e^{F}\omega^{n}?$ (1.3)
However, P. Delanoë [15] showed that when $n=2$, the answer to this question is negative. This result was later extended to all dimensions by Wang and Zhu [52]. A key element of their findings is the construction of a smooth function $\varphi_{0}$ such that $\omega+dJd\varphi_{0}$ lies on the boundary of the set of taming symplectic forms (so its $(1,1)$ part is semipositive definite but not strictly positive), and yet $(\omega+dJd\varphi_{0})^{n}>0$. This outcome is possible because the $(2,0)+(0,2)$ part of $dJd\varphi_{0}$ contributes a strictly positive amount. This suggests that the issue with Gromov’s proposal is that the 2-form $dJd\varphi$ is generally not of type $(1,1)$ with respect to $J$, due to $J$ not being integrable in general. Its $(1,1)$ part is given by
$\sqrt{-1}\partial_{J}\bar{\partial}_{J}\varphi=\frac{1}{2}(dJd\varphi)^{(1,1)}$
which agrees with the standard notation when $J$ is integrable. Thus, one can study the Monge-Ampère equation for non-integrable almost complex structures. Several important works have been conducted when the base manifolds are almost complex. We refer to Harvey-Lawson [27], Tosatti-Wang-Weinkove-Yang [48], Chu-Tosatti-Weinkove [12], and the references therein.
To study the Donaldson tameness problem [19], Tan, Wang, Zhou, and Zhu introduced the operator $\mathcal{D}^{+}_{J}$ on tamed closed almost complex four-manifolds $(M^{4},J)$ in [43]. Using $\mathcal{D}^{+}_{J}$, they resolved this problem under the condition $h_{J}^{-}=b^{+}-1$. The operator $\mathcal{D}^{+}_{J}$ can be viewed as a generalization of $\partial_{J}\bar{\partial}_{J}$. Specifically, when $J$ is integrable, we have $\mathcal{D}^{+}_{J}=2\sqrt{-1}\partial_{J}\bar{\partial}_{J}$. In [51], Wang, Wang, and Zhu used the operator $\mathcal{D}^{+}_{J}$ to study the Nakai-Moishezon criterion for tamed almost Hermitian $4$-manifolds. In [53], Wang, Zhang, Zheng, and Zhu studied the generalized Monge-Ampère equation
$(\omega+\mathcal{D}^{+}_{J}(f))^{2}=e^{F}\omega^{2}$
on almost Kähler surfaces. It is well known that, in the four-dimensional case, the operator $d^{-}_{J}d^{*}$ $d^{-}_{J}d^{*}$ is a self-adjoint, strongly elliptic linear operator and plays a key role in defining the operator $\mathcal{D}^{+}_{J}$. However, in higher dimensions, $d^{-}_{J}d^{*}$ is no longer elliptic. Fortunately, through direct calculation, we find that the principal symbol of $d^{-}_{J}d^{*}$ is the same as that of $\partial_{J}\partial_{J}^{*}+\bar{\partial}_{J}\bar{\partial}_{J}^{*}$ on higher dimensional almost Kähler manifolds. Thus, the operator $d^{-}_{J}d^{*}:{\rm D}(d^{-}_{J}d^{*})\longrightarrow{\rm R}(d^{-}_{J}d^{*})$ is an invertible, formal self-adjoint, and nonnegative operator. Consequently, for any $f\in C^{\infty}(M)$, we have $d^{-}_{J}Jdf\,\bot\,{\rm coker}(d^{-}_{J}d^{*})$, and there exists a unique $\sigma(f)\in{\rm D}(d^{-}_{J}d^{*})$ such that
$d^{-}_{J}d^{*}\sigma(f)=d^{-}_{J}Jdf.$
Using $d^{-}_{J}d^{*}$, we can define the operator $\mathcal{D}^{+}_{J}$ on higher-dimensional almost Kähler manifolds as follows:
$\mathcal{D}^{+}_{J}(f)$ $=$ $dJdf+dd^{*}\sigma(f)$
$=$ $dd^{*}(f\omega)+dd^{*}\sigma(f)$
satisfying $d^{-}_{J}d^{*}(f\omega)+d^{-}_{J}d^{*}\sigma(f)=0$. Let $\mathcal{W}_{J}(f)=d^{*}(f\omega)+d^{*}\sigma(f)$, then
$\left\{\begin{array}[]{ll}d\mathcal{W}_{J}(f)=\mathcal{D}^{+}_{J}(f)&,\\ &\\ d^{*}\mathcal{W}_{J}(f)=0&,\end{array}\right$. (1.4)
which is an elliptic system. Notice that $d_{J}^{-}\mathcal{W}_{J}(f)=0$, similar to the $\bar{\partial}$-problem in classical complex analysis [28], $(\mathcal{W}_{J},d_{J}^{-})$-problem is whether $\mathcal{W}_{J}(f)=A$ has a solution for any $A$ satisfying $d_{J}^{-}A=0$. By using $L^{2}$-method [28], calculating the $L^{2}$-norm of $\mathcal{W}_{J}^{*}A$ and applying Riesz Representation Theorem [60], we solve the $(\mathcal{W}_{J},d^{-}_{J})$-problem in Section 2, which plays a crucial role in the subsequent study of the generalized Monge-Ampère equation.
Let $(M^{2n},\omega,J,g)$ be a closed almost Kähler manifold of dimension $2n$. We define a generalized Monge-Ampère equation on $M^{2n}$ as follows:
$(\omega+\mathcal{D}^{+}_{J}(f))^{n}=e^{F}\omega^{n}$ (1.5)
for some real function $f\in C^{\infty}(M)$ such that
$\omega+\mathcal{D}^{+}_{J}(f)>0,$
where $F\in C^{\infty}(M)$ satisfies
$\int_{M^{2n}}\omega^{n}=\int_{M^{2n}}e^{F}\omega^{n}.$ (1.6)
We establish a uniqueness up to the addition of a constant and local existence theorem for this equation in Section 3.
In Section 4, we find an elliptical system for $\mathcal{D}^{+}_{J}$ operator. For any $f\in C^{\infty}(M)$, if $\omega_{1}=\omega+\mathcal{D}_{J}^{+}(f)>0$, one can define a family of $J$-compatible symplectic forms $\omega_{t}=t\omega_{1}+(1-t)\omega$, $0\leq t\leq 1$ and smooth functions $f_{t}\in C^{\infty}(M)$ by the following equations
$\frac{1}{n}\Delta_{g_{t}}f_{t}=\frac{\omega_{t}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{t}^{n}}.$
Then
$\omega_{1}-\omega=dJdf_{t}+da(f_{t}),$ (1.7)
where
$\left\{\begin{array}[]{ll}d^{*_{t}}a(f_{t})=0,&\\ &\\ d^{-}_{J}a(f_{t})=-d^{-}_{J}Jdf_{t},&\\ &\\ \omega^{n-1}_{t}\wedge da(f_{t})=0.\end{array}\right$. (1.8)
As an application, we study the almost Kähler potentials $f_{0}$, $f_{1}$ and reorganize the result of Tosatti-Weinkove-Yau in [46].
Finally, as in the Kähler case [20], on almost Kähler manifolds, we raise several existence question for four different kinds of special almost Kähler metrics on compact manifolds, working within a fixed symplectic class.
1) Extremal almost Kähler metrics
2) Constant Hermitian scalar curvature almost Kähler metrics
3) Hermite-Einstein almost Kähler metrics with $c_{1}=0$, $c_{1}<0$ and $c_{1}>0$
4) Hermite-Ricci solitons
2 Preliminaries
Let $M$ be an almost complex manifold with an almost complex structure $J$. For any $x\in M$, $T_{x}M\otimes_{\mathbb{R}}\mathbb{C}$ which is the complexification of $T_{x}M$ can be decomposed as follows
$T_{x}M\otimes_{\mathbb{R}}\mathbb{C}=T^{1,0}_{x}+T^{0,1}_{x},$ (2.1)
where $T^{1,0}_{x}$ and $T^{0,1}_{x}$ are the eigenspaces of $J$ corresponding to the eigenvalues $\sqrt{-1}$ and $\sqrt{-1}$, respectively. A complex tangent vector is of type $(1,0)$ (resp. $(0,1)$) if it belongs to $T^{1,0}_{x}$ (resp. $T^{0,1}_{x}$). Let $TM\otimes_{\mathbb{R}}\mathbb{C}$ be the complexification of the tangent bundle. Similarly, let $T^{*}M\otimes_{\mathbb{R}}\mathbb{C}$ denote the complexification of the cotangent bundle $T^{*}M$. $J$ can act on $T^{*}M\otimes_{\mathbb{R}}\mathbb{C}$ as follows:
$\forall\alpha\in T^{*}M\otimes_{\mathbb{R}}\mathbb{C},\,\,\,J\alpha(\cdot)=-\alpha(J\cdot).$
Hence $T^{*}M\otimes_{\mathbb{R}}\mathbb{C}$ has the following decomposition according to the eigenvalues $\mp\sqrt{-1}$:
$T^{*}M\otimes_{\mathbb{R}}\mathbb{C}=\Lambda^{1,0}_{J}\oplus\Lambda^{0,1}_{J}.$ (2.2)
We define exterior bundle $\Lambda^{p,q}_{J}=\Lambda^{p}\Lambda^{1,0}_{J}\otimes\Lambda^{q}\Lambda^{0,1}_{J}$. Let $\Omega^{p,q}_{J}(M)$ denote the space of $C^{\infty}$ sections of the bundle $\Lambda^{p,q}_{J}$. The exterior differential operator acts on $\Omega^{p,q}_{J}$ as follows:
$d\Omega^{p,q}_{J}\subset\Omega^{p-1,q+2}_{J}+\Omega^{p+1,q}_{J}+\Omega^{p,q+1}_{J}+\Omega^{p+2,q-1}_{J}.$ (2.3)
Therefore, the exterior differential operator $d$ decomposes as:
$d=A_{J}\oplus\partial_{J}\oplus\bar{\partial}_{J}\oplus\bar{A}_{J},$ (2.4)
where each component is a derivation with bi-degrees given by
$|A_{J}|=(-1,2),\,\,|\partial_{J}|=(1,0),\,\,|\bar{\partial}_{J}|=(0,1),\,\,|\bar{A}_{J}|=(2,-1).$
The operators $\partial_{J}$ and $\bar{\partial}_{J}$ are of the first order while $A_{J}$ and $\bar{A}_{J}$ are of zero order (see [37]).
Now suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. Let $\Omega^{2}_{\mathbb{R}}(M)$ denote the space of real smooth $2$-forms on $M^{2n}$, i.e., the real $C^{\infty}$ sections of the bundle $\Lambda^{2}_{\mathbb{R}}(M)$. The almost complex structure $J$ acts on the space $\Omega_{\mathbb{R}}^{2}$ as an involution, defined by
$\alpha\longmapsto\alpha(J\cdot,J\cdot),\quad\alpha\in\Omega_{\mathbb{R}}^{2}(M).$ (2.5)
This induces a decomposition of 2-forms into $J$-invariant and $J$-anti-invariant parts (see [19]):
$\Omega_{\mathbb{R}}^{2}=\Omega^{+}_{J}\oplus\Omega^{-}_{J},\quad\alpha=\alpha_{J}^{+}+\alpha_{J}^{-}$
and the corresponding decomposition of vector bundles:
${\Lambda}_{\mathbb{R}}^{2}={\Lambda}_{J}^{+}\oplus{\Lambda}_{J}^{-}.$ (2.6)
We define the following operators:
$d^{+}_{J}$ $=$ $P^{+}_{J}d:\,\,\,\Omega_{\mathbb{R}}^{1}\longrightarrow\Omega^{+}_{J}$,
$d^{-}_{J}$ $=$ $P^{-}_{J}d:\,\,\,\Omega_{\mathbb{R}}^{1}\longrightarrow\Omega^{-}_{J},$ (2.7)
where $P^{\pm}_{J}:\Omega_{\mathbb{R}}^{2}\longrightarrow\Omega^{\pm}_{J}$. A differential $k$-form $B_{k}$ with $k\leq n$ is called primitive if $L^{n-k+1}_{\omega}B_{k}=0$, or equivalently, $\Lambda_{\omega}B_{k}=0$ (see [49, 57]). Here, $L_{\omega}$ is the Lefschetz operator (see [6, 49, 57]), which acts on a $k$-form $A_{k}\in\Omega_{\mathbb{R}}^{k}(M)$ as:
$L_{\omega}(A_{k})=\omega\wedge A_{k}.$
The dual Lefschetz operator is defined by $\Lambda_{\omega}:\Omega_{\mathbb{R}}^{k}(M)\rightarrow\Omega_{\mathbb{R}}^{k-2}(M)$, and it is a contraction map associated with the symplectic form $\omega$. We define the space of primitive $k$-forms as $\Omega^{k}_{0}(M)$. Specifically,
$\Omega^{2}_{0}(M)=\{\alpha\in\Omega_{\mathbb{R}}^{2}(M)\,|\,\omega^{n-1}\wedge\alpha=0\}.$
Thus,
$\Omega^{2}_{0}(M)=\Omega^{-}_{J}(M)\oplus\Omega^{+}_{J,0}(M)$
and
$\Omega_{\mathbb{R}}^{2}(M)$ $=$ $\Omega^{2}_{1}(M)\oplus\Omega^{2}_{0}(M)$ (2.8)
$=$ $\Omega^{2}_{1}(M)\oplus\Omega^{-}_{J}(M)\oplus\Omega^{+}_{J,0}(M),$
where $\Omega^{+}_{J,0}(M)$ is the space of the primitive $J$-invariant 22-forms and
$\Omega^{2}_{1}(M):=\{f\omega\,|\,f\in C^{\infty}(M,\mathbb{R})\}.$
Let
$d^{-}_{J}d^{*}:\Omega^{-}_{J}(M)\longrightarrow\Omega^{-}_{J}(M),$
where $d^{*}=-*_{g}d*_{g}$ and $_{g}$ is the Hodge star operator with respect to the metric $g$. For any $\alpha,\beta\in\Omega^{-}_{J}(M)$, it is straightforward to observe
$d^{-}_{J}d^{*}\alpha,\beta>_{g}=<dd^{*}\alpha,\beta>_{g}=<\alpha,dd^{*}\beta>_{g}=<\alpha,d^{-}_{J}d^{*}\beta>_{g}.$
Hence, $d^{-}_{J}d^{*}$ is a self-adjoint operator. If $\alpha\in\ker(d^{-}_{J}d^{*})\subset\Omega^{-}_{J}(M)$,
$0=<d^{-}_{J}d^{*}\alpha,\alpha>_{g}=<d^{*}\alpha,d^{*}\alpha>_{g}.$
Thus, we obtain
$\ker(d^{-}_{J}d^{*})={\rm coker}(d^{-}_{J}d^{*})=\{\alpha\in\Omega^{-}_{J}(M)\,\,|\,\,d^{*}\alpha=0\}.$
By Weil’s identity [49, 55],
$_{g}\alpha=\frac{1}{(n-2)!}\omega^{n-2}\wedge\alpha.$
So
$d*_{g}\alpha=\frac{1}{(n-2)!}\omega^{n-2}\wedge d\alpha=0.$
Therefore,
$\ker(d^{-}_{J}d^{*})={\rm coker}(d^{-}_{J}d^{*})$ $=$ $\{\alpha\in\Omega^{-}_{J}(M)\,\,|\,\,d^{*}\alpha=0\}$ (2.9)
$=$ $\{\alpha\in\Omega^{-}_{J}(M)\,\,|\,\,\omega^{n-2}\wedge d\alpha=0\}.$
If $n=2$, then $d^{*}\alpha=0$ implies $d\alpha=0$, meaning $\alpha$ is a $J$-anti-invariant harmonic $2$-form in $\Omega^{-}_{J}(M)$ (cf. [31]). Therefore, $\ker(d^{-}_{J}d^{*})=\mathcal{H}^{-}_{J}=\mathcal{H}^{2}\cap\Omega^{-}_{J}$. If $n\geq 3$, it is clear that $\mathcal{H}^{-}_{J}\subsetneq\ker(d^{-}_{J}d^{*})$.
Notice that $C^{\infty}(M)$ is dense in $L^{2}_{2}(M)$, so we can extend $d^{-}_{J}d^{*}$ to become a closed, densely defined operator (see [28]),
$d^{-}_{J}d^{*}:\Lambda^{-}_{J}\otimes L^{2}_{2}(M)\longrightarrow\Lambda^{-}_{J}\otimes L^{2}(M).$
In the sense of distributions, it is straightforward to see that
$\ker(d^{-}_{J}d^{*})=\{\alpha\in\Lambda^{-}_{J}\otimes L^{2}_{2}(M^{2n})\,\,|\,\,d^{-}_{J}d^{*}\alpha=0\}$
is closed. Specifically, let $\{\alpha_{i}\}$ be a sequence in $\ker(d^{-}_{J}d^{*})$ that converges in $L^{2}_{2}$ to some $\alpha\in\Lambda^{-}_{J}\otimes L^{2}_{2}(M)$. Then $\{d^{-}_{J}d^{*}\alpha_{i}\}=\{0\}$ is a constant sequence that converges to 0. Thus, $d^{-}_{J}d^{*}\alpha=0$, and $\alpha\in\ker(d^{-}_{J}d^{*})$, since $d^{-}_{J}d^{*}$ is a closed operator. Let
${\rm D}(d^{-}_{J}d^{*})=\Lambda^{-}_{J}\otimes L^{2}_{2}(M)\setminus\ker(d^{-}_{J}d^{*})$
and
${\rm R}(d^{-}_{J}d^{*})=\Lambda^{-}_{J}\otimes L^{2}(M)\setminus{\rm coker}(d^{-}_{J}d^{*}).$
Then,
$d^{-}_{J}d^{*}:{\rm D}(d^{-}_{J}d^{*})\longrightarrow{\rm R}(d^{-}_{J}d^{*})$
is invertible. For any $f\in L^{2}_{2}(M)$, by direct calculation and Proposition 1.13.1 in [26], we have
$Jdf=d^{*}(f\omega),\,\,d^{-}_{J}Jdf=d^{-}_{J}d^{*}(f\omega),\,\,d^{+}_{J}Jdf=2\sqrt{-1}\partial_{J}\bar{\partial}_{J}f.$
Note that
$d^{-}_{J}Jdf=d^{-}_{J}d^{*}(f\omega)\,\bot\,{\rm coker}(d^{-}_{J}d^{*}),$
so there exists a unique $\sigma(f)\in{\rm D}(d^{-}_{J}d^{*})$ such that
$d^{-}_{J}d^{*}(f\omega+\sigma(f))=0.$
We present the following theorem, which was proved by Lejmi in a $4$-dimensional case [31].
Theorem 2.1.
Let $(M^{2n},\omega,J,g)$ be a closed almost Kähler manifold of dimension $2n$. The operator $d^{-}_{J}d^{*}:{\rm D}(d^{-}_{J}d^{*})\longrightarrow{\rm R}(d^{-}_{J}d^{*})$ is an invertible, formal self-adjoint, and nonnegative operator. Therefore, for any $f\in C^{\infty}(M)$, it holds that $d^{-}_{J}Jdf\,\bot\,{\rm coker}(d^{-}_{J}d^{*})$ and there exists a unique $\sigma_{f}\in{\rm D}(d^{-}_{J}d^{*})$ such that
$d^{-}_{J}d^{*}\sigma(f)=d^{-}_{J}Jdf.$
Remark 2.2.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. Consider the second-order linear differential operator
$P_{J}:\Omega^{2}_{0}(M)$ $\rightarrow$ $\Omega^{2}_{0}(M)$
$\psi$ $\mapsto$ $\Delta_{g}\psi-\frac{1}{n}g(\Delta_{g}\psi,\omega)\omega,$
where $\Delta_{g}$ denotes the Riemannian Laplacian associated with the almost Kähler metric $g$ (with the convention that $\langle\omega,\omega\rangle_{g}=n$). The operator $P_{J}$ is a self-adjoint, strongly elliptic linear operator whose kernel consists of the primitive $g$-harmonic 2-forms (see Lejmi [31], Tan-Wang-Zhou [41]).
For the case $n=2$, Lejmi demonstrated that $P_{J}$ preserves the decomposition
$\Omega^{2}_{0}=\Omega^{+}_{J,0}\oplus\Omega^{-}_{J}.$
and furthermore, for $\psi\in\Omega^{+}_{J,0}$ and $\psi\in\Omega^{-}_{J}$, we have the relations $P_{J}|_{\Omega^{+}_{J,0}}(\psi)=\Delta_{g}\psi$ and $P_{J}|_{\Omega^{-}_{J}}(\psi)=2d^{-}_{J}d^{*}\psi$. Lejmi also noted that $P_{J}|_{\Omega^{-}_{J}}$ is a self-adjoint, strongly elliptic operator acting from $\Omega^{-}_{J}$ to $\Omega^{-}_{J}$ on a closed almost Kähler 4-manifold. In the general case, when $n>2$, the restriction
$P_{J}|_{\Omega^{-}_{J}}:\Omega^{-}_{J}\longrightarrow\Omega^{2}_{0}$
is an elliptic operator, as its symbol is injective, although not invertible [46].
As established in Tan-Wang-Zhou-Zhu [43] and Wang-Wang-Zhu [51], by applying Theorem 2.1, we define the operator $\mathcal{D}^{+}_{J}$ on higher-dimensional almost Kähler manifolds.
Definition 2.3.
Let $\mathcal{D}^{+}_{J}$ be the operator defined follows
$\mathcal{D}^{+}_{J}:L^{2}_{2}(M)\longrightarrow\Lambda^{+}_{J}\otimes L^{2}(M),$
with
$\mathcal{D}^{+}_{J}(f)=dJdf+dd^{*}\sigma(f)=dd^{*}(f\omega)+dd^{*}\sigma(f),$
where $\sigma(f)\in\Omega_{J}^{-}(M^{2n})$ and satisfies the condition
$d^{-}_{J}d^{*}(f\omega)+d^{-}_{J}d^{*}\sigma(f)=0.$
Let $\mathcal{W}_{J}:L^{2}_{2}(M)\longrightarrow\Lambda_{\mathbb{R}}^{1}\otimes L^{2}_{1}(M)$ be defined by
$\mathcal{W}_{J}(f)=d^{*}(f\omega+\sigma(f)).$
Then, we have the following relations:
$d\mathcal{W}_{J}(f)=\mathcal{D}^{+}_{J}(f)\text{ and }d^{-}_{J}\mathcal{W}_{J}(f)=0.$
The function $f$ is called an almost Kähler potential with respect to the almost Kähler metric $g$.
Remark 2.4.
If $J$ is integrable, i.e., $N_{J}=0$, then the operator $\mathcal{D}^{+}_{J}(f)$ simplifies into
$2\sqrt{-1}\partial_{J}\bar{\partial}_{J}f.$
Thus, in the case of integrability, $\mathcal{D}^{+}_{J}$can be viewed as a generalized form of the $\partial\bar{\partial}$ operator.
Denote the space of harmonic 2-forms by $\mathcal{H}^{2}_{dR}$ (cf. [9]). Let
$\mathcal{H}^{-}_{J}=\mathcal{H}^{2}_{dR}\cap\Omega^{-}_{J},\quad h^{-}_{J}=\dim\mathcal{H}^{-}_{J},$
and
$\mathcal{H}^{+}_{J,0}=\mathcal{H}^{2}_{dR}\cap\Omega^{+}_{J,0},\quad h^{+}_{J,0}=\dim\mathcal{H}^{+}_{J,0}$.
Here, $\mathcal{H}^{-}_{J}$ and $\mathcal{H}^{+}_{J,0}$ are the harmonic representations of their respective real de Rham cohomology groups on the manifold $M^{2n}$ (see Draghici-Li-Zhang [23] for $n=2$). As in Theorem 4.3 for the complex de Rham cohomology groups in [13], , we have the inequality
$h^{-}_{J}+h^{+}_{J,0}\leq b^{2}-1.$
If $h^{-}_{J}+h^{+}_{J,0}=b^{2}-1$, then
$\mathcal{H}^{2}_{dR}=Span\{\omega\}\oplus\mathcal{H}^{+}_{J,0}\oplus\mathcal{H}^{-}_{J}.$
Note that if $n=2$, then $\mathcal{H}^{2,0}_{dR}=\{0\}=\mathcal{H}^{0,2}_{dR}$ (cf. [13, Lemma 5.6]). As in the case of almost Kähler 4-manifolds (see Tan-Wang-Zhang-Zhu [40] and Tan-Wang-Zhou [42]), we can define
$\mathcal{H}^{\perp}_{J,0}=\mathbb{R}\cdot\omega\oplus\{\alpha_{f}=f\omega+d^{-}_{J}(v_{f}+\bar{v}_{f})\in\mathcal{H}^{2}_{dR}\,\,|\,\,f\in C^{\infty}(M),\,v_{f}\in\Omega^{0,1}_{J}(M)\}.$
Then, $h^{\perp}_{J,0}=\dim\mathcal{H}^{\perp}_{J,0}\geq 1$, and we have the inclusion
$\mathcal{H}^{\perp}_{J,0}\oplus\mathcal{H}^{+}_{J,0}\oplus\mathcal{H}^{-}_{J}\subseteq\mathcal{H}^{2}_{dR}.$
It is straightforward to see that
$\ker\mathcal{W}_{J}=Span_{\mathbb{R}}\{f_{2},\cdot\cdot\cdot,f_{h^{\perp}_{J,0}-1}\,\,|\,\,\alpha_{f_{i}}\in\mathcal{H}^{\perp}_{J,0}\}.$
If $n=2$, we have $\dim\mathcal{H}^{\perp}_{J,0}=b^{+}-\dim\mathcal{H}^{-}_{J}$, $0\leq\dim\mathcal{H}^{-}_{J}\leq b^{+}-1$. In this case, the direct sum is
$\mathcal{H}^{\perp}_{J,0}\oplus\mathcal{H}^{+}_{J,0}\oplus\mathcal{H}^{-}_{J}=\mathcal{H}^{2}_{dR},$
where $b^{+}$ is the self-dual second Betti number of $M^{4}$ (cf. Tan-Wang-Zhou [42]).
If $n\geq 3$, for $\alpha_{f}\in\mathcal{H}^{\perp}_{J,0}$, we have $d^{*}\alpha_{f}=0$ and
$d\alpha_{f}=\omega\wedge df+dd^{-}_{J}(v_{f}+\bar{v}_{f})=0.$
Thus, we obtain
$0$ $=$ $d*_{g}\alpha_{f}$
$=$ $d*_{g}[f\omega+d^{-}_{J}(v_{f}+\bar{v}_{f})]$
$=$ $d[\frac{1}{(n-1)!}f\omega^{n-1}+\frac{1}{(n-2)!}\omega^{n-2}\wedge d^{-}_{J}(v_{f}+\bar{v}_{f})]$
$=$ $\frac{1}{(n-1)!}\omega^{n-1}\wedge df+\frac{1}{(n-2)!}\omega^{n-2}\wedge dd^{-}_{J}(v_{f}+\bar{v}_{f})$
$=$ $\frac{1}{(n-1)!}\omega^{n-1}\wedge df-\frac{1}{(n-2)!}\omega^{n-2}\wedge\omega\wedge df$
$=$ $(\frac{1}{(n-1)!}-\frac{1}{(n-2)!})\omega^{n-1}\wedge df.$
By Corollary 2.7 in Yan [57], the operator $L_{\omega}^{n-1}:\Omega_{\mathbb{R}}^{1}\rightarrow\Omega_{\mathbb{R}}^{2n-1}$ is an isomorphism. Therefore, we conclude that $df=0$ and $f=c$. Consequently, $d^{-}_{J}(v_{f}+\bar{v}_{f})\in\mathcal{H}^{2}_{dR}$, and further, $d^{-}_{J}(v_{f}+\bar{v}_{f})\in\mathcal{H}^{-}_{J}$. Thus, we conclude that $d^{-}_{J}(v_{f}+\bar{v}_{f})=0$, and therefore, $\mathcal{H}^{\perp}_{J,0}=\mathbb{R}\cdot\omega$. Hence, $\ker\mathcal{W}_{J}=\mathbb{R}$.
Notice that $d_{J}^{-}\mathcal{W}_{J}(f)=0$, as $\bar{\partial}$-problem in classical complex analysis [28], $(\mathcal{W}_{J},d_{J}^{-})$-problem is whether $\mathcal{W}_{J}(f)=A$ has a solution for any $A$ satisfying $d_{J}^{-}A=0$. If we can use the theory of Hilbert space, considering
$L_{2}^{2}(M)\stackrel{{\scriptstyle\mathcal{W}_{J}}}{{\rightarrow}}\wedge_{\mathbb{R}}^{1}\otimes L_{1}^{2}(M)\stackrel{{\scriptstyle d_{J}^{-}}}{{\rightarrow}}\wedge_{J}^{-}\otimes L^{2}(M),$ (2.10)
then the above problem is equivalent to whether the kernel of $d_{J}^{-}$ is equal to the image of $\mathcal{W}_{J}$. Recall the definition of $\mathcal{W}_{J}$: $f\in L^{2}_{2}(M)$, $\sigma(f)\in{\rm D}(d^{-}_{J}d^{*})$ such that
$\mathcal{W}_{J}(f)=d^{*}(f\omega+\sigma(f))$
satisfying $d^{-}_{J}\mathcal{W}_{J}(f)=0$, $d^{*}\mathcal{W}_{J}(f)=0$ and
$d\mathcal{W}_{J}(f)=d^{+}_{J}\mathcal{W}_{J}(f)\in\Lambda^{+}_{J}\otimes L^{2}(M).$
Without loss of generality, we may assume that if $A\in\Omega^{1}_{\mathbb{R}}(M)$, $d^{*}A=0$ and $d^{-}_{J}A=0$, then
$<\mathcal{W}_{J}(f),A>_{g}$ $=$ $\int_{M^{2n}}A\wedge*_{g}d^{*}(f\omega+\sigma(f))$
$=$ $-\int_{M^{2n}}A\wedge d(f\frac{\omega^{n-1}}{(n-1)!}+\sigma(f)\wedge\frac{\omega^{n-2}}{(n-2)!})$
$=$ $-\int_{M^{2n}}dA\wedge(f\frac{\omega^{n-1}}{(n-1)!}+\sigma(f)\wedge\frac{\omega^{n-2}}{(n-2)!})$
$=$ $-\int_{M^{2n}}d^{+}_{J}A\wedge f\frac{\omega^{n-1}}{(n-1)!}$
$=$ $<f,\mathcal{W}_{J}^{*}A>_{g}.$
Thus, the formal $L^{2}$-adjoint operator of $\mathcal{W}_{J}$ is
$\mathcal{W}_{J}^{*}A=\frac{-n\omega^{n-1}\wedge d^{+}_{J}A}{\omega^{n}}=-(\Lambda_{\omega}d^{+}_{J}A).$ (2.11)
As done on closed almost Kähler $4$-manifolds, by using $L^{2}$-method [28], calculating the $L^{2}$-norm of $\mathcal{W}_{J}^{*}A$ and applying Riesz Representation Theorem [60], it is easy to get the following theorem (cf. [43, Appendix A.3] or [51, Section 6]):
Theorem 2.5.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. One can solve $(\mathcal{W}_{J},d^{-}_{J})$-problem.
Suppose $\psi\in\Lambda^{1,1}_{\mathbb{R}}\otimes L^{2}(M)$ is $d$-exact, hence
$\psi=d(u_{\psi}+\bar{u}_{\psi})=d^{+}_{J}(u_{\psi}+\bar{u}_{\psi}),\,\,\,i.e.,\,\,\,d^{-}_{J}(u_{\psi}+\bar{u}_{\psi})=0,$ (2.12)
for some $u_{\psi}\in\Lambda^{0,1}_{J}\otimes L^{2}_{1}(M)$. By Theorem 2.5, there exists $f\in L_{2}^{2}(M)$ such that $\mathcal{W}_{J}(f)=u_{\psi}+\bar{u}_{\psi}$. Hence,
$\psi=d(u_{\psi}+\bar{u}_{\psi})=d\mathcal{W}_{J}(f)=\mathcal{D}^{+}_{J}(f).$
We summarize the above discussion to obtain the following corollary:
Corollary 2.6.
Suppose that $\psi\in\Lambda^{+}_{J}\otimes L^{2}(M)$ is $d$-exact, that is, there is $u_{\psi}\in\Lambda^{0,1}_{J}\otimes L^{2}_{1}(M)$ such that $\psi=d(u_{\psi}+\bar{u}_{\psi})$. Then $\psi$ is $\mathcal{D}^{+}_{J}-$exact, that is, there exists $f_{\psi}\in L^{2}_{2}(M)$ such that $\psi=\mathcal{D}^{+}_{J}(f_{\psi})$.
With the above lemma, we can easily get the following proposition (cf. Tan-Wang-Zhou-Zhu [43]):
Proposition 2.7.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. $\mathcal{D}^{+}_{J}:L^{2}_{2}(M)\longrightarrow\Lambda^{+}_{J}\otimes L^{2}(M)$ has closed range.
Proof.
Let $\{f_{i}\}$ be a sequence of real functions on $M^{2n}$ in $L^{2}_{2}(M)$. By Definition 2.3, $\{\mathcal{W}_{J}(f_{i})\}$ is a sequence of real $1$-forms on $M^{2n}$ with coefficients in $L^{2}_{1}$ such that
$\psi_{i}=d\mathcal{W}_{J}(f_{i})=\mathcal{D}^{+}_{J}(f_{i})\in\Lambda^{+}_{J}\otimes L^{2}(M)$
is converging in $L^{2}$ to some $\psi\in\Lambda^{+}_{J}\otimes L^{2}(M)$. It is clear that $d^{*}\mathcal{W}_{J}(f_{i})=0$ and $\mathcal{W}_{J}(f_{i})$ is perpendicular to the harmonic $1$-forms. It is well known that $d+d^{*}$ is an elliptic system (cf. Donaldson-Kronheimer [21]). Hence there exists a constant $C$ such that
[…]
$=$ $C\|d\mathcal{W}_{J}(f_{i})\|^{2}_{L^{2}(M)}<Const.$
Hence, $\{\mathcal{W}_{J}(f_{i})\}$ is bounded in $L^{2}_{1}(M)$, so a subsequence of $\{\mathcal{W}_{J}(f_{i})\}$ converges weakly in $L^{2}_{1}$ to some $\mathcal{W}_{J}\in\Lambda^{1}_{\mathbb{R}}\otimes L^{2}_{1}(M)$. Since $d\mathcal{W}_{J}(f_{i})\in\Lambda^{+}_{J}\otimes L^{2}(M)$, it follows that
$d\mathcal{W}_{J}=\psi\in\Lambda^{+}_{J}\otimes L^{2}(M).$
By Lemma 2.6, there exists $f\in L^{2}_{2}(M)$ such that $\mathcal{D}^{+}_{J}(f)=d\mathcal{W}_{J}(f)=\psi$. This completes the proof of Proposition 2.7. ∎
3 Generalized Monge-Ampère Equation on Almost Kähler Manifolds
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. Define
$\mathcal{D}^{+}_{J}:C^{\infty}(M)\longrightarrow\Omega^{+}_{J}(M),$
by
$\mathcal{D}^{+}_{J}(f)=d\mathcal{W}_{J}(f)=dJdf+dd^{*}\sigma(f)$
satisfying $d^{-}_{J}\mathcal{W}_{J}(f)=0$, where $\mathcal{W}_{J}(f)=Jdf+d^{*}\sigma(f)$ and $\sigma(f)\in\Omega^{-}_{J}(M)$. It is easy to see that $\mathcal{D}^{+}_{J}$ is a linear operator and $d^{*}\mathcal{W}_{J}(f)=0$ with the fact that $Jdf=d^{*}(f\omega)$.
It is well known that Monge-Ampère equations are a class of important fully nonlinear equations profoundly related to many fields of analysis and geometry (see [24, 58]). The solution of the Monge-Ampère equation has been studied extensively. There are many existence, uniqueness and regularity results of the Monge-Ampère equation under different conditions and we refer the reader to Tosatti-Weinkove [47], Chu-Tosatti-Weinkove [12], Demailly-Pali [14], Pliś [38], Zhang-Zhang [61] and others. As done in almost Kähler surface in Wang-Zhang-Zheng-Zhu [53], we can define a generalized Monge-Ampère equation on $(M^{2n},\omega,J,g)$,
$(\omega+\mathcal{D}^{+}_{J}(f))^{n}=e^{F}\omega^{n}$ (3.1)
for some real function $f\in C^{\infty}(M)$ with
$\omega+\mathcal{D}^{+}_{J}(f)>0,$
where $F\in C^{\infty}(M)$ satisfies
$\int_{M^{2n}}\omega^{n}=\int_{M^{2n}}e^{F}\omega^{n}.$ (3.2)
If $J$ is integrable, $\mathcal{D}^{+}_{J}(f)=2\sqrt{-1}\partial_{J}\bar{\partial}_{J}f$. Then Equation (3.1) evolves into the famous Monge-Ampère equation (cf. [7, 59]), which is called the generalized Monge-Ampère equation on almost Kähler $2n$-manifolds.
By Definition 2.3, the generalized Monge-Ampère equation (3.1) is equivalent to the following Calabi-Yau equation for $1$-forms:
$(\omega+da)^{n}=e^{F}\omega^{n},$ (3.3)
where $a\in\Omega_{\mathbb{R}}^{1}(M)$, $d^{*}a=0$ and $d^{-}_{J}a=0$. In fact, we can choose $a=\mathcal{W}_{J}(f)$, $f\in C^{\infty}(M)$. For form-type Calabi-Yau equations, see Weinkove [56], Tosatti-Weinkove-Yau [46], Fu-Wang-Wu [24], Delanoë [15], Wang-Zhu [52] and others.
We want to consider the local theory of the Calabi-Yau equation on almost Kähler $2n$-manifolds (cf. [15, 52]).
Definition 3.1.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. The sets $A$, $B$, $A_{+}$ and $B_{+}$ are defined as follows:
$A:=\{u\in\Omega^{1}_{\mathbb{R}}(M)\,\,|\,\,d^{-}_{J}u=0,\,d^{*}u=0\};$
$A_{+}:=\{u\in A\,\,|\,\,\omega(u)=\omega+du>0\};$
$B:=\{f\in C^{\infty}(M)\,\,|\,\,\int_{M^{2n}}f\omega^{n}=\int_{M^{2n}}\omega^{n}\};$
$B_{+}:=\{f\in B\,\,|\,\,f>0\,\,on\,\,M^{2n}\}.$
We define an operator $\mathcal{F}$ from $A$ to $B$ as follows:
$u\mapsto\mathcal{F}(u),$
where
$\mathcal{F}(u)\omega^{n}=(\omega(u))^{n},\,\,\,\omega(u)=\omega+du.$ (3.4)
Restricting the operator $\mathcal{F}$ to $A_{+}$, we get $\mathcal{F}(A_{+})\subset B_{+}$. Thus, the existence of a solution to Equation (3.1) is equivalent to that the restricted operator
$\mathcal{F}|_{A_{+}}:A_{+}\rightarrow B_{+}$ (3.5)
is surjective. Moerover, we have the following result (cf. [15, Proposition 5] or [52, Proposition 2.4]):
Proposition 3.2.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$.
(1) If $u\in A_{+}$, then $\mathcal{F}(u)>0$;
(2) define
$B_{\varepsilon}(0):=\{u\in A\,|\,\|u\|_{C^{1}}\leq\varepsilon\},$
where $C^{1}$ is $C^{1}$-norm introduced by the metric $g$; if $\varepsilon<<1$, then $B_{\varepsilon}(0)\subset A_{+}$.
Suppose that $u_{0}\in A_{+}$. By Proposition 3.2, there exists a small neighborhood $U(u_{0})$ in $A$ such that $U(u_{0})\subset A_{+}$, and if $u_{1}\in A$, $\varepsilon<<1$, we can obtain that $\mathcal{F}(u_{0}+\varepsilon u_{1})>0$.
By solvability of $(\mathcal{W}_{J},d^{-}_{J})$-problem (see Theorem 2.5), for any $\phi\in A$, there exists $f_{\phi}\in C^{\infty}(M)$ such that $\mathcal{W}_{J}(f_{\phi})=\phi$. For any $u\in A_{+}$, since $A_{+}$ is an open subset of $A$, it is easy to see that the tangent space at $u$, $T_{u}A_{+}$, is $A$. For $\phi\in T_{u}A_{+}$, define $L(u)(\phi)$ by
$L(u)(\phi)=\frac{d}{dt}\mathcal{F}(u+t\phi)|_{t=0}.$
By simple calculation, we can obtain
$L(u)(\phi)$ $=$ $\frac{n\omega(u)^{n-1}\wedge\mathcal{D}^{+}_{J}(f_{\phi})}{\omega^{n}}.$
It is easy to see that $L(u)$ is a linear elliptic system on $A$ (cf. [15, 52]). Moreover, we have $\ker L(u)=\{0\}$. Indeed, for $u\in A_{+}$,
$\omega(u)=\omega+du=\omega+\mathcal{D}^{+}_{J}(f_{u})>0,$
that is, a symplectic form compatible with $J$ on $M^{2n}$, where $\mathcal{W}_{J}(f_{u})=u$. Let $g_{u}(\cdot,\cdot)=\omega(u)(\cdot,J\cdot)$. Then $(\omega(u),J,g_{u})$ is an almost Kähler structure on $M^{2n}$. By the primitive decomposition,
$\Omega^{+}_{J}(M)=\Omega^{2}_{1}(u)(M)\oplus\Omega^{+}_{J,0}(u)(M),$
where
$\Omega^{2}_{1}(u):=\{f\omega(u)\,|\,f\in C^{\infty}(M)\},$
and
$\Omega^{+}_{J,0}(u)(M)=\{\alpha\in\Omega^{+}_{J}(M)\,|\,\omega(u)^{n-1}\wedge\alpha=0\}.$
If $L(u)(\phi)=0$, then $\mathcal{D}^{+}_{J}(f_{\phi})$ is a primitive 2-form on $M^{2n}$. By Weil’s identity [49],
$_{g_{u}}\mathcal{D}^{+}_{J}(f_{\phi})=-\mathcal{D}^{+}_{J}(f_{\phi})\wedge\frac{\omega(u)^{n-2}}{(n-2)!}.$
Thus,
$<\mathcal{D}^{+}_{J}(f_{\phi}),\mathcal{D}^{+}_{J}(f_{\phi})>_{g_{u}}$ $=$ $\int_{M^{2n}}\mathcal{D}^{+}_{J}(f_{\phi})\wedge*_{g_{u}}\mathcal{D}^{+}_{J}(f_{\phi})$
$=$ $-\int_{M^{2n}}\mathcal{D}^{+}_{J}(f_{\phi})\wedge\mathcal{D}^{+}_{J}(f_{\phi})\wedge\frac{\omega(u)^{n-2}}{(n-2)!}$
$=$ $0.$
Therefore, $\mathcal{D}^{+}_{J}(f_{\phi})=d\mathcal{W}_{J}(f_{\phi})=0$. On the other hand, we have $d^{*}\mathcal{W}_{J}(f_{\phi})=0$. So $\mathcal{W}_{J}(f_{\phi})$ is a harmonic $1$-form [9]. By the definition of $\mathcal{W}_{J}(f_{\phi})$, we know that $\mathcal{W}_{J}(f_{\phi})$ is a $d^{*}$-exact form. Therefore, by Hodge decomposition, $\mathcal{W}_{J}(f_{\phi})=0$, that is, $\phi=0$. Hence, we have the following lemma (cf. [15, Proposition 1] or [52, Lemma 2.5]):
Lemma 3.3.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. Then the restricted operator
$\mathcal{F}:A_{+}\longrightarrow B_{+}$
is elliptic type on $A_{+}$. Moreover, the tangent map, $d\mathcal{F}(u)=L(u)$, of $\mathcal{F}$ at $u\in A_{+}$ is a linear elliptic system on $A$ and $\ker L(u)=\{0\}$.
Obviously, $A_{+}\subset A$ is a convex open set. Suppose that
$\mathcal{F}(u_{0})=\mathcal{F}(u_{1})$
for $u_{0},u_{1}\in A_{+}$. Let $u_{t}=tu_{1}+(1-t)u_{0}$, $t\in[0,1]$. So
$\int_{0}^{1}\frac{d}{dt}[\mathcal{F}(u_{t})]dt=0.$
Then
$L(u_{0})(u_{1}-u_{0})\omega^{n}=\int_{0}^{1}n\omega(u_{t})^{n-1}dt\wedge\mathcal{D}^{+}_{J}(f_{u_{1}}-f_{u_{0}})=0.$
Therefore $\mathcal{F}:A_{+}\rightarrow B_{+}$ is an injectivity map, and by nonlinear analysis (cf. Aubin [4]), we obtain the following result (cf. [15, Theorem 2] or [52, Proposition 2.6]):
Theorem 3.4.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. Then the restricted operator
$\mathcal{F}:A_{+}\rightarrow\mathcal{F}(A_{+})\subset B_{+}$
is a diffeomorphic map.
Let $F\in C^{\infty}(M)$ satisfying $e^{F}\in\mathcal{F}(A_{+})\subset B_{+}$, and
$\int_{M^{2n}}\omega^{n}=\int_{M^{2n}}e^{F}\omega^{n}.$
By the above theorem, there exists $u\in A_{+}$ such that $\mathcal{F}(u)=e^{F}$ and $\mathcal{F}(u)\omega^{n}=\omega(u)^{n}$. Then
$e^{F}\omega^{n}=\omega(u)^{n}=(\omega+du)^{n}=(\omega+d\mathcal{W}_{J}(f_{u}))^{n}=(\omega+\mathcal{D}^{+}_{J}(f_{u}))^{n}.$
Hence, with Theorem 3.4, we have the following local existence theorem for solution of generalized Monge-Ampère equation on the closed almost Kähler manifold $(M^{2n},\omega,J,g)$.
Theorem 3.5.
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold. Let $F\in C^{\infty}(M,\mathbb{R})$ satisfying
$\int_{M^{2n}}e^{F}\omega^{n}=\int_{M^{2n}}\omega^{n},$
and
$e^{F}\in\mathcal{F}(A_{+})\subset B_{+}.$
Then, there exists a smooth function $f\in C^{\infty}(M)$ satisfying
$(\omega+\mathcal{D}^{+}_{J}(f))^{n}=e^{F}\omega^{n}.$
The remainder of this section is devoted to studying a uniqueness theorem for the solution of generalized Monge-Ampère equation on closed almost Kähler manifold $(M^{2n},\omega,J,g)$. If there are two solutions $f_{0},f_{1}$ for equation (3.1), then
$(\omega+\mathcal{D}^{+}_{J}(f_{0}))^{n}=(\omega+\mathcal{D}^{+}_{J}(f_{1}))^{n}=e^{F}\omega^{n}.$
Let $f_{t}=tf_{1}+(1-t)f_{0}$, $t\in[0,1]$.
$\mathcal{D}^{+}_{J}(f_{t})$ $=$ $d\mathcal{W}_{J}(f_{t})$
$=$ $td\mathcal{W}_{J}(f_{1})+(1-t)d\mathcal{W}_{J}(f_{0})$
$=$ $td\phi_{1}+(1-t)d\phi_{0}$
$=$ $d\phi_{t},$
where $\phi_{0}=\mathcal{W}_{J}(f_{0})$, $\phi_{1}=\mathcal{W}_{J}(f_{1})$ and $\phi_{t}=t\phi_{1}+(1-t)\phi_{0}$ are all in $A_{+}$. So by the definition of operator $L$, we have
$0$ $=$ $\int_{0}^{1}\frac{d}{dt}(\omega+\mathcal{D}^{+}_{J}(f_{t}))^{n}dt$
$=$ $\int_{0}^{1}\frac{d}{dt}(\omega+d\phi_{t})^{n}dt$
$=$ $\int_{0}^{1}\frac{d}{dt}(\omega(\phi_{t}))^{n}dt$
$=$ $\int_{0}^{1}\frac{d}{dt}(\mathcal{F}(\phi_{t}))^{n}dt$
$=$ $L(\phi_{0})(\phi_{1}-\phi_{0})\omega^{n}.$
By Lemma 3.3, we know that $\ker L(u)=\{0\}$ for $u\in A_{+}$. Hence, $\phi_{1}=\phi_{0}$ and $\mathcal{W}_{J}(f_{1})=\mathcal{W}_{J}(f_{0})$. So we obtain a uniqueness theorem for the generalized Monge-Ampère equation up to $\ker\mathcal{W}_{J}$ (cf. Calabi [7] and Weinkove [56]).
Theorem 3.6.
The generalized Monge-Ampère equation (3.1) on almost Kähler $2n$-manifold has at most one solution up to $\ker\mathcal{W}_{J}$.
4 An elliptical system for $\mathcal{D}_{J}^{+}$ operator
Suppose that $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold of dimension $2n$. By the definition of $\mathcal{D}_{J}^{+}$, for any $f\in C^{\infty}(M)$,
$\mathcal{D}_{J}^{+}(f)=dJdf+dd^{*}\sigma(f),$
where $d^{*}=-*_{g}d*_{g}$, $\sigma(f)\in\Omega^{-}_{J}(M)$ and
$d_{J}^{-}Jdf+d_{J}^{-}d^{*}\sigma(f)=0.$
If $\omega_{1}=\omega+\mathcal{D}_{J}^{+}(f)>0$, then $\omega_{1}$ is a $J$-compatible symplectic form. Let
$\omega_{t}=t\omega_{1}+(1-t)\omega,\,\,\,0\leq t\leq 1.$
Then $\omega_{t}$ is a family of $J$-compatible symplectic forms in the same symplectic class $\omega]$ and $g_{t}(\cdot,\cdot)=\omega_{t}(\cdot,J\cdot)$ is a family of almost Kähler metrics. It is easy to see that $(\omega,J,g)=(\omega_{0},J,g_{0})$. By direct calculation and Proposition 1.13.1 in [26], we have $d^{*_{t}}(f\omega_{t})=Jdf$, where $d^{*_{t}}=-*_{g_{t}}d*_{g_{t}}$. Then
$\mathcal{D}_{J}^{+}(f)$ $=$ $dJdf+dd^{*}\sigma(f)$ (4.6)
$=$ $dd^{*_{t}}(f\omega_{t})+da^{\prime}_{t}(f),$
where $a^{\prime}_{t}(f)=d^{*}\sigma(f)+d\varphi_{t}$ satisfying
$d^{*_{t}}a^{\prime}_{t}(f)=0,\,\,\,da^{\prime}_{t}(f)=dd^{*}\sigma(f),$
and $\varphi_{t}\in C^{\infty}(M)$. Hence, $Jdf+a^{\prime}_{t}(f)$ is $d^{*_{t}}$ closed.
Define
$\Omega_{1}^{2,t}(M)=\{f\omega_{t}\mid f\in C^{\infty}(M)\}.$
Recall the decomposition, we have
$\Omega^{2}(M)=\Omega_{1}^{2,t}(M)\oplus\Omega_{J}^{-}(M)\oplus\Omega_{J,0}^{+,t}(M)$
where
$\Omega_{J,0}^{+,t}(M)=\{\beta\in\Omega_{J}^{+}(M)\mid\omega_{t}^{n-1}\wedge\beta=0\}.$
A differential $k$-form $B_{k}$ with $k\leq n$ on $M^{2n}$ is called $\omega_{t}$-primitive if it satisfies $\omega_{t}^{n-k+1}\wedge B_{k}=0$ [55, 49]. Let $B_{k}$ with $k<n$, then $dB_{k}=B^{0}_{k+1}+\omega_{t}\wedge B_{k-1}^{1}$ [49, Lemma 2.4]. For $\omega_{t}$-primitive k-forms,
$_{g_{t}}\frac{1}{r!}\omega_{t}^{r}\wedge B_{k}=(-1)^{\frac{k(k+1)}{2}}\frac{1}{(n-k-r)!}\omega_{t}^{n-k-r}\mathcal{J}(B_{k}),$ (4.7)
where
$\mathcal{J}=\sum_{p,q}(\sqrt{-1})^{p-q}\prod^{p,q}$
projects a k-form on its (p,q) parts times the multiplicative factor $(\sqrt{-1})^{p-q}$ [55, 49]. By (4.7), it follows that for any $\alpha\in\Omega_{J}^{-}(M)$ and $\beta\in\Omega_{J,0}^{+,t}(M)$, we have
$_{g_{t}}\alpha=\frac{\omega_{t}^{n-2}}{(n-2)!}\wedge\alpha,\quad*_{g_{t}}\beta=-\frac{\omega_{t}^{n-2}}{(n-2)!}\wedge\beta.$
Define a smooth function $f_{t}\in C^{\infty}(M)$ as follows:
$\frac{1}{n}\Delta_{g_{t}}f_{t}=\frac{\omega_{t}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{t}^{n}},$ (4.8)
where $\Delta_{g_{t}}$ is the Laplacian of the Levi-Civita connection with respect to the almost Kähler metric $g_{t}$. In general, $f\neq f_{0}$. Using the result of Tosatti-Weinkove-Yau [46, Lemma 2.5], we can easily obtain
$\omega_{t}^{n-1}\wedge dJdf_{t}=\frac{1}{n}\Delta_{g_{t}}f_{t}\cdot\omega_{t}^{n}.$
On the other hand, by a pair $(r,k)$ corresponding to the space
$\mathcal{L}^{r,k}(t)=\left\{A\in\Omega^{2r+k}(M)\mid A=\frac{1}{r!}L_{\omega_{t}}^{r}B_{k}\text{ with }\Lambda_{\omega_{t}}B_{k}=0\right\}.$ (4.9)
We have the result [50] that $d$ acting on $\mathcal{L}^{r,k}$ leads to most two terms
$d:\mathcal{L}^{r,k}(t)\longrightarrow\mathcal{L}^{r,k+1}(t)\oplus\mathcal{L}^{r+1,k-1}(t).$ (4.10)
Indeed we can define the decomposition of $d$ into linear differential operators $(\partial_{+,t},\partial_{-,t})$ with respect to $\omega_{t}$ by writing
$d=\partial_{+,t}+L_{\omega_{t}}\partial_{-,t}.$
By Lemma 2.5 in [50], we find that on a symplectic manifold $(M^{2n},\omega_{t})$, the symplectic differential operator $(\partial_{+,t},\partial_{-,t})$ satisfies the following:
  1. (i)
    $(\partial_{+,t})^{2}=(\partial_{-,t})^{2}=0$;
  2. (ii)
    $L_{\omega_{t}}(\partial_{+,t}\partial_{-,t})=-L_{\omega_{t}}(\partial_{-,t}\partial_{+,t})$;
  3. (iii)
    $\partial_{+,t},L_{\omega_{t}}]=[L_{\omega_{t}}\partial_{-,t},L_{\omega_{t}}]=0$.
For any $1$-form $b$, we have
$db=\partial_{+,t}b+\omega_{t}\wedge\partial_{-,t}b,$ (4.11)
where $\partial_{+,t}b\in\Omega_{J}^{-}(M)\oplus\Omega_{J,0}^{+,t}(M)$ and $\partial_{-,t}b\in C^{\infty}(M)$. By (4.8), we have
$\frac{\omega_{t}^{n-1}\wedge d(Jdf+a_{t}^{\prime}(f))}{\omega_{t}^{n}}=\frac{\omega_{t}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{t}^{n}}=-\frac{1}{n}\Delta_{g_{t}}f_{t}.$
By (4.11), we have
$\omega_{1}-\omega=d(Jdf+a_{t}^{\prime}(f))=\partial_{+,t}(Jdf+a_{t}^{\prime}(f))+\omega_{t}\wedge\partial_{-,t}(Jdf+a_{t}^{\prime}(f)).$
Then, we will get
$\partial_{-,t}(Jdf+a_{t}^{\prime}(f))=-\frac{1}{n}\Delta_{g_{t}}f_{t}.$ (4.12)
Similarly,
$dJdf_{t}=\partial_{+,t}Jdf_{t}+\omega_{t}\wedge\partial_{-,t}Jdf_{t}.$
Moreover,
$P^{+}_{J}(\partial_{+,t}Jdf_{t})\in\Omega_{J,0}^{+,t}(M),\,\,\,\partial_{-,t}Jdf_{t}=-\frac{1}{n}\Delta_{g_{t}}f_{t}.$ (4.13)
By (4.12) and (4.13), we have
$\omega_{1}-\omega-dJdf_{t}$ $=$ $d(Jdf+a_{t}^{\prime}(f))-dJdf_{t}$
$=$ $\partial_{+,t}(Jdf+a_{t}^{\prime}(f))-\partial_{+,t}Jdf_{t},$
and
$P^{-}_{J}(\omega_{1}-\omega-dJdf_{t})=-d^{-}_{J}Jdf_{t}.$
Let $a(f_{t})=Jdf+a_{t}^{\prime}(f)-Jdf_{t}$. Then
$\omega_{1}-\omega=dJdf_{t}+da(f_{t})=\mathcal{D}_{J,t}^{+}(f_{t}),\,\,\,0\leq t\leq 1,$
where
$\left\{\begin{array}[]{ll}d^{*_{t}}a(f_{t})=0,&\\ &\\ d^{-}_{J}a(f_{t})=-d^{-}_{J}Jdf_{t},&\\ &\\ \omega^{n-1}_{t}\wedge da(f_{t})=0.\end{array}\right$.
Theorem 4.1.
For any $f\in C^{\infty}(M)$, if $\omega_{1}=\omega+\mathcal{D}_{J}^{+}(f)>0$, one can define a family of $J$-compatible symplectic forms $\omega_{t}=t\omega_{1}+(1-t)\omega$, $0\leq t\leq 1$ and smooth functions $f_{t}\in C^{\infty}(M)$ by the following equations
$\frac{1}{n}\Delta_{g_{t}}f_{t}=\frac{\omega_{t}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{t}^{n}}.$
Then
$\omega_{1}-\omega=dJdf_{t}+da(f_{t})=\mathcal{D}_{J,t}^{+}(f_{t}),$ (4.14)
where
$\left\{\begin{array}[]{ll}d^{*_{t}}a(f_{t})=0,&\\ &\\ d^{-}_{J}a(f_{t})=-d^{-}_{J}Jdf_{t},&\\ &\\ \omega^{n-1}_{t}\wedge da(f_{t})=0.\end{array}\right$. (4.15)
Remark 4.2.
By Theorem 2.1, we can find $\sigma(f_{t})\in\Omega^{-}_{J}(M)$ satisfying
$d^{-}_{J}Jdf_{t}+d^{-}_{J}d^{*_{t}}\sigma(f_{t})=0.$
But in general, $a(f_{t})\neq d^{*_{t}}\sigma(f_{t})$. In fact, we have
$d^{-}_{J}(a(f_{t})-d^{*_{t}}\sigma(f_{t}))=0.$
By Theorem 2.5, one can find $f^{\prime}_{t}\in C^{\infty}(M)$ such that
$a(f_{t})-d^{*_{t}}\sigma(f_{t})=\mathcal{W}_{J}(f^{\prime}_{t})=Jdf^{\prime}_{t}+d^{*_{t}}\sigma(f^{\prime}_{t}).$
Thus, $a(f_{t})=d^{*_{t}}\sigma(f_{t})+Jdf^{\prime}_{t}+d^{*_{t}}\sigma(f^{\prime}_{t})$.
The remainder of this section is devoted to studying the almost Kähler potentials $f_{0}$ and $f_{1}$. By elementary linear algebra, using simultaneous diagonalization (see McDuff-Salamon [36]), for any $p\in M^{2n}$, it is possible to find complex coordinates $z_{1},\cdot\cdot\cdot,z_{n}$ on $M^{2n}$ near $p$ such that:
Lemma 4.3.
$g(p)=2(|dz^{1}|^{2}+\cdot\cdot\cdot+|dz^{n}|^{2})$;
$g_{1}(p)=2(a_{1}|dz^{1}|^{2}+\cdot\cdot\cdot+a_{n}|dz^{n}|^{2})$;
$\omega(p)=\sqrt{-1}(dz^{1}\wedge d\bar{z}^{1}+\cdot\cdot\cdot+dz^{n}\wedge d\bar{z}^{n})$;
$\omega_{1}(p)=\sqrt{-1}(a_{1}dz^{1}\wedge d\bar{z}^{1}+\cdot\cdot\cdot+a_{n}dz^{n}\wedge d\bar{z}^{n})$, where $0<a_{1}\leq\cdot\cdot\cdot\leq a_{n}$.
By using orthonormal coordinate system [9] and the result of Tosatti-Weinkove-Yau [46, Lemma 2.5], we have
$\Delta^{c}_{g}f_{0}(p)=\sqrt{-1}\sum^{n}_{j=1}(dJdf_{0})^{(1,1)}(p)(v_{j},\bar{v}_{j}),$ (4.16)
where $v_{j}=\partial/\partial z_{j}$ at $p$ and $\Delta^{c}_{g}$ is the complex Laplacian of the Hermitian canonical connection with respect to the almost Kähler metric $g$ at $p$. Also by Tosatti-Weinkove-Yau [46, Lemma 2.6], since $(M^{2n},\omega,J,g)$ is a closed almost Kähler manifold, then
$\Delta_{g}f_{0}=\Delta^{c}_{g}f_{0},$ (4.17)
where $\Delta_{g}$ is the Laplacian of the Levi-Civita connection with respect to the almost Kähler metric $g$. Thus, we can relate $a_{1},\cdot\cdot\cdot,a_{n}$ to $e^{F}$ and $\Delta_{g}f_{0}=\Delta^{c}_{g}f_{0}$. If $\omega_{1}^{n}=e^{F}\omega^{n}$ (that is Equation (3.1)), we have the following lemma:
Lemma 4.4.
For any $p\in M^{2n}$, find complex coordinates $z_{1},\cdot\cdot\cdot,z_{n}$ near $p$ such that:
$\Pi^{n}_{j=1}a_{j}=e^{F(p)},\,\,\,\frac{\partial^{2}f_{0}}{\partial z_{j}\partial\bar{z}_{j}}(p)=a_{j}-1,\,\,\,\frac{\partial^{2}f_{0}}{\partial z_{i}\partial\bar{z}_{j}}(p)=0\,\,\,for\,\,\,i\neq j$
and
$\Delta_{g}f_{0}(p)=n-\sum^{n}_{j=1}a_{j}.$
By the previous lemmas, it is easy to obtain the following lemma:
Lemma 4.5.
At $p\in M^{2n}$, we have
$|d\mathcal{W}_{J}(f)(p)|^{2}_{g}=|\mathcal{D}^{+}_{J}(f)(p)|^{2}_{g}=2\sum^{n}_{j=1}(a_{j}-1)^{2},\,\,\,|g_{1}(p)|^{2}_{g}=2\sum^{n}_{j=1}a_{j}^{2}$
and
$|g_{1}^{-1}(p)|^{2}_{g}=2\sum^{n}_{j=1}a_{j}^{-2}.$
By the first and last equations in Lemma 4.4 and the fact that the geometric mean $(a_{1}\cdot\cdot\cdot a_{n})^{\frac{1}{n}}$ is less than or equal to the arithmetic mean $\frac{1}{n}(a_{1}+\cdot\cdot\cdot+a_{n})$, it is easy to obtain the following inequality:
$\Delta_{g}f_{0}(p)\leq n-ne^{\frac{F(p)}{n}}<n.$ (4.18)
Since any almost Kähler metric is Gauduchon, it is nature to extend Proposition $2.3$ in [12] to almost Kähler setting.
Proposition 4.6.
Let $(M^{2n},\omega,J,g)$ be a closed almost Kähler manifold. Then there is a constant $C>0$ depending only on $(M^{2n},\omega,J,g)$ such that every smooth function $f_{0}$ on $M^{2n}$ which satisfies
$\omega+\mathcal{D}^{+}_{J}(f)=\omega+dJdf_{0}+da(f_{0})>0,\,\,\,\sup_{M^{2n}}f_{0}=0,$ (4.19)
also satisfies
$\int_{M^{2n}}(-f_{0})\omega^{n}<C.$ (4.20)
Proof.
Notice that $\Delta_{g}$ is the canonical Laplacian of $\omega$, which is an elliptic second order differential operator with kernel consisting of just constants. Standard linear PDE theory (cf. [2, Appendix A]) shows that there exists a Green function $G$ for $\Delta_{g}$ which satisfies $G(x,y)\geq-C$ and $\|G(x,\cdot)\|_{L^{1}(M,g)}\leq C$ for a constant $C>0$, and
$f_{0}(x)=\frac{1}{\int_{M^{2n}}\omega^{n}}\int_{M^{2n}}f_{0}\omega^{n}+\int_{M^{2n}}\Delta_{g}f_{0}(y)G(x,y)\omega^{n}(y)$ (4.21)
for all smooth functions $f_{0}$ and all $x\in M^{2n}$. On the other hand, with the third equation of (4.15), we have
$\int_{M^{2n}}\Delta_{g}f_{0}\omega^{n}=-n\int_{M^{2n}}\omega^{n-1}\wedge(dJdf_{0}+da(f_{0}))=0.$
Therefore, we are free to add a large uniform constant to $G(x,y)$ to make it nonnegative, while preserving the same Green formula.
If $f_{0}$ satisfies (4.19), that is $\mathcal{D}^{+}_{J}(f)>-\omega$, then by (4.8), we have
$\Delta_{g}f_{0}<n.$
Let $x_{0}\in M^{2n}$ be a point such that $f_{0}(x_{0})=0$. By Green formula (4.21), we obtain
$\int_{M^{2n}}(-f_{0})\omega^{n}$ $=$ $\int_{M^{2n}}\omega^{n}\cdot\int_{M^{2n}}\Delta_{g}f_{0}(y)G(x_{0},y)\omega^{n}(y)$
$<$ $n\int_{M^{2n}}\omega^{n}\cdot\int_{M^{2n}}G(x_{0},y)\omega^{n}(y)$
$<$ $C.$
Hence, $\|f_{0}\|_{L^{1}(g)}$ is bounded.
Formula (4.15) is similar to (5.4) in [46]. The kernel of (4.15) consists of the harmonic $1$-forms. If $a(f_{t})$ is in the kernel of (4.15), we have
$\|da(f_{t})\|^{2}_{L^{2}(g_{t})}$ $=$ $\int_{M^{2n}}da(f_{t})\wedge*_{t}da(f_{t})$
$=$ $-\frac{1}{(n-2)!}\int_{M^{2n}}da(f_{t})\wedge da(f_{t})\wedge\omega^{n-2}_{t}$
$=$ $0.$
Since $d^{*_{t}}a(f_{t})=0$, we see that $a(f_{t})$ is harmonic with respect to $g_{t}$. For $n=2$, one can use self-dual equation to prove Theorem 4.1. Moreover, recall that $f_{1}$ is defined by
$\frac{1}{n}\Delta_{g_{1}}f_{1}=\frac{\omega_{1}^{n-1}\wedge(\omega_{1}-\omega)}{\omega_{1}^{n}}$
which can be rewritten as
$\Delta_{g_{1}}f_{1}=n-\frac{1}{2}{\rm tr}_{g_{1}}g.$ (4.22)
Notice that $f_{1}$ is the same as $\varphi$ defined in [46]. As the discussion of Tosatti-Weinkove-Yau in [46], we have
Theorem 4.7.
(cf. [46, Theorem 1.3]) Let $(M^{2n},\omega,J,g)$ be a closed almost Kähler manifold of dimension $2n$. Let $\alpha>0$ be given. Let $F\in C^{\infty}(M)$ satisfying
$\int_{M^{2n}}e^{F}\omega^{n}=\int_{M^{2n}}\omega^{n}.$
Then if
$\omega_{1}$ $=$ $\omega+\mathcal{D}^{+}_{J}(f)$
$=$ $dJdf_{1}+da(f_{1})$
is an almost Kähler form cohomologous to $\omega$ solving
$\omega_{1}^{n}=e^{F}\omega^{n},$
there are $C^{\infty}$ a priori bounds on $\omega_{1}$ depending only on $\omega,J,F,\alpha$ and
$I_{\alpha}(f_{1})=\int_{M^{2n}}e^{-\alpha f_{1}}dvol_{g}.$
5 Further Remarks and Questions
If $\omega$ is a Kähler form, let $K_{\omega}$ be the space of $\omega$-compatible complex structures, $AK_{\omega}$ the space of $\omega$-compatible almost complex structures. It is easy to see that $K_{\omega}\subsetneq AK_{\omega}$, where $AK_{\omega}$ is viewed as a contractible Fréchet manifold equipped with a formal Kähler structure. Suppose that $(\omega,J,g)$ is an almost Kähler metric on closed manifold $M^{2n}$. Let $\mathcal{H}(\omega,J)$ be the almost Kähler potential space, that is,
$\mathcal{H}(\omega,J):=\{f\in C^{\infty}(M)\,\,|\,\,\omega+\mathcal{D}^{+}_{J}(f)>0\}.$ (5.1)
It is similar to the Kähler potential space, $\mathcal{H}(\omega,J)$ admits a natural Riemann metric of non-positive sectional curvature in the same sense [17]. Recall that, for Kähler setting, it was discovered by Mabuchi [35] and rediscovered by Semmes [39] and Donaldson [17]. The real numbers act on $\mathcal{H}(\omega,J)$, by addition of constants, and we define $\mathcal{M}_{[\omega]}=\mathcal{H}(\omega,J)/\mathbb{R}$, which can be viewed as the space of almost Kähler metrics on $M^{2n}$, in the given cohomology class. By the method of Mabuchi [35], we can show that in fact $\mathcal{M}_{[\omega]}$ is isometric to the Riemann product $\mathcal{H}_{0}\times\mathbb{R}$, and both he and Donaldson point out that understanding geodesics in these spaces is important for study of the space of almsot Kähler metrics. They raise the obvious question of wether any pair points in $\mathcal{H}(\omega,J)$ (or $\mathcal{M}_{[\omega]}$) can connected by a smooth geodesic. Lempert and Vivas give a negative answer in [33]. Chen proves that the space is at least convex by $C^{1,1}$ geodesics [10]. Darvas and Lempert show that the regularity that Chen obtains cannot be improved [22]. Since $K_{\omega}\subsetneq AK_{\omega}$, and $\mathcal{H}(\omega,J)$ is the almsot Kähler potential space, it is natural to investigate almost Kähler geometry by using $\mathcal{D}^{+}_{J}$ operator and the generalized Monge-Ampère equation. For almost Kähler setting, geodesics in $\mathcal{H}(\omega,J)$ are related to a generalized Monge-Ampère equation (cf. [39]) as follows. Let $S=\{s\in\mathbb{C}\,\,|\,\,0<{\rm Im}s<1\}$, and let $\tilde{\omega}$ the pullback of $\omega$ by the projection $\bar{S}\times M^{2n}\rightarrow M^{2n}$. With any smooth curve $0,1]\ni t\mapsto v_{t}\in\mathcal{H}(\omega,J)$, associate the smooth function $u(s,x)=v_{{\rm Im}s}(x)$, $(s,x)\in\bar{S}\times M^{2n}$. Then $t\mapsto v_{t}$ is a geodesic if and only if $u$ satisfies
$(\omega+\mathcal{D}^{+}_{\tilde{J}}(u))^{n+1}=0,$ (5.2)
where $\tilde{J}=(J_{st},J)$ is an almost complex structure on $\mathbb{C}\times M^{2n}$, $\tilde{J}$ is the standard complex structure. In order to study almost Kähler geometry, we need extend Futaki invariant [25], Mabuchi functional [34, 35], Tian $\alpha$-integral [44] and Ding functional [16].
As in Kähler case [20], on almost Kähler manifolds, we can consider existence question for four different kinds of special almost Kähler metrics, working within a fixed symplectic class on a compact manifold.
1) Extremal almost Kähler metrics
As in Kähler setting, extremal almost Kähler metrics are critical points of Calabi functional
$(\omega,J)\mapsto\int_{M^{2n}}R(J)^{2}\frac{\omega^{n}}{n!},$ (5.3)
where $R(J)$ is the Hermitian scalar curvature with respect the almost Kähler metric. It is interesting to consider the uniqueness of extremal almost Kähler metrics. For Kähler case, see [5]. Also, one can consider the relation between extremal Kähler metrics and extremal almost Kähler metrics. In the toric case, the existence of an extremal Kähler metric is conjecturallye equivalent to the existence of nonintegrable extremal almost Kähler metric [18] (see also [3]).
2) Constant Hermitian scalar curvature almost Kähler metrics (CHSC for short)
These metrics are just those whose Hermitian scalar curvature $R$ is constant. The constant Hermitian scalar curvature almost Kähler metrics are centainly extremal. Recently, Keller and Lejmi study $L^{2}$-norm of the Hermitian scalar curvature [30].
3) Hermite-Einstein almost Kähler metrics with $c_{1}=0$, $c_{1}<0$ and $c_{1}>0$
An almost Kähler metric $(\omega,J,g)$ is called Hermite-Einstein (HEAK for short) if the Hermite-Ricci form $\rho$ is a (constant) multiple of the symplectic form $\omega$, i.e.
$\rho=\frac{R}{2n}\omega,$
so the Hermitian scalar curvature $R$ is constant (cf. [32]). The form $\rho$ is a de Rham representative of $2\pi c_{1}(M,J)$ in $H^{2}(M,\mathbb{R})$ where $c_{1}(M,J)$ is the first (real) Chern class. If we suppose that $\omega$ and $\tilde{\omega}$ are symplectic forms compatible with the same almost-complex structure $J$ and satisfy $\tilde{\omega}^{n}=e^{F}\omega^{n}$ for some real-valued function $F$ then
$\tilde{\rho}=\rho-\frac{1}{2}dJdF,$ (5.4)
[…]
$v$ $=$ $M(f)$ (5.5)
$=$ $\begin{vmatrix}1+\mathcal{D}^{+}_{J}(f)_{1\bar{1}}&\mathcal{D}^{+}_{J}(f)_{1\bar{2}}&\cdots&\mathcal{D}^{+}_{J}(f)_{1\bar{n}}\\ \mathcal{D}^{+}_{J}(f)_{2\bar{1}}&1+\mathcal{D}^{+}_{J}(f)_{2\bar{2}}&\cdots&\mathcal{D}^{+}_{J}(f)_{2\bar{n}}\\ \vdots&\vdots&\cdots&\vdots\\ \mathcal{D}^{+}_{J}(f)_{n\bar{1}}&\mathcal{D}^{+}_{J}(f)_{n\bar{2}}&\cdots&1+\mathcal{D}^{+}_{J}(f)_{n\bar{n}}\\ \end{vmatrix}$.
Here, by Theorem 4.1, we have
$\mathcal{D}^{+}_{J}(f)$ $=$ $dJdf+dd^{*}\sigma(f)$
$=$ $dJdf_{0}+da(f_{0}),$
where
$\frac{1}{n}\Delta_{g}f_{0}=\frac{\omega^{n-1}\wedge(\tilde{\omega}-\omega)}{\omega^{n}},$
and
$\left\{\begin{array}[]{ll}d^{*}a(f_{0})=0&,\\ &\\ d^{-}_{J}Jdf_{0}+d^{-}_{J}a(f_{0})=0&,\\ &\\ \omega^{n-1}\wedge da(f_{0})=0&.\end{array}\right$. (5.6)
Accordingly, we may replace $M(f)$ by $M(f_{0})$. Further more, if $\tilde{g}$ is a HEAK metric, that is, $\tilde{\rho}=\lambda\tilde{\omega}$, then there exists an $F\in C^{\infty}(M)$ such that
$\rho$ $=$ $\lambda\omega+d\alpha$
$=$ $\lambda\omega+dJdF+da(F).$
Then
$\lambda dJdf_{0}$ $=$ $\tilde{\rho}-\rho+dJdF+da(F)-\lambda da(f_{0})$ (5.7)
$=$ $-\frac{1}{2}(dJd\log M(f_{0}))^{(1,1)}+dJdF+da(F)-\lambda da(f_{0}).$
Then (5.7) implies that $\lambda f_{0}=\frac{1}{2}\log M(f_{0})-F+c$ since
$\lambda dJdf_{0}\wedge\omega^{n-1}=(-\frac{1}{2}dJd\log M(f_{0})+dJdF)\wedge\omega^{n-1}$
and
$\lambda\Delta_{g}f_{0}=\frac{1}{2}\Delta_{g}\log M(f_{0})-\Delta_{g}F.$
Therefore the existence of HEAK metrics on almost Kähler manifold $(M^{2n},\omega,J,g)$ is equivalent to solving
$\left\{\begin{array}[]{ll}\log M(f_{0})=f_{0}+F,&~if~~c_{1}(M,J)<0;\\ &\\ \log M(f_{0})=F+c,&~if~~c_{1}(M,J)=0;\\ &\\ \log M(f_{0})=-f_{0}+F,&~if~~c_{1}(M,J)>0.\end{array}\right$. (5.8)
Similar to Kähler case, we have the following statements for HEAK metrics. If $c_{1}(M,J)=0$, the existence of HEAK metrics on almost Kähler manifold $(M^{2n},\omega,J,g)$ is equivalent the existence of the solution of the generalized Monge-Ampère equation (3.1). If $c_{1}(M,J)>0$, this is an important question which is regared as the generalized Yau-Tian-Donaldson conjecture for symplectic Fano manifolds. For Kähler Fano manifolds, see Chen-Donaldson-Sun[11], Tian [45] and others.
4) The generalzied almost Kähler-Ricci solitons (GeAKRS for short)
For Kähler-Ricci solitons, there are a lot of paper for this topic, see [62, 54] and others. Inoue [29] fixed a compact subgroup $G$ of $Ham(M,\omega)$ and consider the group $Ham^{G}(M,\omega)$ of Hamiltonian symplectomorphisms commuting with $G$. Let $AK^{G}_{\omega}$ be the space of $G$-invariant almost complex structures compatible with $\omega$. He proved that the action of $Ham^{G}(M^{,}\omega)$ on $AK^{G}_{\omega}$ with moment map given by
$R^{G}_{\xi}(J)=R^{G}(J)-n-2\Delta_{g}f_{\xi}+2f_{\xi}-2|f_{\xi}|^{2}_{g},$ (5.9)
where $f_{\xi}$ is a Hamiltonian potential of $\xi$ which is a fixed element in the center of the Lie algebra of $G$ on a compact symplectic Fano manifold. Inoue proved that the zeros of this moment map correspond to Kähler-Ricci solitons (cf. [29, Proposition 3.2]). If an almost Kähler metric $(\omega,J,g)$ satisfies the condition $R^{G}_{\xi}(J)=0$, we call it generalized almost-Kähler-Ricci soliton (GeAKRS for short). For the study of GeAKRS, see [1].
Acknowledgements.  The third author is very grateful to his advisor Z. Lü for his support; the authors thank Haisheng Liu for some helpful discussions.